### v1

# Enhancing Object Detection Accuracy in Underwater Sonar Images through Deep Learning-based Denoising

Abstract
Sonar image object detection is crucial for underwater robotics and other applications. However, various types of noise in sonar images can affect the accuracy of object detection. Denoising, as a critical preprocessing step, aims to remove noise while retaining useful information to improve detection accuracy. Although deep learning-based denoising algorithms perform well on optical images, their application to underwater sonar images remains underexplored.
This paper systematically evaluates the effectiveness of several deep learning-based denoising algorithms, originally designed for optical images, in the context of underwater sonar image object detection. We apply nine trained denoising models to images from five open-source sonar datasets, each processing different types of noise. We then test the denoised images using four object detection algorithms. The results show that different denoising models have varying effects on detection performance. By combining the strengths of multiple denoising models, the detection results can be optimized, thus more effectively suppressing noise.
Additionally, we adopt a multi-frame denoising technique, using different outputs generated by multiple denoising models as multiple frames of the same scene for further processing to enhance detection accuracy. This method, originally designed for optical images, leverages complementary noise-reduction effects. Experimental results show that denoised sonar images improve the performance of object detection algorithms compared to the original sonar images.
Index Terms:
Image Denoising, Underwater Sonar Images, Object Detection, Multi-frame DenoisingI Introduction
Underwater sonar imaging plays an indispensable role in marine exploration and various ocean industries, providing valuable insights into underwater environments. Unlike optical imaging, where light propagation is restricted, sonar systems utilize sound waves that travel farther, allowing them to cover larger underwater areas. This makes sonar images an ideal choice for applications such as seabed mapping, underwater object detection, and navigation. However, despite the advantages of sonar imaging, its image quality is often severely compromised by noise, which negatively impacts the accuracy of downstream tasks, such as object detection.
In sonar images, noise can originate from various factors, including environmental interference, sensor imperfections, and the inherent characteristics of sound wave propagation in water. Common types of sonar image noise include Gaussian noise, speckle noise, and Poisson noise. Gaussian noise typically arises from random fluctuations in sensor readings or environmental changes. Speckle noise, caused by sound wave scattering, manifests as granular interference, which can obscure object boundaries. Poisson noise, a type of noise induced by the Poisson distribution in statistics, primarily stems from the natural variability of the signal, particularly when the signal generation process is random. These noise types severely degrade the signal-to-noise ratio (SNR) of sonar images, resulting in poor image visibility and consequently affecting the performance of object detection tasks.
A standard approach for improving the quality of sonar images and enhancing object detection accuracy is through denoising. Denoising algorithms aim to reduce or eliminate noise while preserving critical features in the image. Traditional denoising methods, such as spatial filtering and thresholding, often perform poorly in underwater sonar image processing due to the complexity of the noise. These methods may blur or distort important details, such as object contours and features, which are vital for accurate detection. As a result, an increasing number of studies have turned to more advanced techniques, particularly deep learning-based methods, which can adapt to the unique characteristics of underwater sonar images.
Deep learning-based denoising technologies, especially self-supervised learning models, have demonstrated significant potential in handling image noise. These algorithms do not rely on large amounts of labeled data but instead learn to denoise by utilizing the inherent structure of the data itself. These models are particularly well-suited for scenarios where clear ground-truth images are unavailable, such as in underwater sonar imaging. In addition to directly denoising the original images, another promising approach is to treat multiframe denoised images as different exposures of the same scene and apply multi-frame image denoising techniques to further reduce noise. This technique integrates information from multiframe denoised images, thereby reducing noise while preserving more details and structural information. Multi-frame image denoising has been widely studied and applied in the field of optical images, demonstrating significant advantages in improving image quality. Although this method has been relatively less explored in underwater sonar images, its potential in underwater environments is equally noteworthy, especially when dealing with sonar images contaminated by noise.
This paper proposes an integrated approach to systematically evaluate the impact of denoising and multi-frame image denoising strategies on underwater sonar image object detection performance. We not only focus on how denoising techniques can improve image quality but also explore the application of multi-frame image denoising and assess its effectiveness in enhancing object detection accuracy and robustness.
We conducted experiments on several underwater datasets, including FDD[39], UATD[38], SCTD[40], MarineDebris[41], and 2020Contest[42]. These datasets feature diverse underwater scenes and various target types, enabling an effective evaluation of the performance of denoising, object detection, and image fusion methods . We combined several advanced deep learning techniques, employing multiple denoising algorithms and object detection models to assess their performance on underwater sonar images. Specifically, we used three self-supervised denoising algorithms Neighbor2Neighbor[2], Blind2Unblind[5], and DSPNet[6], aimed at removing noise from sonar images and restoring critical image information. For object detection, we evaluated four different models: SSD-MobileNetV2[27][47], Faster R-CNN[31], SSD300[27], and YOLOX[45]. These models feature distinct network architectures, providing valuable insights into the performance of different object detection frameworks on denoised sonar images. Additionally, we employed the OPD[46] multi-frame image denoising algorithm to further denoise multiple denoised images, enhancing target visibility and reducing noise, thereby improving detection accuracy.
The contributions of our work are as follows:
Through comprehensive experiments, we conducted a comparative analysis of the differences in object detection results among the original dataset, the dataset with applied denoising algorithms, and the dataset with both denoising and multi-frame image denoising algorithms. This provides important reference for subsequent research in the field of underwater sonar image object detection.
We analyzed the effects of three advanced deep learning denoising algorithms on object detection in sonar images, offering valuable insights for future research in this domain.
We innovatively treated the denoised sonar images as multiple exposures of a single scene and input them into a multi-frame image denoising algorithm for integration. Experimental results validated the effectiveness of this approach.
II Related Work
II-A Object Detection
The methods for object detection in underwater sonar images can be categorized into traditional methods, CNN-based methods, and other machine learning methods. Traditional methods mainly rely on image processing and feature extraction techniques to accomplish object detection through manually designed rules and algorithms. Deep learning methods, particularly object detection techniques based on convolutional neural networks (CNN), have gradually demonstrated strong capabilities in sonar image processing. These methods learn features directly from data in an end-to-end manner, showing excellent adaptability and robustness. In addition to CNNs, traditional machine learning methods also play a role in object detection for underwater sonar images. These methods often combine handcrafted feature extraction with classifiers, offering flexibility but with performance limited by the quality of feature design.
Traditional Methods. Over the past decades, sonar image detection has predominantly relied on artificial feature extraction, mainly based on pixel values, features, and gray thresholds. Traditional underwater target detection methods typically depend on pixel value features, gray thresholds, or prior information regarding corresponding targets. For instance, [32] presents a hierarchical system that detects objects in underwater videos by transforming the input image into the spatial frequency domain and utilizing a frame-difference algorithm, demonstrating robust performance and accuracy across various environments. Another study [33] introduces a two-dimensional object segmentation method based on the OS-CFAR technique, designed to effectively extract targets from side-scan sonar data, showing improved resilience to noise and validating its effectiveness through experiments on real images. Additionally, a symbolic pattern analysis method [34] is proposed for robust feature extraction from sidescan sonar images, aimed at detecting underwater targets like mines, with real-time object recognition validated on actual naval operational data. Finally, another study [35] showcases preliminary results in change detection by surveying different seafloor environments, with the goal of developing an automated tool for detecting recently placed objects on the seabed.
CNN-Based Methods. In recent years, CNN-based methods have gained widespread adoption in underwater object detection. For instance, [24] employs CNN for object recognition in forward-looking sonar images, which significantly surpasses existing template matching methods, while requiring fewer learnable parameters to achieve high performance that generalizes well to unseen data. [15] introduces YOLOv3-DPFIN, an improved YOLOv3 [16] algorithm that employs a dual-path feature fusion network for real-time, noise-resistant underwater target detection. This method enhances feature extraction capabilities and multi-scale prediction accuracy, significantly boosting detection speed and precision. [38] proposes an improved YOLOv5-based target detection model for forward-looking sonar images, utilizing transfer learning and CoordConv to enhance feature extraction, resulting in improved detection accuracy and speed. Moreover, the detection network UFIDNet proposed in [18] integrates speckle noise reduction and scene prior knowledge to enhance detection performance, incorporating an auxiliary branch and feature selection strategy to optimize accuracy and efficiency. On the other hand, [14] presents EPL-UFLSID, a pseudo label-driven algorithm for underwater forward-looking sonar image object detection, which leverages high-quality pseudo labels to guide feature extraction for improved detection outcomes. Meanwhile, [19] introduces a semi-supervised framework based on a multibranch convolutional encoder-decoder network (MB-CEDN) for multitarget detection and segmentation in circular-scan synthetic aperture sonar (CSAS) images, significantly enhancing both detection and segmentation accuracy. Yang et al. [62] proposed a lightweight underwater object detection network, FLSD-Net, which enhances the detection accuracy and speed of small underwater objects by addressing overfitting and generalization issues, thereby providing greater robustness for practical applications.The following two studies explore the application of improved YOLOv7 models in underwater object detection. Qin et al. [63] introduced the YOLOv7C model, which incorporates an attention mechanism and the Multi-GnBlock module to enhance object detection performance in underwater sonar images with complex backgrounds. Additionally, pruning techniques were applied to improve real-time performance. Wen et al. [65], focusing on side-scan sonar images, integrated the Swin-Transformer and CBAM modules to strengthen the model’s attention to target regions. Furthermore, they innovatively introduced a feature scale factor, significantly improving detection accuracy.Both studies demonstrate the significant advantages of the improved YOLOv7 models in underwater object detection tasks.
Other Machine Learning Methods. In traditional methods for underwater object detection, [22] presents a cascade of boosted classifiers based on the Viola and Jones algorithm, emphasizing in-situ learning of target responses and local clutter. This approach enhances detection accuracy while minimizing false alarm rates in complex terrains, making it suitable for real-time applications. The method not only processes images rapidly but also adapts effectively to various complex environments. Additionally, reference [12] introduces an automatic detection method that combines clustering, segmentation, and feature extraction to achieve efficient and low-error detection without complex training processes.
II-B Denoising
Image denoising technology is crucial in the field of image processing, especially in noisy environments such as underwater sonar images. With the rapid development of deep learning, convolutional neural network-based methods have gradually become mainstream in the image denoising field. Deep learning methods offer significant advantages in denoising performance. For instance, RED-Net [48] proposed a deep fully convolutional network architecture with skip-layer connections to enhance image denoising performance. DnCNN [49] trains deep convolutional neural networks to effectively handle images with various noise intensities, making it particularly suitable for blind denoising tasks. FFDNet [50] introduces a noise level map, allowing the same network to handle images with varying noise levels, significantly improving inference speed. Additionally, CBDNet [51] adopts a more realistic noise model, combining synthetic and real noise images for training, effectively handling complex noise in real-world applications. In 2022, Chunwei Tian et al. [52] proposed MWDCNN, which combines convolutional neural networks with wavelet transform techniques and utilizes a multi-stage design to improve denoising performance. This method introduces dynamic convolution, allowing linear combinations between different convolution kernels to balance denoising performance and computational cost. In 2024, they proposed CTNet [53], which combines serial blocks, parallel blocks, and residual blocks, integrating the Transformer mechanism to further improve denoising performance. This model is more effective in handling complex noise patterns and is suitable for denoising tasks in large-scale datasets.
Unsupervised and self-supervised learning methods offer new ideas for image denoising, as they do not rely on clean images for training. For example, the Noise2Noise [54] method trains using pairs of noisy images, enabling effective recovery of clean images from noisy inputs. Noise2Self [55] and Noise2Void [56] utilize self-supervised learning, which does not require clean images or paired images. Instead, they perform denoising by exploiting the statistical properties of noise in the images. The Noisy-As-Clean [57] method treats noisy images as “clean” during training, effectively reducing the discrepancy between the noise and image priors, thereby improving denoising performance. Furthermore, Neighbor2Neighbor [2] and Blind2Unblind [5] are two recently proposed self-supervised denoising methods that can perform efficient denoising without the need for clean images.
For specific denoising issues in medical imaging and other domains, researchers have proposed plug-and-play denoising network modules. For example, LIDnet [58] integrates denoising tasks with lesion detection tasks and introduces an interest-region perception loss based on lesion detection, along with a collaborative training strategy, significantly improving denoising results and detection accuracy. Similarly, the DN-DETR [59] method incorporates a denoising training strategy to enhance the convergence speed and performance of the object detection method DETR.
In the field of underwater sonar image denoising, the UFIDNet [60] network proposed a method that combines speckle noise reduction auxiliary branches with feature selection strategies, significantly improving the accuracy and speed of underwater forward-looking sonar object detection. Meanwhile, Wang, X. et al. [61] proposed a denoising method based on the golden ratio for non-local spatial information and an adaptive cultural algorithm (NACA), showing better effectiveness and adaptability in both denoising and object detection tasks.
Overall, the denoising technology for underwater sonar images is still in the early stages of development. Future research needs to delve deeper into how to handle the various complex types of noise in underwater sonar images and better apply the mature denoising techniques developed for optical images to sonar images.
III Background
III-A Sonar Images
In seawater, due to changes in the medium, light waves and radio waves experience significant attenuation, resulting in very limited propagation distances. In contrast, sound waves propagate much better in water, covering a broader marine area. This makes sonar images more suitable for ocean exploration and various industries.
Sonar images can be classified into two types: side-scan sonar images and forward-looking sonar images. Side-scan sonar images are mainly used for exploring the seabed, generating images that present a lateral cross-section of the underwater terrain, typically used for target identification and underwater object exploration. On the other hand, forward-looking sonar images are used for scanning objects in front of the underwater vehicle, typically forming images in front of submarines or underwater robots, providing a more intuitive three-dimensional view to assist in identifying obstacles or targets ahead.
Typically, sonar images consist of seven main components: the sonar pulse source, surface reflections, water column clutter, seabed reflections, water column, target objects, and shadows. The process of image formation by sonar equipment generally includes the following steps: the transducer array emits a sonar pulse and listens for returning echoes. The returned sound waves are converted from acoustic energy into electrical energy, which is transmitted via a towed cable to the recording and display unit onboard the surface vessel. At the recording and display unit, the signal is processed into the desired format and then sent to monitors and recorders, resulting in a recorded image. The main characteristics of sonar images are as follows:
Low Resolution. Due to the long wavelength of sound waves, sonar images typically have lower resolution, resulting in blurry details.
Noise Interference. Sonar images are often affected by various noise sources, such as speckle noise and multipath effects, creating noticeable noise patterns.
Low Signal-to-Noise Ratio (SNR). Given the complex underwater environment, sonar imaging often has a low SNR, making the contours of objects blurry and harder to identify.
Depth Information. Sonar images provide valuable depth information about objects, facilitating the identification and localization of targets in a 3D underwater environment.
Fig. 1 presents two images: the top one is a side-scan sonar image, while the bottom one is a forward-looking sonar image. In the side-scan sonar image, the black area in the center represents sonar shadows (as the detector emits sound waves to both sides, the area directly beneath receives no echoes, forming a black region). The bright yellow areas on both sides indicate the returned sonar signals, revealing details of the seabed terrain, such as sand dunes or other geological features. Since side-scan sonar scans the seabed from the sides, it provides a wide coverage area, making it suitable for underwater terrain surveying and target detection. The forward-looking sonar image contains numerous radial streaks, which may be noise caused by underwater target reflections or sound wave interference. The primary purpose of forward-looking sonar is to detect obstacles ahead of underwater vehicles or divers, providing real-time navigation information.
III-B Noise in sonar images
Sonar images are affected by a variety of noise types, including Gaussian noise, Poisson noise, speckle noise, and other environmental noises. Each type of noise impacts the quality of sonar images and the accuracy of object detection in distinct ways.
Gaussian Noise. Gaussian noise typically arises from sensor instability and environmental factors, and it is randomly distributed throughout the image. This type of noise significantly affects the overall image quality, especially under low-light conditions. Due to its randomness, Gaussian noise is evenly spread across the image, blurring details of the target and making precise object recognition and detection more challenging.
Poisson Noise.Poisson noise occurs primarily in low-light or low-signal scenarios, with its intensity proportional to the signal strength. It is especially noticeable in underexposed regions of the image. In sonar images, Poisson noise can reduce the visibility of the signal, blurring target edges and making image processing and object detection more difficult.
Speckle Noise. Speckle noise is a common noise type in sonar images, generally caused by sound wave reflection and scattering. It appears as granular interference, giving the image an uneven texture and causing blurred target boundaries. Speckle noise significantly impacts image detail, especially in high-resolution images, and can impair the accuracy of object detection.
Other Noises. Apart from the primary noise types, sonar images may also contain other noises, such as multipath reflection noise and environmental noise. Multipath reflection noise arises from variations in the reflection paths of sound waves, potentially causing image blurring or false targets. Environmental noise may stem from underwater biological activity, ship vibrations, and other external interferences, adding to the complexity of noise in sonar images. These various types of noise present significant challenges for object detection in underwater sonar images.
III-C Typical Object Detection Algorithms
The development of object detection algorithms has gone through several important stages, gradually transitioning from traditional detection methods to real-time detection technologies based on deep learning. YOLO Series. YOLO [24] was proposed by Joseph et al. in 2015 and is the first one-stage detector in the deep learning era. YOLO is extremely fast, following a completely different paradigm from two-stage detectors. It applies a single neural network to the full image, dividing the image into regions and predicting bounding boxes and probabilities for each region simultaneously. Despite its significant improvement in detection speed, YOLO suffers from a drop in localization accuracy, especially for small objects. Subsequent versions of YOLO (such as YOLOv4 [25] and YOLOv7 [26]) focus on addressing this issue. YOLOv7 introduces optimized structures like dynamic label assignment and model structure reparameterization, outperforming most existing object detectors in speed and accuracy. YOLOv5 [44] builds on YOLOv4 by optimizing model architecture and hyperparameters, making training and deployment more streamlined while enhancing speed and accuracy. In addition, YOLOX [45], a newer member of the YOLO family, incorporates anchor-free mechanisms and decoupled heads, further boosting its detection capabilities while maintaining high efficiency. SSD Series. SSD [27] was proposed by Liu et al. in 2015. Its main contribution is the introduction of multi-reference and multi-resolution detection techniques, significantly improving the detection accuracy of a one-stage detector, especially for small objects. SSD detects objects of different scales on different layers of the network, whereas previous detectors only ran detection on their top layers. R-CNN Series. The idea behind R-CNN [28] is straightforward. It begins with the extraction of a set of object proposals (candidate boxes) using selective search. Each proposal is then rescaled to a fixed size and fed into a CNN model pretrained on ImageNet (e.g., AlexNet[43]) to extract features. Finally, linear SVM classifiers are used to predict the presence of an object within each region and recognize object categories. However, the drawbacks are evident: redundant feature computations on a large number of overlapping proposals (over 2000 boxes from one image) lead to an extremely slow detection speed. Later in the same year, SPPNet [29] was proposed to solve this problem. In 2015, Girshick proposed Fast R-CNN [30], a further improvement of R-CNN and SPPNet. Fast R-CNN enables simultaneous training of a detector and a bounding box regressor under the same network configurations. Although Fast R-CNN successfully integrates the advantages of R-CNN and SPPNet, its detection speed is still limited by proposal detection. Ren et al. proposed Faster R-CNN [31] shortly after Fast R-CNN. Faster R-CNN is the first near-real-time deep learning detector. Its main contribution is the introduction of a Region Proposal Network (RPN), enabling nearly cost-free region proposals.
IV Methodology
The variety and abundance of natural image datasets have fueled the development of advanced denoising methods, especially self-supervised algorithms. These algorithms do not rely on large-scale labeled datasets, making them well-suited for noise removal tasks in scenarios with scarce labeled data. Underwater sonar imaging faces significant challenges due to the diverse sonar devices, complex underwater environments, and varying noise patterns. These noise types, such as speckle noise, gaussian Noise, and poisson noise, significantly degrade the quality of sonar images, complicating subsequent analysis and object detection tasks. Unlike natural image datasets, which often include clean or minimally noisy images, obtaining noise-free sonar images is exceptionally difficult. This limitation necessitates the adaptive application of existing denoising methods. Consequently, these algorithms require tailored evaluation and adjustment to verify their effectiveness in sonar imaging. We designed a method to verify whether applying denoising algorithms originally developed for optical images to sonar image datasets can improve object detection accuracy and whether performing multi-frame denoising on images processed by different algorithms can further enhance detection precision.
Fig. 2 presents the proposed method designed to validate the above questions. This section introduces the proposed methodology for validating the impact of denoising techniques and fusion strategies on underwater sonar image object detection. The workflow is divided into three major steps: baseline object detection, object detection on denoised datasets, and object detection on multi-frame denoised datasets. Each step is detailed below to explain the overall process and its rationale.
First, multiple original underwater sonar image datasets (Dataset1, Dataset2, …, DatasetX) are processed using several object detection algorithms (Object Detection Algorithm1, Algorithm2, …, AlgorithmX) to establish a baseline for comparison. Next, the original datasets are denoised using state-of-the-art self-supervised denoising algorithms (Self-Supervised Denoising Algorithm1, Algorithm2, …, AlgorithmX), generating denoised datasets (Denoised Dataset1, Denoised Dataset2, …, Denoised DatasetX). These denoised datasets are subsequently evaluated using the same object detection algorithms. Finally, multi-frame image denoising techniques are employed to further denoise the denoised dataset, generating multi-frame denoised datasets (Multi-frame Denoised Dataset 1, Multi-frame Denoised Dataset 2, …, Multi-frame Denoised Dataset X). Object detection is then performed on these multi-frame denoised datasets to assess potential improvements in detection accuracy and robustness. This method systematically explores how denoising and multi-frame denoising can enhance detection outcomes while preserving critical information in sonar images.
IV-A Object Detection on Original Datasets
The first step involves processing original underwater sonar image datasets to establish a baseline for object detection performance. To establish a performance baseline, we collect sonar image datasets using various sonar devices and in multiple underwater scenarios. These datasets represent the diversity of noise types and environmental conditions encountered in real-world applications. The effectiveness of object detection algorithms is evaluated on these datasets, enabling us to identify the challenges posed by different noise patterns and device characteristics.
IV-B Object Detection on Denoised Datasets
Given that underwater sonar images often lack clean, noise-free ground truths, self-supervised denoising algorithms are utilized. These algorithms (e.g., Self-Supervised Denoising Algorithm1, Algorithm2, …, AlgorithmX) are specifically chosen for their ability to learn noise patterns without relying on clean ground-truth images. The denoised datasets are subsequently processed using the same object detection algorithms applied in the baseline step. Comparing detection results on denoised datasets with the baseline allows us to quantify each denoising algorithm’s effectiveness. This step examines the impact of noise reduction on object detection performance and evaluates whether certain algorithms perform better for specific sonar noise types.
IV-C Enhancing Datasets with Multi-Frame Denoising Models
To further improve object detection performance, we consider the denoised datasets as multiple frames of the same scene captured in sequence. This approach is based on the premise that applying different denoising algorithms to the same image produces multiple versions, each with unique strengths in noise suppression and detail preservation, resembling observations of the same scene under varying conditions. Therefore, we utilize these denoised datasets as training data to train a deep learning-based multi-frame denoising model, enabling it to extract stable features from diverse denoised versions while enhancing noise adaptability.
Once trained, this model is applied to the original sonar dataset for denoising, aiming to reduce noise while preserving critical information essential for object detection. Finally, object detection is performed on the model-denoised dataset to evaluate its impact on detection accuracy and robustness. This multi-frame denoising strategy not only leverages the complementary strengths of different denoising algorithms but also enhances the model’s generalization ability to complex underwater sonar noise, providing a more effective data preprocessing solution for underwater object detection.
By systematically comparing the results from original, denoised, and multi-frame denoised datasets, this methodology provides valuable insights into the interplay between denoising, multi-frame denoising, and object detection performance in underwater sonar imagery.
V Experiments
In this section, we conduct experiments based on the method proposed in Section 3 to evaluate the following two questions.
1.Whether applying denoising algorithms originally designed for optical images to sonar image datasets can improve object detection accuracy.
2.Whether treating the denoised dataset as multi-frame images and applying a multi-frame denoising algorithm can further improve object detection accuracy.
V-A Experimental Settings
To fairly evaluate the proposed questions, we utilized open-source underwater sonar image datasets, including FDD[39], UATD[38], SCTD[40], MarineDebris[41], and 2020Contest[42]. In SCTD [40], there are three types of targets: 461 shipwrecks , 90 aircraft, and 45 human figures. The contours and details of SCTD images are clear; however, interferences such as fragmentation, fractures, and geometric distortions can introduce additional challenges for detection. The UATD [38] dataset was collected using the Tritech Gemini 1200ik sonar from lakes and shallow waters. It contains over 9,000 multi-beam forward-looking sonar (MFLS) images. The 2020Contest [42] dataset includes both forward-looking and side-scan sonar images. The training set contains 2,288 images, with 1,216 side-scan sonar images, 800 forward-looking sonar images, and 272 negative samples with no targets. We selected the forward-looking sonar images for the experiment, as they provide a real-time, directional view of the forward environment, making them suitable for rapid detection and obstacle avoidance tasks, while side-scan sonar images are more suited for large-area, planar surveys, with lower accuracy and real-time performance. The Marine Debris dataset [41] was collected in the Ocean Systems Lab Water Tank at Heriot-Watt University, using an ARIS Explorer 3000 Forward-Looking Sonar operating at a 3.0 MHz frequency. The dataset includes various categories of marine debris, such as bottle, can, chain, drink carton, hook, propeller, shampoo bottle, standing bottle, tire, and valve. Each image is annotated with bounding boxes, and the annotations are stored in a corresponding JSON file. The Forward Looking Sonar Detection Dataset [39] is collected using the Oculus M1200d sonar, and it includes three categories: boat, plane, and victim, with a total of 3,752 images.
We chose three denoising methods [2][5][6] to comprehensively validate whether denoising techniques can improve object detection accuracy in underwater sonar images. Neighbor2Neighbor [2] provides a self-supervised approach to training an image denoising model using only noisy images, avoiding the need for large-scale clean-noisy image pairs. This method focuses on generating effective training pairs and reducing noise while maintaining the similarity of pixel appearances, making it suitable for underwater sonar images with varying noise distributions. Blind2Unblind [5] addresses the issue of information loss in self-supervised denoising by using a global-aware mask mapper and a re-visible loss function, which allows the network to capture important details in noisy images. This method performs well on both synthetic and real-world datasets, showing superior performance in denoising underwater sonar images. DSPNet [6] addresses the problem of speckle noise in sonar imaging, which significantly affects object detection. By using a feature pyramid network (FPN) and atrous spatial pyramid pooling (ASPP), DSPNet effectively reduces speckle noise while preserving key image details. The multi-scale mixed loss function further enhances its performance, making it suitable for underwater sonar images.
In the multi-frame denoising approach, we chose the multi-frame denoising algorithm OPD [65] because it provides a unified, unsupervised method that effectively leverages multiple noisy frames to enhance image quality. OPD learns through mutual supervision among frames, capturing inter-frame correlations and reducing noise without relying on clean labels. This strategy ensures adaptive information preservation, allowing critical details in different frames to be retained while enhancing robustness against noise.
We conducted object detection experiments using MMDetection, an open-source object detection toolbox developed by OpenMMLab, implemented based on PyTorch. This library provides a highly convenient platform for researchers and engineers to train and infer models.
To comprehensively evaluate the impact of denoising algorithms on object detection performance in underwater sonar images, we carefully selected four representative object detection algorithms: SSD-MobileNet v2, Faster R-CNN, SSD300, and YOLOX. These algorithms cover a diverse range of detection paradigms, enabling a robust assessment of how denoising influences detection accuracy, robustness, and computational efficiency.
First, SSD-MobileNet v2 is included due to its lightweight structure and efficiency, making it highly suitable for real-time applications on resource-constrained devices. Its use of MobileNet v2 as a backbone provides a balanced trade-off between detection accuracy and computational cost. This allows us to assess how denoising improves the performance of lightweight models designed for practical deployment in mobile or embedded systems.
Second, Faster R-CNN is chosen for its role as a representative of two-stage detection frameworks. Its region proposal network (RPN) enables precise detection and localization of objects, especially in scenarios with complex backgrounds or overlapping targets, such as in sonar images. By including Faster R-CNN, we aim to evaluate how denoising affects high-accuracy models that prioritize precision over speed.
Third, SSD300, as a single-stage detection algorithm, bridges the gap between speed and accuracy. With its ability to handle multi-scale feature maps for detecting objects of varying sizes, SSD300 is particularly relevant for sonar images, where targets may vary in size and clarity due to environmental noise. This inclusion helps us analyze the impact of denoising on models that balance real-time performance and moderate precision.
Finally, YOLOX represents the latest advancements in one-stage object detection, offering improved accuracy and speed compared to traditional YOLO models. Its advanced feature extraction and anchor-free design make it highly effective for detecting objects in challenging environments. Using YOLOX enables us to investigate the influence of denoising on state-of-the-art models that excel in both speed and detection performance.
By evaluating these four algorithms, we can comprehensively study the interplay between denoising and object detection across various detection paradigms. This selection ensures that our analysis captures the effects of denoising on models with different computational demands, detection strategies, and target applications, thereby providing valuable insights for improving object detection in underwater sonar imagery.
The experiments were conducted on a PC running Ubuntu 22.04.5 LTS operating system. The PC is equipped with an Intel(R) Core(TM) i9-14900K processor with 32 core. It has 128 GB of RAM and two NVIDIA GeForce RTX 3090 GPUs, each with 24 GB of memory, totaling 48 GB. The GPU driver version is 535.183.01, and the CUDA version used is 12.2.
Our experiments can be broadly divided into three parts. The first part involves performing object detection on the original underwater sonar datasets. The second part focuses on object detection on the denoised underwater sonar datasets. The third part conducts object detection on the underwater sonar datasets after multi-frame denoising.
V-B object detection on the original underwater sonar datasets
The detection results on the five datasets, FDD, 2020contest, MarineDebris, SCTD, and UATD, using four detection algorithms with the MMDetection toolkit are shown in Figures 3–7.
The analysis of the experimental results on five underwater sonar datasets reveals significant performance variations across different object detection algorithms, highlighting the critical impact of noise characteristics and target features in sonar images. For the UATD and FDD datasets, all algorithms exhibit high detection performance, with YOLOX and Faster R-CNN achieving remarkable mAP@0.5:0.95 scores of 0.545 and 0.845, respectively. These results demonstrate the potential of deep learning models in scenarios with relatively clear target features. Specifically, on the FDD dataset, YOLOX achieved the highest performance at high IoU thresholds (mAP@0.75 of 0.974) while maintaining a high AR, indicating a balanced trade-off between precision and recall. This superior performance is likely attributed to the distinctive target features and lower noise levels in the UATD and FDD datasets, which facilitate the models’ ability to extract and utilize critical information effectively.
In contrast, the performance of all algorithms declines significantly on the Marine and 2020contest datasets, with much lower mAP@0.5:0.95 scores compared to UATD and FDD. This decline can be attributed to higher noise levels, more complex backgrounds, or smaller target sizes, which collectively increase the difficulty of detection. On the MarineDebris dataset, YOLOX achieves the best performance in mAP@0.5 (0.96), but it lags behind Faster R-CNN in mAP@0.75 and AR, suggesting that while YOLOX excels at detecting targets under low IoU thresholds, it may struggle with precise boundary detection under stricter IoU conditions. On the 2020contest dataset, Faster R-CNN and YOLOX achieve comparable mAP@0.5:0.95 scores of 0.271 and 0.278, respectively. However, YOLOX’s advantage in AR indicates stronger recall capabilities, whereas Faster R-CNN appears more adept at detecting high-confidence targets in complex noise conditions.
On the SCTD dataset, the overall performance of all algorithms is relatively low, indicating the presence of extreme noise levels or blurred target features, which further complicates detection. On this dataset, SSD300 achieves the highest mAP@0.75 score (0.61), demonstrating its capability for detection under moderate IoU thresholds. Although YOLOX maintains some advantages in mAP@0.5, its mAP@0.5:0.95 performance is slightly inferior, reflecting potential limitations in its generalization to multi-scale targets in high-noise, low-contrast scenarios. These findings highlight the distinctive strengths of different algorithms when dealing with complex sonar data, emphasizing the importance of selecting appropriate algorithms for specific datasets. Furthermore, the results underline the critical challenge of balancing noise reduction and target feature preservation in underwater sonar object detection.
V-C object detection on the denoised underwater sonar datasets
We reviewed the state-of-the-art deep learning denoising algorithms and ultimately selected three that met our experimental requirements, enabling us to carry out a comprehensive experiment. The three deep learning-based denoising algorithms we selected each have unique strengths that meet our diverse experimental requirements for denoising performance. To begin with, Neighbor2Neighbor [2] achieves effective noise reduction by training on pairs of neighboring pixels within the same noisy image. This approach not only reduces noise significantly but also enhances model performance through a regularizer, thereby minimizing dependency on noise distribution assumptions. This enables it to achieve strong denoising without the need for clean images, making it robust for real-world noisy data. Additionally, Blind2Unblind [5] introduces an innovative transition from blind to non-blind denoising. By employing a global perception mask mapper and a re-visible loss, it improves denoising performance, allowing for precise noise removal while preserving critical image details. This balance between noise reduction and detail preservation makes it highly effective in scenarios requiring high fidelity. Finally, DSPNet [6] is specifically designed for the reduction of specks of noise. It applies a logarithmic transformation to convert multiplicative noise into an additive form, while leveraging feature pyramid networks and atrous spatial pyramid pooling to improve image quality and retain essential details. Each of these algorithms offers distinct strengths, providing a comprehensive basis for our experiments and comparative analysis of denoising performance.
The Blind2Unblind algorithm can be divided into four parameters: (1) Gaussian noise with , (2) Gaussian noise with , (3) Poisson noise with , and (4) Poisson noise with . A denoising model can be trained for each type of noise and its corresponding parameters. Similarly, the Neighbor2Neighbor algorithm can also be divided into four parameters: (1) Gaussian noise with , (2) Gaussian noise with , (3) Poisson noise with , and (4) Poisson noise with . In addition to these, the denoising model trained by DSPNet is also included. In total, there are nine denoising models. In the inference stage, the dataset is input into the trained models to obtain the denoised underwater sonar image datasets.
Tables 2–29 present the detection results of five datasets evaluated across four object detection algorithms, including mAP (), mAP_50 (), mAP_75 (), mAP_s, mAP_m, mAP_l, AR, AR_s, and AR_l, where s, m, and l represent small targets (objects with areas smaller than ), medium targets (areas between and ), and large targets (areas larger than ), respectively. The term ”noisy” denotes detections performed on the original dataset without denoising.
The appendix Tables 2–5 present the detection results of the SCTD dataset using four object detection algorithms: SSD-MobileNetV2, Faster R-CNN, SSD300, and YOLOX. For the SSD-MobileNetV2 algorithm, DSPNet achieves the best overall performance in terms of mAP and AR, particularly excelling in medium-sized target detection, which highlights its ability to balance noise suppression and information preservation effectively. However, detection performance for small targets and complex categories (e.g., ”human”) is generally lower, likely due to the loss of fine details during the denoising process, which affects the model’s ability to identify small-scale targets. Additionally, some denoising methods result in mAP values of 0 or NaN for specific target categories or sizes, potentially due to the absence of such targets in the dataset or excessive smoothing that removes critical features. Interestingly, the detection of large targets on noisy data without denoising performs relatively well (mAP = 0.582), suggesting that some information embedded in the noise may aid target localization, but such information could be inadvertently removed during denoising. For Faster R-CNN and SSD300, the impact of denoising methods is even more pronounced, particularly in improving mAP_50. For example, in Faster R-CNN, the ”b2ub-g5-50” method achieves the highest mAP_50 (0.783) for the ”ship” category, while ”b2ub-p30” demonstrates strong performance in high IoU thresholds for the ”human” category (mAP_75 = 0.424). In SSD300, DSPNet achieves the highest overall mAP (0.494), while B2UB-P30 stands out under high IoU thresholds (mAP_75 = 0.639). Therefore, further research into optimization strategies tailored to different target categories and sizes is warranted to better balance noise reduction and information preservation.
Fig. 8 shows the mAP@0.5 and AR results of object detection on the original SCTD dataset and the SCTD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). For SSD-MobileNetV2, the b2ub-g25 model provides the highest mAP@0.5 value, but the AR performance is not the best. For Faster R-CNN, nei-p30 performs best in mAP@0.5 (0.642) and also provides a good AR value (0.431). On SSD300, models such as b2ub-g25 and nei-p30 show improvements in mAP@0.5, but the AR values do not follow a consistent trend. For YOLOX, the b2ub-g25 model performs strongly in mAP@0.5 (0.68), but the AR value does not reach the best performance, with the highest AR value observed in the nei-g25 model (0.482).
Tables 6-9 present the evaluation results of the 2020Contest dataset under four object detection algorithms. Overall, denoising algorithms generally lead to improvements in one or more evaluation metrics. For SSD-MobileNetV2, the performance on small object detection was initially poor, nearly zero, but after processing with eight denoising models, the mAP for small objects generally improved. In Faster R-CNN, although some denoising algorithms did not significantly increase the mAP, they showed a noticeable improvement in the AR for large objects. For SSD300, among all the denoising algorithms, nei-p30 performed the best, with an mAP of 0.223, showing a 0.006 improvement over the noisy images. The performance of b2ub-g5-50 after denoising was relatively stable, with an mAP of 0.219. Despite the small increase, there was an improvement in the recall rate for large objects (AR_l = 0.434). In contrast, b2ub-p5-50 showed a significant drop in mAP to 0.137, indicating that some denoising algorithms might over-denoise, leading to information loss and negatively impacting detection performance. YOLOX outperformed the other algorithms, with an overall mAP of 0.278 for noisy images. After denoising, b2ub-g25 performed the best, with improvements in all metrics except for mAP_m, which slightly declined.
Fig. 9 shows the results of mAP@0.5 and AR from object detection on the original 2020Contest dataset and the 2020Contest dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). For mAP@0.5 (represented by the bar chart), the results of most models are relatively close, with nei-p30 achieving the highest mAP@0.5 across all algorithms, ranging from 0.564 to 0.589. The b2ub-p30 model also performs consistently well across all detection algorithms, especially with a value of 0.585 on YOLOX. The dspnet model shows generally lower mAP@0.5 compared to other models, particularly on SSD-MobileNetV2, where it is 0.381. For AR (represented by the line chart), nei-p30 has the highest AR value across all algorithms, ranging from 0.369 to 0.373, indicating a good balance between accuracy and recall. The AR values of b2ub-p5-50 and b2ub-p30 are similar, ranging from 0.365 to 0.371, while dspnet has relatively lower AR values. Overall, nei-p30 performs the best in both mAP@0.5 and AR, achieving the highest scores in almost all cases. Other models such as b2ub-p30 and nei-g25 show good performance in mAP@0.5, but their AR values are slightly lower. The performance of b2ub-p5-50 and nei-p5-50 models varies across different algorithms.
Table 10-17 show the results of four object detection algorithms on the MarinDebris underwater sonar dataset under different denoising algorithms and parameter settings. For the SSD-MobileNetV2 algorithm, the performance of the dspnet and b2ub-g25 denoising models was relatively poor, with most metrics showing a decline. On the other hand, b2ub-p30, b2ub-p5-50, and nei-p5-50 performed better, with three or more metrics in mAP higher than the baseline, and AR also outperforming the baseline. In Faster R-CNN object detection, compared to the original noisy images, the DSPNet denoising algorithm shows improvements in mAP_50, as well as in the mAP and AR of large objects, but its overall performance is still relatively poor. The three denoising models, b2ub-g5-50, b2ub-p30, and nei-g5-50, perform better, with improvements in 7 metrics across all categories. When using SSD300 for object detection, except for the dspnet denoising model, which shows minimal improvement in the detection results, the remaining eight denoising models all show significant improvements in object detection performance. Seven or more metrics exhibit improvements compared to the baseline, with nei-g25 performing the best, as all ten evaluation metrics outperform the baseline. In YOLOX, multiple denoising models significantly improved the object detection results. Among them, b2ub-p30 increased the mAP from 0.675 to 0.695 compared to the original noisy images, with seven metrics higher than the baseline. Both b2ub-g5-50, b2ub-p5-50, and nei-g25 also showed seven metrics higher than the baseline. The dspnet denoising model performed poorly, with a decrease in all metrics.
Fig. 10 shows the results of mAP@0.5 and AR for object detection on the original MarinDebris dataset and the MarinDebris dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). As seen in the figure, the noisy dataset, serving as the baseline, exhibits high mAP@0.5 values across all algorithms, with little noticeable improvement from the denoising models. Overall, the b2ub-p5-50 model performs exceptionally well in both mAP@0.5 and AR, particularly in YOLOX and SSD-MobileNetV2, where it improves detection accuracy while maintaining a high recall rate. In contrast, dspnet and nei-g5-50 show slightly lower mAP@0.5 values, especially for YOLOX, where dspnet’s mAP@0.5 is 0.938, slightly lower than other models.
Table 18-21 show the results of detection using four object detection algorithms on the FDD underwater sonar dataset. When evaluating with the SSD-MobileNetV2 object detection algorithm, the original noisy data exhibited a relatively high overall mAP (0.779), especially reaching a value close to 1 at IoU=0.50. The nei-p5-50 denoising model performed the best, with improvements in all seven metrics. The second-best model was b2ub-g5-50, which showed improvements in mAP and AR for medium and large targets. On Faster R-CNN, the b2ub-g5-50 and b2ub-p30 denoising models performed better, with 8 metrics outperforming the baseline, while the other metrics were on par with the baseline. When using SSD300 for object detection, the b2ub-g5-50 denoising model shows the best improvement in object detection metrics, outperforming the baseline in 9 metrics, while the remaining AR_s is on par with the baseline metrics. On YOLOX, the b2ub-p30 and b2ub-p5-50 denoising models show the best improvement in object detection metrics. In all categories, seven metrics outperform the baseline, and the mAP of b2ub-p30 increased by 0.14.
Fig. 11 shows the mAP@0.5 and AR results of object detection on the original FDD dataset and the FDD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). The mAP@0.5 values (represented by the bar chart) remain high across all models, approaching 1. The b2ub-p5-50 model achieves the highest AR value (0.875) in YOLOX, indicating an improvement in recall performance with this model.
Tables 22-29 show the results of object detection using four detection algorithms on the UATD underwater sonar dataset. When evaluating with the SSD-MobileNetV2 object detection algorithm, the nei-p30 denoising model performs the best, with ten metrics surpassing the baseline, followed by b2ub-g5-50, which outperforms the baseline in nine metrics, while mAP_75 remains equal to the baseline. On Faster R-CNN, the nei-p30 denoising model again performs best, with all metrics exceeding the baseline. When using SSD300 for object detection, the b2ub-g5-50, b2ub-p30, b2ub-p5-50, nei-g5-50, and nei-p5-50 models all perform well, surpassing the baseline in seven metrics. However, they generally fall short of the baseline in metrics related to small objects. On YOLOX, b2ub-p5-50 and nei-p5-50 show good performance, with eight metrics exceeding the baseline. After denoising with the b2ub-g5-50 model, the mAP_l and AR_l detection metrics are lower than the baseline, showing poor detection performance for large objects. After denoising with the b2ub-p5-50 model, the mAP_s and AR_s detection metrics are lower than the baseline, showing poor detection performance for small objects.
Fig. 12 shows the mAP@0.5 and AR results of object detection on the original UATD dataset and the UATD dataset processed by nine denoising models using four object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, YOLOX). Among all the denoising models, the nei-p30 model stands out in both mAP@0.5 and AR metrics, achieving good performance across nearly all detection algorithms. Meanwhile, the b2ub-p5-50 and nei-p5-50 models also demonstrate stable performance across the different algorithms.
In this chapter, we conducted a comprehensive study on the impact of various deep learning-based denoising algorithms on underwater sonar image datasets for object detection tasks. Three denoising algorithms—Neighbor2Neighbor, Blind2Unblind, and DSPNet—were selected based on their strengths in handling different types of noise in sonar images. Neighbor2Neighbor effectively reduces noise by leveraging pixel pairs within noisy images, while Blind2Unblind employs an innovative transition from blind to non-blind denoising, allowing for detailed noise removal without sacrificing important image informati on. DSPNet is designed specifically for speckle noise reduction and improves image quality by using advanced techniques like feature pyramid networks.
We trained nine distinct denoising models based on different noise types and parameters, then evaluated their impact on object detection performance using multiple object detection algorithms (SSD-MobileNetV2, Faster R-CNN, SSD300, and YOLOX). The results, presented across several datasets, demonstrate that denoising algorithms can improve detection performance, especially for medium and large targets, while occasionally hindering the detection of small targets due to the loss of fine details.
In particular, models such as b2ub-g25 and nei-p30 showed notable improvements in mAP@0.5 and AR metrics, with some variations depending on the dataset and detection algorithm used. While some models excelled in certain scenarios (e.g., DSPNet with SSD-MobileNetV2 for medium targets), others like the b2ub-p5-50 model achieved the highest recall values in YOLOX and SSD-MobileNetV2. However, over-denoising was observed in some cases, which led to a reduction in performance due to the loss of crucial image details.
The findings suggest that while denoising algorithms can significantly improve detection metrics, a careful balance between noise reduction and detail preservation is crucial. Further research should focus on optimizing denoising models for specific target categories and sizes to better support downstream object detection tasks in underwater sonar image analysis.
V-D object detection on the underwater sonar datasets after multi-frame denoising
The selection of denoised datasets is a critical step in ensuring improved performance in subsequent object detection tasks. Given the varying characteristics of different object detection algorithms, it is essential to establish appropriate selection criteria tailored to each method. This allows for the identification of the most effective denoised datasets, which can be further optimized using image fusion techniques to enhance detection performance.
For the denoised SCTD datasets, we evaluate their performance based on key metrics, including mAP, mAP@0.5, mAP@0.75, and AR, and apply distinct selection criteria for different object detection algorithms:
YOLOX: Denoised datasets that outperform the original noisy dataset across all four metrics are selected. The datasets meeting this criterion include b2ub-g25, b2ub-p30, nei2nei-g25, nei2nei-g5-50, and nei2nei-p30, resulting in a total of five selected datasets.
SSD300: Denoised datasets that demonstrate improvements in at least two of the four metrics compared to the original dataset are selected, yielding five datasets.
Faster R-CNN: Only denoised datasets that exhibit improvements across all four metrics are selected, resulting in four datasets.
SSD_Mobilenet: Denoised datasets that show improvements in at least two of the four metrics are selected, with four datasets meeting this criterion.
By applying these criteria, we ensure that the selected denoised datasets provide substantial improvements in object detection performance, laying the foundation for further optimization through image fusion.
Fig. 13 presents the detection results of a sample from the SCTD dataset using the YOLOX algorithm, generated with MMDetection. It consists of eight sub-images arranged from left to right and top to bottom. The first image represents the ground truth, the second image shows the detection result on the original noisy dataset, followed by the results from different denoising methods: B2UB-G25, B2UB-P30, Nei2Nei-G25, Nei2Nei-G5-50, Nei2Nei-P30, and OPD. The confidence scores for object detection are displayed in each image, with OPD achieving the highest confidence among all methods.
| YOLOX | ||||
| Denoising | mAP | mAP_50 | mAP_75 | AR |
| noisy | 0.349 | 0.621 | 0.276 | 0.43 |
| b2ub-g25 | 0.384 | 0.68 | 0.411 | 0.475 |
| b2ub-p30 | 0.364 | 0.696 | 0.4 | 0.474 |
| nei-g25 | 0.423 | 0.682 | 0.485 | 0.482 |
| nei-g5-50 | 0.386 | 0.67 | 0.357 | 0.434 |
| nei-p30 | 0.381 | 0.674 | 0.402 | 0.445 |
| OPD-5 | 0.429 | 0.682 | 0.504 | 0.497 |
| SSD300 | ||||
| Denoising | mAP | mAP_50 | mAP_75 | AR |
| noisy | 0.461 | 0.756 | 0.61 | 0.543 |
| b2ubg25 | 0.46 | 0.813 | 0.47 | 0.552 |
| b2ub-g5-50 | 0.461 | 0.775 | 0.537 | 0.557 |
| b2ubp30 | 0.493 | 0.811 | 0.639 | 0.547 |
| neig25 | 0.442 | 0.782 | 0.48 | 0.536 |
| dspnet | 0.494 | 0.776 | 0.55 | 0.557 |
| OPD-5 | 0.426 | 0.763 | 0.462 | 0.51 |
| Faster R-CNN | ||||
| Denoising | mAP | mAP_50 | mAP_75 | AR |
| noisy | 0.297 | 0.554 | 0.222 | 0.408 |
| nei-g25 | 0.369 | 0.613 | 0.429 | 0.461 |
| nei-g5-50 | 0.302 | 0.557 | 0.305 | 0.416 |
| nei-p30 | 0.351 | 0.642 | 0.327 | 0.431 |
| nei-p5-50 | 0.327 | 0.589 | 0.325 | 0.414 |
| OPD-4 | 0.317 | 0.545 | 0.309 | 0.432 |
| SSD_Mobilenet | ||||
| Denoising | mAP | mAP_50 | mAP_75 | AR |
| noisy | 0.484 | 0.733 | 0.546 | 0.561 |
| b2ub-g5-50 | 0.461 | 0.775 | 0.537 | 0.557 |
| nei-g25 | 0.369 | 0.613 | 0.429 | 0.461 |
| nei-p30 | 0.351 | 0.642 | 0.327 | 0.431 |
| dspnet | 0.494 | 0.776 | 0.55 | 0.557 |
| OPD-4 | 0.489 | 0.772 | 0.525 | 0.573 |
The experimental results of the OPD algorithm on the SCTD dataset are shown in Table 1, where four object detection algorithms were applied to the multi-frame denoised SCTD dataset. The highest values in each column are marked in red, and the second-highest values are marked in blue.
OPD shows the most significant improvement when applied to the YOLOX object detection algorithm, with mAP of 0.429, mAP_75 of 0.504, and AR of 0.497. These three metrics are higher than those of the other common denoising methods, and all four metrics outperform the original noisy dataset. However, under SSD300 object detection, the OPD-denoised mAP, mAP_50, mAP_75, and AR are relatively low, particularly in mAP and mAP_75, where they are significantly lower than other denoising methods. In Faster R-CNN and SSD_Mobilenet, OPD performs somewhat modestly, but still, three of its metrics outperform those of the original noisy dataset.
Due to the structural characteristics of different object detection models and the way they process input data, the performance of the same denoising method varies across different models. Based on the above experiments, it was found that the OPD multi-frame denoising algorithm notably improves the object detection results of YOLOX, but the metrics decrease on SSD300. Compared to YOLOX, the SSD300 model is more focused on multi-scale feature maps and detection frameworks. While SSD300 can also extract features, its multi-scale processing mechanism makes it more susceptible to the effects of noise on different scale feature maps. Therefore, in this structure, OPD may lead to the loss of some useful details, especially in terms of detection precision at higher IoU thresholds. In other words, although OPD aims to remove noise, the denoising process may remove some details that are crucial for the multi-scale feature processing in SSD300.
This subsection investigates the performance of the SCTD dataset after multi-frame denoising under different object detection algorithms and establishes selection criteria for denoised datasets based on four key evaluation metrics. Experimental results indicate that the OPD method achieves the most significant improvement in the YOLOX object detection task, whereas certain metrics decline in the SSD300 task. Different detection models exhibit varying adaptability to denoised data, which is closely related to their feature extraction and multi-scale processing mechanisms. Therefore, when selecting a denoising method, it is crucial to consider the characteristics of the object detection model to ensure noise removal while preserving critical information, thereby optimizing detection performance.
VI Conclusion
This paper systematically conducts academic experiments and analysis on the application of various deep learning-based denoising algorithms in underwater sonar images to explore the impact of denoising on object detection performance. By applying different types of denoising treatments to underwater sonar images and processing the denoised images with four typical object detection algorithms, we find that deep learning denoising algorithms developed for optical images can, in some scenarios, improve object detection performance. However, the impact of different denoising algorithms on detection results varies significantly, and the noise types in underwater sonar images are in a complex mixed state.
To maximize object detection performance, this paper uses a set of images processed by different denoising algorithms as input to train a multi-frame denoising model, treating them as multiple shots of the same scene to further reduce noise. This approach has been extensively studied in the field of optical images and is explored for underwater sonar images in this paper. The experimental results validate that this method can effectively retain target information and remove noise more efficiently when applied to the YOLOX algorithm, thus improving the accuracy and robustness of object detection.
In conclusion, through comprehensive experimental analysis, this study provides an experimental evaluation of denoising algorithms in the field of underwater sonar image object detection, offering important references for future research. This paper analyzes the impact of three advanced deep learning-based denoising algorithms on object detection and innovatively applies outputs from multiple denoising algorithms as multi-frame images for further denoising. The experimental results provide meaningful insights for denoising and object detection of underwater sonar images.
Appendix A Appendix Section 1
| Denoising | SSD-MobileNetV2 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | ship | 0.552 | 0.782 | 0.693 | 0 | 0.507 | 0.582 | - | - | - | - |
| aircraft | 0.479 | 0.877 | 0.407 | nan | 0.652 | 0.445 | - | - | - | - | |
| human | 0.421 | 0.539 | 0.539 | 0.6 | 0.365 | 0.468 | - | - | - | - | |
| all | 0.484 | 0.733 | 0.546 | 0.3 | 0.508 | 0.498 | 0.561 | 0.3 | 0.531 | 0.64 | |
| dspnet | ship | 0.485 | 0.771 | 0.553 | 0 | 0.35 | 0.529 | - | - | - | - |
| aircraft | 0.541 | 0.855 | 0.699 | nan | 0.802 | 0.503 | - | - | - | - | |
| human | 0.434 | 0.58 | 0.405 | 0.4 | 0.579 | 0.328 | - | - | - | - | |
| all | 0.487 | 0.735 | 0.552 | 0.2 | 0.577 | 0.453 | 0.586 | 0.2 | 0.596 | 0.581 | |
| b2ub- g25 | ship | 0.521 | 0.78 | 0.63 | 0 | 0.465 | 0.56 | - | - | - | - |
| aircraft | 0.528 | 0.86 | 0.522 | nan | 0.653 | 0.494 | - | - | - | - | |
| human | 0.389 | 0.639 | 0.473 | 0.4 | 0.294 | 0.558 | - | - | - | - | |
| all | 0.479 | 0.76 | 0.542 | 0.2 | 0.471 | 0.537 | 0.548 | 0.2 | 0.471 | 0.537 | |
| b2ub- g5-50 | ship | 0.537 | 0.786 | 0.663 | 0 | 0.471 | 0.572 | - | - | - | - |
| aircraft | 0.487 | 0.888 | 0.437 | nan | 0.702 | 0.441 | - | - | - | - | |
| human | 0.43 | 0.573 | 0.401 | 0.5 | 0.437 | 0.417 | - | - | - | - | |
| all | 0.484 | 0.749 | 0.5 | 0.25 | 0.537 | 0.477 | 0.58 | 0.25 | 0.554 | 0.635 | |
| b2ub- p30 | ship | 0.504 | 0.754 | 0.589 | 0 | 0.528 | 0.533 | - | - | - | - |
| aircraft | 0.515 | 0.805 | 0.498 | nan | 0.702 | 0.485 | - | - | - | - | |
| human | 0.39 | 0.473 | 0.446 | 0.7 | 0.395 | 0.303 | - | - | - | - | |
| all | 0.47 | 0.677 | 0.511 | 0.35 | 0.542 | 0.44 | 0.53 | 0.35 | 0.563 | 0.492 | |
| b2ub- p5-50 | ship | 0.538 | 0.769 | 0.644 | 0 | 0.531 | 0.569 | - | - | - | - |
| aircraft | 0.484 | 0.857 | 0.51 | nan | 0.751 | 0.434 | - | - | - | - | |
| human | 0.343 | 0.47 | 0.337 | 0.5 | 0.351 | 0.303 | - | - | - | - | |
| all | 0.455 | 0.699 | 0.497 | 0.25 | 0.545 | 0.435 | 0.521 | 0.25 | 0.572 | 0.495 | |
| nei- g25 | ship | 0.508 | 0.766 | 0.577 | 0 | 0.491 | 0.534 | - | - | - | - |
| aircraft | 0.51 | 0.918 | 0.427 | nan | 0.603 | 0.495 | - | - | - | - | |
| human | 0.443 | 0.548 | 0.482 | 0.6 | 0.48 | 0.38 | - | - | - | - | |
| all | 0.487 | 0.744 | 0.495 | 0.3 | 0.525 | 0.47 | 0.596 | 0.3 | 0.544 | 0.669 | |
| nei- g5-50 | ship | 0.516 | 0.783 | 0.603 | 0 | 0.466 | 0.55 | - | - | - | - |
| aircraft | 0.368 | 0.739 | 0.347 | nan | 0.552 | 0.326 | - | - | - | - | |
| human | 0.35 | 0.487 | 0.446 | 0.8 | 0.314 | 0.303 | - | - | - | - | |
| all | 0.411 | 0.67 | 0.465 | 0.4 | 0.444 | 0.393 | 0.477 | 0.4 | 0.452 | 0.46 | |
| nei- p30 | ship | 0.528 | 0.806 | 0.666 | 0 | 0.513 | 0.554 | - | - | - | - |
| aircraft | 0.487 | 0.877 | 0.396 | nan | 0.652 | 0.463 | - | - | - | - | |
| human | 0.387 | 0.527 | 0.372 | 0.5 | 0.409 | 0.343 | - | - | - | - | |
| all | 0.468 | 0.737 | 0.478 | 0.25 | 0.525 | 0.454 | 0.57 | 0.25 | 0.547 | 0.632 | |
| nei- p5-50 | ship | 0.547 | 0.788 | 0.679 | 0 | 0.509 | 0.584 | - | - | - | - |
| aircraft | 0.484 | 0.866 | 0.505 | nan | 0.651 | 0.456 | - | - | - | - | |
| human | 0.389 | 0.534 | 0.377 | 0.4 | 0.401 | 0.414 | - | - | - | - | |
| all | 0.473 | 0.729 | 0.521 | 0.2 | 0.52 | 0.485 | 0.588 | 0.2 | 0.54 | 0.676 | |
| Denoising | Faster r-cnn | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | ship | 0.405 | 0.716 | 0.4 | 0 | 0.466 | 0.415 | - | - | - | - |
| aircraft | 0.23 | 0.453 | 0.192 | nan | 0.7 | 0.149 | - | - | - | - | |
| human | 0.257 | 0.492 | 0.076 | 0 | 0.203 | 0.495 | - | - | - | - | |
| all | 0.297 | 0.554 | 0.222 | 0 | 0.456 | 0.353 | 0.408 | 0 | 0.461 | 0.502 | |
| dspnet | ship | 0.379 | 0.712 | 0.339 | 0 | 0.403 | 0.392 | - | - | - | - |
| aircraft | 0.135 | 0.534 | 0.056 | nan | 0.431 | 0.103 | - | - | - | - | |
| human | 0.236 | 0.441 | 0.141 | 0.5 | 0.173 | 0.374 | - | - | - | - | |
| all | 0.25 | 0.563 | 0.179 | 0.25 | 0.336 | 0.29 | 0.419 | 0.25 | 0.41 | 0.511 | |
| b2ub- g25 | ship | 0.419 | 0.72 | 0.476 | 0 | 0.468 | 0.431 | - | - | - | - |
| aircraft | 0.209 | 0.407 | 0.218 | nan | 0.429 | 0.161 | - | - | - | - | |
| human | 0.303 | 0.505 | 0.228 | 0 | 0.236 | 0.59 | - | - | - | - | |
| all | 0.311 | 0.544 | 0.307 | 0 | 0.378 | 0.394 | 0.445 | 0 | 0.439 | 0.549 | |
| b2ub- g5-50 | ship | 0.427 | 0.783 | 0.501 | 0 | 0.418 | 0.448 | - | - | - | - |
| aircraft | 0.223 | 0.418 | 0.189 | nan | 0.559 | 0.183 | - | - | - | - | |
| human | 0.202 | 0.351 | 0.252 | 0 | 0.244 | 0.262 | - | - | - | - | |
| all | 0.284 | 0.518 | 0.314 | 0 | 0.407 | 0.298 | 0.444 | 0 | 0.513 | 0.521 | |
| b2ub- p30 | ship | 0.418 | 0.718 | 0.452 | 0 | 0.364 | 0.452 | - | - | - | - |
| aircraft | 0.208 | 0.396 | 0.169 | nan | 0.635 | 0.154 | - | - | - | - | |
| human | 0.324 | 0.518 | 0.424 | 0 | 0.244 | 0.617 | - | - | - | - | |
| all | 0.317 | 0.544 | 0.348 | 0 | 0.414 | 0.408 | 0.416 | 0 | 0.452 | 0.524 | |
| b2ub- p5-50 | ship | 0.414 | 0.708 | 0.469 | 0 | 0.37 | 0.439 | - | - | - | - |
| aircraft | 0.243 | 0.449 | 0.282 | nan | 0.65 | 0.167 | - | - | - | - | |
| human | 0.291 | 0.458 | 0.337 | 0 | 0.284 | 0.455 | - | - | - | - | |
| all | 0.316 | 0.538 | 0.362 | 0 | 0.435 | 0.354 | 0.41 | 0 | 0.432 | 0.481 | |
| nei- g25 | ship | 0.453 | 0.764 | 0.499 | 0 | 0.509 | 0.468 | - | - | - | - |
| aircraft | 0.281 | 0.519 | 0.343 | nan | 0.825 | 0.201 | - | - | - | - | |
| human | 0.374 | 0.554 | 0.446 | 0 | 0.284 | 0.7 | - | - | - | - | |
| all | 0.369 | 0.613 | 0.429 | 0 | 0.54 | 0.456 | 0.461 | 0 | 0.549 | 0.55 | |
| nei- g5-50 | ship | 0.38 | 0.725 | 0.341 | 0 | 0.417 | 0.397 | - | - | - | - |
| aircraft | 0.192 | 0.391 | 0.13 | nan | 0.751 | 0.11 | - | - | - | - | |
| human | 0.334 | 0.554 | 0.446 | 0 | 0.244 | 0.667 | - | - | - | - | |
| all | 0.302 | 0.557 | 0.305 | 0 | 0.471 | 0.391 | 0.416 | 0 | 0.53 | 0.498 | |
| nei- p30 | ship | 0.412 | 0.72 | 0.416 | 0 | 0.444 | 0.427 | - | - | - | - |
| aircraft | 0.287 | 0.653 | 0.316 | nan | 0.8 | 0.189 | - | - | - | - | |
| human | 0.355 | 0.554 | 0.25 | 0 | 0.203 | 0.789 | - | - | - | - | |
| all | 0.351 | 0.642 | 0.327 | 0 | 0.482 | 0.468 | 0.431 | 0 | 0.489 | 0.557 | |
| nei- p5-50 | ship | 0.36 | 0.748 | 0.258 | 0 | 0.353 | 0.38 | - | - | - | - |
| aircraft | 0.288 | 0.466 | 0.381 | nan | 0.584 | 0.239 | - | - | - | - | |
| human | 0.334 | 0.554 | 0.337 | 0 | 0.284 | 0.534 | - | - | - | - | |
| all | 0.414 | 0.589 | 0.325 | 0 | 0.407 | 0.384 | 0.414 | 0 | 0.471 | 0.473 | |
| Denoising | ssd300 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | ship | 0.498 | 0.775 | 0.604 | 0 | 0.495 | 0.522 | - | - | - | - |
| aircraft | 0.488 | 0.862 | 0.592 | nan | 0.403 | 0.5 | - | - | - | - | |
| human | 0.398 | 0.632 | 0.632 | 0.6 | 0.244 | 0.645 | - | - | - | - | |
| all | 0.461 | 0.756 | 0.61 | 0.3 | 0.38 | 0.556 | 0.543 | 0.3 | 0.391 | 0.669 | |
| dspnet | ship | 0.466 | 0.725 | 0.546 | 0 | 0.45 | 0.494 | - | - | - | - |
| aircraft | 0.549 | 0.905 | 0.55 | nan | 0.553 | 0.554 | - | - | - | - | |
| human | 0.468 | 0.697 | 0.554 | 0.5 | 0.317 | 0.8 | - | - | - | - | |
| all | 0.494 | 0.776 | 0.55 | 0.25 | 0.44 | 0.616 | 0.557 | 0.25 | 0.46 | 0.661 | |
| b2ub- g25 | ship | 0.495 | 0.756 | 0.575 | 0 | 0.483 | 0.519 | - | - | - | - |
| aircraft | 0.471 | 0.934 | 0.3 | nan | 0.553 | 0.466 | - | - | - | - | |
| human | 0.413 | 0.748 | 0.535 | 0.3 | 0.333 | 0.783 | - | - | - | - | |
| all | 0.46 | 0.813 | 0.47 | 0.15 | 0.456 | 0.589 | 0.552 | 0.15 | 0.481 | 0.664 | |
| b2ub- g5-50 | ship | 0.511 | 0.77 | 0.577 | 0 | 0.474 | 0.539 | - | - | - | - |
| aircraft | 0.504 | 0.873 | 0.558 | nan | 0.552 | 0.5 | - | - | - | - | |
| human | 0.368 | 0.68 | 0.477 | 0.3 | 0.261 | 0.766 | - | - | - | - | |
| all | 0.461 | 0.775 | 0.537 | 0.15 | 0.429 | 0.602 | 0.557 | 0.15 | 0.487 | 0.662 | |
| b2ub- p30 | ship | 0.483 | 0.723 | 0.56 | 0 | 0.524 | 0.502 | - | - | - | - |
| aircraft | 0.528 | 0.977 | 0.694 | nan | 0.403 | 0.55 | - | - | - | - | |
| human | 0.466 | 0.733 | 0.663 | 0.7 | 0.304 | 0.711 | - | - | - | - | |
| all | 0.493 | 0.811 | 0.639 | 0.35 | 0.41 | 0.587 | 0.547 | 0.35 | 0.428 | 0.648 | |
| b2ub- p5-50 | ship | 0.473 | 0.754 | 0.503 | 0 | 0.45 | 0.5 | - | - | - | - |
| aircraft | 0.497 | 0.868 | 0.334 | nan | 0.702 | 0.48 | - | - | - | - | |
| human | 0.381 | 0.694 | 0.168 | 0.5 | 0.215 | 0.766 | - | - | - | - | |
| all | 0.45 | 0.772 | 0.335 | 0.25 | 0.456 | 0.582 | 0.53 | 0.25 | 0.474 | 0.642 | |
| nei- g25 | ship | 0.514 | 0.784 | 0.6 | 0 | 0.511 | 0.541 | - | - | - | - |
| aircraft | 0.502 | 0.952 | 0.451 | nan | 0.503 | 0.505 | - | - | - | - | |
| human | 0.31 | 0.609 | 0.389 | 0.15 | 0.284 | 0.618 | - | - | - | - | |
| all | 0.442 | 0.782 | 0.48 | 0.075 | 0.433 | 0.555 | 0.536 | 0.1 | 0.438 | 0.656 | |
| nei- g5-50 | ship | 0.483 | 0.758 | 0.569 | 0 | 0.47 | 0.507 | - | - | - | - |
| aircraft | 0.399 | 0.81 | 0.38 | nan | 0.504 | 0.396 | - | - | - | - | |
| human | 0.398 | 0.667 | 0.408 | 0.4 | 0.302 | 0.8 | - | - | - | - | |
| all | 0.427 | 0.745 | 0.452 | 0.2 | 0.425 | 0.568 | 0.548 | 0.2 | 0.466 | 0.654 | |
| nei- p30 | ship | 0.497 | 0.751 | 0.51 | 0 | 0.48 | 0.528 | - | - | - | - |
| aircraft | 0.431 | 0.813 | 0.332 | nan | 0.403 | 0.435 | - | - | - | - | |
| human | 0.378 | 0.677 | 0.462 | 0.3 | 0.299 | 0.655 | - | - | - | - | |
| all | 0.435 | 0.747 | 0.435 | 0.15 | 0.394 | 0.54 | 0.529 | 0.15 | 0.413 | 0.632 | |
| nei- p5-50 | ship | 0.481 | 0.77 | 0.599 | 0 | 0.494 | 0.506 | - | - | - | - |
| aircraft | 0.492 | 0.901 | 0.434 | nan | 0.552 | 0.507 | - | - | - | - | |
| human | 0.347 | 0.663 | 0.168 | 0.4 | 0.203 | 0.711 | - | - | - | - | |
| all | 0.44 | 0.778 | 0.401 | 0.2 | 0.417 | 0.575 | 0.52 | 0.2 | 0.422 | 0.644 | |
| Denoising | YOLOX | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | ship | 0.362 | 0.636 | 0.331 | 0 | 0.407 | 0.37 | - | - | - | - |
| aircraft | 0.413 | 0.755 | 0.268 | nan | 0.55 | 0.377 | - | - | - | - | |
| human | 0.272 | 0.473 | 0.228 | 0.8 | 0.171 | 0.269 | - | - | - | - | |
| all | 0.349 | 0.621 | 0.276 | 0.4 | 0.376 | 0.339 | 0.43 | 0.4 | 0.404 | 0.431 | |
| dspnet | ship | 0.367 | 0.699 | 0.401 | 0 | 0.449 | 0.377 | - | - | - | - |
| aircraft | 0.361 | 0.704 | 0.267 | nan | 0.75 | 0.281 | - | - | - | - | |
| human | 0.178 | 0.368 | 0.119 | 0.4 | 0.139 | 0.297 | - | - | - | - | |
| all | 0.302 | 0.591 | 0.262 | 0.2 | 0.446 | 0.318 | 0.427 | 0.2 | 0.467 | 0.465 | |
| b2ub- g25 | ship | 0.444 | 0.728 | 0.531 | 0 | 0.464 | 0.465 | - | - | - | - |
| aircraft | 0.437 | 0.824 | 0.473 | nan | 0.75 | 0.37 | - | - | - | - | |
| human | 0.272 | 0.487 | 0.228 | 0.6 | 0.203 | 0.301 | - | - | - | - | |
| all | 0.384 | 0.68 | 0.411 | 0.3 | 0.472 | 0.379 | 0.475 | 0.3 | 0.489 | 0.502 | |
| b2ub- g5-50 | ship | 0.416 | 0.727 | 0.398 | 0 | 0.473 | 0.424 | - | - | - | - |
| aircraft | 0.409 | 0.69 | 0.325 | nan | 0.752 | 0.341 | - | - | - | - | |
| human | 0.258 | 0.446 | 0.282 | 0.3 | 0.284 | 0.269 | - | - | - | - | |
| all | 0.361 | 0.621 | 0.335 | 0.15 | 0.503 | 0.345 | 0.426 | 0.15 | 0.516 | 0.413 | |
| b2ub- p30 | ship | 0.373 | 0.704 | 0.408 | 0 | 0.365 | 0.393 | - | - | - | - |
| aircraft | 0.373 | 0.737 | 0.369 | nan | 0.701 | 0.313 | - | - | - | - | |
| human | 0.345 | 0.648 | 0.424 | 0.7 | 0.244 | 0.395 | - | - | - | - | |
| all | 0.364 | 0.696 | 0.4 | 0.35 | 0.436 | 0.367 | 0.474 | 0.35 | 0.447 | 0.51 | |
| b2ub- p5-50 | ship | 0.429 | 0.718 | 0.444 | 0 | 0.479 | 0.438 | - | - | - | - |
| aircraft | 0.408 | 0.821 | 0.254 | nan | 0.525 | 0.39 | - | - | - | - | |
| human | 0.189 | 0.361 | 0.119 | 0.1 | 0.249 | 0.303 | - | - | - | - | |
| all | 0.342 | 0.633 | 0.272 | 0.05 | 0.418 | 0.377 | 0.439 | 0.1 | 0.453 | 0.455 | |
| nei- g25 | ship | 0.467 | 0.728 | 0.584 | 0 | 0.565 | 0.48 | - | - | - | - |
| aircraft | 0.498 | 0.872 | 0.587 | nan | 0.65 | 0.472 | - | - | - | - | |
| human | 0.303 | 0.446 | 0.282 | 0.5 | 0.284 | 0.303 | - | - | - | - | |
| all | 0.423 | 0.682 | 0.485 | 0.25 | 0.5 | 0.418 | 0.482 | 0.25 | 0.516 | 0.478 | |
| nei- g5-50 | ship | 0.423 | 0.723 | 0.427 | 0 | 0.515 | 0.436 | - | - | - | - |
| aircraft | 0.523 | 0.907 | 0.526 | nan | 0.75 | 0.482 | - | - | - | - | |
| human | 0.211 | 0.38 | 0.119 | 0.4 | 0.162 | 0.303 | - | - | - | - | |
| all | 0.386 | 0.67 | 0.357 | 0.2 | 0.476 | 0.407 | 0.434 | 0.2 | 0.476 | 0.451 | |
| nei- p30 | ship | 0.423 | 0.683 | 0.573 | 0 | 0.507 | 0.436 | - | - | - | - |
| aircraft | 0.492 | 0.893 | 0.515 | nan | 0.65 | 0.459 | - | - | - | - | |
| human | 0.229 | 0.446 | 0.119 | 0.433 | 0.122 | 0.337 | - | - | - | - | |
| all | 0.381 | 0.674 | 0.402 | 0.217 | 0.427 | 0.411 | 0.445 | 0.25 | 0.429 | 0.482 | |
| nei- p5-50 | ship | 0.362 | 0.677 | 0.401 | 0 | 0.401 | 0.381 | - | - | - | - |
| aircraft | 0.465 | 0.888 | 0.396 | nan | 0.703 | 0.434 | - | - | - | - | |
| human | 0.211 | 0.446 | 0.212 | 0.3 | 0.223 | 0.202 | - | - | - | - | |
| all | 0.346 | 0.67 | 0.336 | 0.15 | 0.442 | 0.339 | 0.423 | 0.15 | 0.486 | 0.409 | |
| Denoising | ssd_mobilenetv2 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | object/all | 0.186 | 0.392 | 0.158 | 0.004 | 0.024 | 0.296 | 0.297 | 0.047 | 0.158 | 0.414 |
| dspnet | object/all | 0.177 | 0.381 | 0.152 | 0.014 | 0.027 | 0.295 | 0.293 | 0.058 | 0.16 | 0.405 |
| b2ub-g25 | object/all | 0.176 | 0.389 | 0.163 | 0.021 | 0.021 | 0.291 | 0.296 | 0.062 | 0.161 | 0.408 |
| b2ub-g5-50 | object/all | 0.186 | 0.39 | 0.163 | 0.011 | 0.019 | 0.298 | 0.294 | 0.077 | 0.145 | 0.411 |
| b2ub-p30 | object/all | 0.171 | 0.364 | 0.15 | 0.005 | 0.021 | 0.301 | 0.296 | 0.046 | 0.158 | 0.413 |
| b2ub-p5-50 | object/all | 0.177 | 0.388 | 0.145 | 0.017 | 0.027 | 0.3 | 0.302 | 0.067 | 0.172 | 0.411 |
| nei-g25 | object/all | 0.191 | 0.396 | 0.17 | 0.006 | 0.015 | 0.308 | 0.299 | 0.061 | 0.159 | 0.415 |
| nei-g5-50 | object/all | 0.184 | 0.39 | 0.162 | 0.014 | 0.014 | 0.305 | 0.299 | 0.064 | 0.15 | 0.418 |
| nei-p30 | object/all | 0.18 | 0.392 | 0.152 | 0.016 | 0.023 | 0.304 | 0.302 | 0.07 | 0.159 | 0.418 |
| nei-p5-50 | object/all | 0.175 | 0.382 | 0.141 | 0.026 | 0.026 | 0.293 | 0.298 | 0.073 | 0.166 | 0.407 |
| Denoising | Faster r-cnn | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | object/all | 0.271 | 0.562 | 0.243 | 0.155 | 0.104 | 0.368 | 0.365 | 0.222 | 0.264 | 0.443 |
| dspnet | object/all | 0.262 | 0.546 | 0.226 | 0.11 | 0.1 | 0.365 | 0.353 | 0.221 | 0.232 | 0.44 |
| b2ub-g25 | object/all | 0.248 | 0.523 | 0.214 | 0.078 | 0.085 | 0.353 | 0.354 | 0.141 | 0.242 | 0.449 |
| b2ub-g5-50 | object/all | 0.264 | 0.553 | 0.232 | 0.127 | 0.099 | 0.369 | 0.362 | 0.225 | 0.236 | 0.452 |
| b2ub-p30 | object/all | 0.271 | 0.572 | 0.235 | 0.146 | 0.111 | 0.365 | 0.365 | 0.251 | 0.26 | 0.439 |
| b2ub-p5-50 | object/all | 0.269 | 0.567 | 0.226 | 0.121 | 0.107 | 0.375 | 0.365 | 0.213 | 0.246 | 0.454 |
| nei-g25 | object/all | 0.266 | 0.55 | 0.23 | 0.105 | 0.104 | 0.371 | 0.354 | 0.183 | 0.232 | 0.448 |
| nei-g5-50 | object/all | 0.261 | 0.548 | 0.219 | 0.119 | 0.094 | 0.37 | 0.363 | 0.223 | 0.225 | 0.459 |
| nei-p30 | object/all | 0.271 | 0.565 | 0.235 | 0.147 | 0.103 | 0.372 | 0.369 | 0.248 | 0.247 | 0.453 |
| nei-p5-50 | object/all | 0.26 | 0.56 | 0.23 | 0.125 | 0.101 | 0.365 | 0.35 | 0.204 | 0.229 | 0.439 |
| Denoising | ssd300 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | object/all | 0.217 | 0.491 | 0.178 | 0.071 | 0.059 | 0.312 | 0.352 | 0.18 | 0.264 | 0.427 |
| dspnet | object/all | 0.213 | 0.468 | 0.188 | 0.065 | 0.055 | 0.309 | 0.345 | 0.163 | 0.272 | 0.415 |
| b2ub-g25 | object/all | 0.215 | 0.485 | 0.17 | 0.06 | 0.063 | 0.308 | 0.344 | 0.14 | 0.259 | 0.425 |
| b2ub-g5-50 | object/all | 0.219 | 0.496 | 0.177 | 0.066 | 0.066 | 0.313 | 0.359 | 0.196 | 0.269 | 0.434 |
| b2ub-p30 | object/all | 0.218 | 0.502 | 0.159 | 0.069 | 0.063 | 0.313 | 0.355 | 0.19 | 0.278 | 0.425 |
| b2ub-p5-50 | object/all | 0.137 | 0.346 | 0.095 | 0.05 | 0.042 | 0.201 | 0.291 | 0.136 | 0.242 | 0.343 |
| nei-g25 | object/all | 0.216 | 0.502 | 0.179 | 0.078 | 0.062 | 0.311 | 0.349 | 0.192 | 0.268 | 0.42 |
| nei-g5-50 | object/all | 0.218 | 0.488 | 0.173 | 0.066 | 0.062 | 0.314 | 0.35 | 0.15 | 0.274 | 0.425 |
| nei-p30 | object/all | 0.223 | 0.518 | 0.18 | 0.077 | 0.071 | 0.314 | 0.36 | 0.196 | 0.283 | 0.428 |
| nei-p5-50 | object/all | 0.219 | 0.492 | 0.178 | 0.081 | 0.056 | 0.318 | 0.355 | 0.23 | 0.273 | 0.42 |
| Denoising | YOLOX | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | object/all | 0.278 | 0.582 | 0.231 | 0.091 | 0.114 | 0.381 | 0.376 | 0.158 | 0.285 | 0.463 |
| dspnet | object/all | 0.264 | 0.551 | 0.232 | 0.102 | 0.091 | 0.367 | 0.354 | 0.151 | 0.257 | 0.441 |
| b2ub-g25 | object/all | 0.281 | 0.584 | 0.242 | 0.129 | 0.107 | 0.383 | 0.383 | 0.2 | 0.288 | 0.466 |
| b2ub-g5-50 | object/all | 0.274 | 0.579 | 0.236 | 0.104 | 0.116 | 0.375 | 0.369 | 0.152 | 0.273 | 0.457 |
| b2ub-p30 | object/all | 0.275 | 0.585 | 0.226 | 0.097 | 0.107 | 0.381 | 0.367 | 0.15 | 0.264 | 0.459 |
| b2ub-p5-50 | object/all | 0.275 | 0.586 | 0.233 | 0.098 | 0.12 | 0.375 | 0.371 | 0.142 | 0.286 | 0.456 |
| nei-g25 | object/all | 0.27 | 0.564 | 0.223 | 0.103 | 0.095 | 0.372 | 0.369 | 0.152 | 0.283 | 0.452 |
| nei-g5-50 | object/all | 0.272 | 0.567 | 0.236 | 0.112 | 0.098 | 0.372 | 0.37 | 0.166 | 0.286 | 0.45 |
| nei-p30 | object/all | 0.272 | 0.589 | 0.22 | 0.111 | 0.112 | 0.369 | 0.373 | 0.165 | 0.3 | 0.448 |
| nei-p5-50 | object/all | 0.268 | 0.57 | 0.223 | 0.098 | 0.1 | 0.373 | 0.366 | 0.138 | 0.27 | 0.457 |
| Denoising | ssd_mobilenetv2 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | Bottle | 0.739 | 0.981 | 0.948 | 0.2 | 0.735 | nan | - | - | - | - |
| Can | 0.34 | 0.652 | 0.314 | 0.348 | 0.352 | nan | - | - | - | - | |
| Chain | 0.687 | 0.949 | 0.834 | 0.217 | 0.457 | 0.738 | - | - | - | - | |
| Drink-carton | 0.567 | 0.887 | 0.699 | 0.595 | 0.537 | nan | - | - | - | - | |
| Hook | 0.727 | 0.956 | 0.956 | nan | 0.727 | nan | - | - | - | - | |
| Propeller | 0.708 | 0.989 | 0.882 | 0.5 | 0.715 | 0.8 | - | - | - | - | |
| Shampoo-bottle | 0.71 | 0.968 | 0.968 | nan | 0.708 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.788 | 1 | 1 | nan | 0.788 | nan | - | - | - | - | |
| Tire | 0.805 | 0.99 | 0.955 | 0.8 | 0.81 | 0.3 | - | - | - | - | |
| Valve | 0.725 | 0.927 | 0.901 | 0.383 | 0.751 | 0 | - | - | - | - | |
| Wall | 0.713 | 0.916 | 0.841 | 0.304 | 0.7 | 0.84 | - | - | - | - | |
| all | 0.683 | 0.929 | 0.845 | 0.418 | 0.662 | 0.58 | 0.743 | 0.571 | 0.729 | 0.596 | |
| dspnet | Bottle | 0.749 | 0.979 | 0.917 | 0.3 | 0.746 | nan | - | - | - | - |
| Can | 0.331 | 0.565 | 0.358 | 0.407 | 0.339 | nan | - | - | - | - | |
| Chain | 0.691 | 0.946 | 0.861 | 0.6 | 0.478 | 0.732 | - | - | - | - | |
| Drink-carton | 0.555 | 0.955 | 0.556 | 0.544 | 0.588 | nan | - | - | - | - | |
| Hook | 0.693 | 0.947 | 0.834 | nan | 0.693 | nan | - | - | - | - | |
| Propeller | 0.633 | 0.98 | 0.747 | 0.4 | 0.642 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.681 | 0.957 | 0.957 | nan | 0.681 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.786 | 1 | 1 | nan | 0.786 | nan | - | - | - | - | |
| Tire | 0.797 | 0.99 | 0.936 | 0.6 | 0.803 | 0.5 | - | - | - | - | |
| Valve | 0.701 | 0.917 | 0.847 | 0 | 0.733 | 0 | - | - | - | - | |
| Wall | 0.715 | 0.933 | 0.805 | 0.365 | 0.693 | 0.832 | - | - | - | - | |
| all | 0.667 | 0.924 | 0.802 | 0.402 | 0.653 | 0.561 | 0.729 | 0.469 | 0.724 | 0.575 | |
| b2ub- g25 | Bottle | 0.745 | 0.981 | 0.952 | 0.2 | 0.74 | nan | - | - | - | - |
| Can | 0.336 | 0.617 | 0.298 | 0.335 | 0.356 | nan | - | - | - | - | |
| Chain | 0.688 | 0.952 | 0.847 | 0.317 | 0.425 | 0.736 | - | - | - | - | |
| Drink-carton | 0.572 | 0.951 | 0.599 | 0.588 | 0.565 | nan | - | - | - | - | |
| Hook | 0.721 | 0.95 | 0.929 | nan | 0.721 | nan | - | - | - | - | |
| Propeller | 0.678 | 0.984 | 0.83 | 0.5 | 0.684 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.724 | 0.951 | 0.951 | nan | 0.716 | 0.9 | - | - | - | - | |
| Standing-bottle | 0.782 | 1 | 1 | nan | 0.782 | nan | - | - | - | - | |
| Tire | 0.803 | 0.99 | 0.944 | 0.6 | 0.81 | 0.3 | - | - | - | - | |
| Valve | 0.691 | 0.919 | 0.873 | 0.2 | 0.729 | 0 | - | - | - | - | |
| Wall | 0.699 | 0.919 | 0.813 | 0.28 | 0.688 | 0.827 | - | - | - | - | |
| all | 0.676 | 0.929 | 0.822 | 0.377 | 0.656 | 0.577 | 0.74 | 0.475 | 0.73 | 0.593 | |
| b2ub- g5-50 | Bottle | 0.746 | 0.976 | 0.943 | 0.2 | 0.742 | nan | - | - | - | - |
| Can | 0.346 | 0.629 | 0.39 | 0.384 | 0.355 | nan | - | - | - | - | |
| Chain | 0.693 | 0.95 | 0.837 | 0.65 | 0.489 | 0.728 | - | - | - | - | |
| Drink-carton | 0.551 | 0.93 | 0.664 | 0.585 | 0.516 | nan | - | - | - | - | |
| Hook | 0.727 | 0.944 | 0.944 | nan | 0.727 | nan | - | - | - | - | |
| Propeller | 0.68 | 0.985 | 0.836 | 0.5 | 0.685 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.722 | 0.968 | 0.968 | nan | 0.725 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.798 | 1 | 0.871 | nan | 0.798 | nan | - | - | - | - | |
| Tire | 0.815 | 0.989 | 0.958 | 0.8 | 0.823 | 0 | - | - | - | - | |
| Valve | 0.7 | 0.911 | 0.886 | 0 | 0.735 | 0 | - | - | - | - | |
| Wall | 0.715 | 0.93 | 0.828 | 0.304 | 0.701 | 0.852 | - | - | - | - | |
| all | 0.681 | 0.928 | 0.83 | 0.428 | 0.663 | 0.497 | 0.744 | 0.505 | 0.737 | 0.511 | |
| b2ub- p30 | Bottle | 0.749 | 0.984 | 0.973 | 0.35 | 0.743 | nan | - | - | - | - |
| Can | 0.371 | 0.649 | 0.357 | 0.375 | 0.388 | nan | - | - | - | - | |
| Chain | 0.683 | 0.95 | 0.825 | 0.601 | 0.44 | 0.723 | - | - | - | - | |
| Drink-carton | 0.547 | 0.931 | 0.558 | 0.527 | 0.59 | nan | - | - | - | - | |
| Hook | 0.739 | 0.951 | 0.951 | nan | 0.739 | nan | - | - | - | - | |
| Propeller | 0.689 | 0.991 | 0.895 | 0.7 | 0.692 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.738 | 0.958 | 0.958 | nan | 0.735 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.791 | 1 | 1 | nan | 0.791 | nan | - | - | - | - | |
| Tire | 0.804 | 0.99 | 0.947 | 0.6 | 0.812 | 0.2 | - | - | - | - | |
| Valve | 0.687 | 0.906 | 0.877 | 0 | 0.725 | 0 | - | - | - | - | |
| Wall | 0.723 | 0.926 | 0.852 | 0.325 | 0.716 | 0.835 | - | - | - | - | |
| all | 0.684 | 0.93 | 0.836 | 0.435 | 0.67 | 0.543 | 0.744 | 0.503 | 0.741 | 0.559 | |
| b2ub- p5-50 | Bottle | 0.749 | 0.986 | 0.965 | 0.25 | 0.745 | nan | - | - | - | - |
| Can | 0.492 | 0.838 | 0.501 | 0.528 | 0.486 | nan | - | - | - | - | |
| Chain | 0.686 | 0.963 | 0.84 | 0.734 | 0.554 | 0.718 | - | - | - | - | |
| Drink-carton | 0.641 | 0.978 | 0.832 | 0.663 | 0.619 | nan | - | - | - | - | |
| Hook | 0.695 | 0.954 | 0.935 | nan | 0.695 | nan | - | - | - | - | |
| Propeller | 0.663 | 0.974 | 0.773 | 0.6 | 0.667 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.702 | 0.957 | 0.843 | nan | 0.724 | 0.4 | - | - | - | - | |
| Standing-bottle | 0.725 | 1 | 0.871 | nan | 0.725 | nan | - | - | - | - | |
| Tire | 0.809 | 0.97 | 0.937 | 0.3 | 0.822 | 0 | - | - | - | - | |
| Valve | 0.689 | 0.918 | 0.821 | 0 | 0.72 | 0 | - | - | - | - | |
| Wall | 0.717 | 0.936 | 0.843 | 0.371 | 0.719 | 0.794 | - | - | - | - | |
| all | 0.688 | 0.952 | 0.833 | 0.431 | 0.679 | 0.435 | 0.761 | 0.515 | 0.753 | 0.461 | |
| nei- g25 | Bottle | 0.742 | 0.981 | 0.961 | 0.3 | 0.737 | nan | - | - | - | - |
| Can | 0.332 | 0.64 | 0.298 | 0.349 | 0.347 | nan | - | - | - | - | |
| Chain | 0.689 | 0.924 | 0.825 | 0.751 | 0.486 | 0.725 | - | - | - | - | |
| Drink-carton | 0.568 | 0.984 | 0.558 | 0.578 | 0.568 | nan | - | - | - | - | |
| Hook | 0.725 | 0.954 | 0.929 | nan | 0.725 | nan | - | - | - | - | |
| Propeller | 0.674 | 0.984 | 0.805 | 0.6 | 0.678 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.725 | 0.971 | 0.971 | nan | 0.724 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.72 | 1 | 1 | nan | 0.72 | nan | - | - | - | - | |
| Tire | 0.806 | 0.99 | 0.969 | 0.5 | 0.811 | 0.6 | - | - | - | - | |
| Valve | 0.713 | 0.925 | 0.881 | 0.25 | 0.742 | 0 | - | - | - | - | |
| Wall | 0.725 | 0.928 | 0.837 | 0.336 | 0.714 | 0.843 | - | - | - | - | |
| all | 0.674 | 0.935 | 0.821 | 0.458 | 0.659 | 0.611 | 0.739 | 0.557 | 0.727 | 0.629 | |
| nei- g5-50 | Bottle | 0.739 | 0.98 | 0.921 | 0.167 | 0.737 | nan | - | - | - | - |
| Can | 0.32 | 0.648 | 0.304 | 0.313 | 0.345 | nan | - | - | - | - | |
| Chain | 0.696 | 0.952 | 0.797 | 0.213 | 0.511 | 0.737 | - | - | - | - | |
| Drink-carton | 0.567 | 0.965 | 0.685 | 0.569 | 0.569 | nan | - | - | - | - | |
| Hook | 0.723 | 0.954 | 0.954 | nan | 0.723 | nan | - | - | - | - | |
| Propeller | 0.692 | 0.99 | 0.853 | 0.3 | 0.705 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.721 | 0.958 | 0.894 | nan | 0.713 | 0.9 | - | - | - | - | |
| Standing-bottle | 0.752 | 1 | 1 | nan | 0.752 | nan | - | - | - | - | |
| Tire | 0.799 | 0.98 | 0.929 | 0.5 | 0.81 | 0 | - | - | - | - | |
| Valve | 0.68 | 0.911 | 0.814 | 0 | 0.713 | 0 | - | - | - | - | |
| Wall | 0.714 | 0.92 | 0.828 | 0.353 | 0.695 | 0.839 | - | - | - | - | |
| all | 0.673 | 0.932 | 0.816 | 0.302 | 0.661 | 0.513 | 0.734 | 0.396 | 0.732 | 0.528 | |
| nei- p30 | Bottle | 0.751 | 0.981 | 0.963 | 0.167 | 0.749 | nan | - | - | - | - |
| Can | 0.354 | 0.624 | 0.357 | 0.344 | 0.379 | nan | - | - | - | - | |
| Chain | 0.684 | 0.95 | 0.8 | 0.251 | 0.455 | 0.733 | - | - | - | - | |
| Drink-carton | 0.578 | 0.958 | 0.65 | 0.578 | 0.585 | nan | - | - | - | - | |
| Hook | 0.701 | 0.944 | 0.914 | nan | 0.701 | nan | - | - | - | - | |
| Propeller | 0.632 | 0.964 | 0.77 | 0.6 | 0.638 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.735 | 0.964 | 0.964 | nan | 0.729 | 0.9 | - | - | - | - | |
| Standing-bottle | 0.736 | 1 | 1 | nan | 0.736 | nan | - | - | - | - | |
| Tire | 0.808 | 0.99 | 0.956 | 0.6 | 0.813 | 0.5 | - | - | - | - | |
| Valve | 0.72 | 0.944 | 0.901 | 0.2 | 0.751 | 0 | - | - | - | - | |
| Wall | 0.703 | 0.925 | 0.817 | 0.284 | 0.691 | 0.84 | - | - | - | - | |
| all | 0.673 | 0.931 | 0.826 | 0.378 | 0.657 | 0.595 | 0.738 | 0.465 | 0.731 | 0.614 | |
| nei- p5-50 | Bottle | 0.755 | 0.982 | 0.961 | 0.5 | 0.752 | nan | - | - | - | - |
| Can | 0.343 | 0.644 | 0.305 | 0.333 | 0.365 | nan | - | - | - | - | |
| Chain | 0.688 | 0.968 | 0.804 | 0.501 | 0.431 | 0.734 | - | - | - | - | |
| Drink-carton | 0.577 | 0.953 | 0.648 | 0.574 | 0.594 | nan | - | - | - | - | |
| Hook | 0.736 | 0.948 | 0.929 | nan | 0.736 | nan | - | - | - | - | |
| Propeller | 0.681 | 0.983 | 0.834 | 0.6 | 0.686 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.736 | 0.958 | 0.958 | nan | 0.734 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.803 | 1 | 1 | nan | 0.803 | nan | - | - | - | - | |
| Tire | 0.803 | 0.98 | 0.938 | 0.5 | 0.813 | 0 | - | - | - | - | |
| Valve | 0.681 | 0.914 | 0.887 | 0 | 0.712 | 0 | - | - | - | - | |
| Wall | 0.717 | 0.93 | 0.848 | 0.304 | 0.718 | 0.83 | - | - | - | - | |
| all | 0.684 | 0.933 | 0.828 | 0.414 | 0.668 | 0.494 | 0.745 | 0.437 | 0.74 | 0.512 |
| Denoising | Faster r-cnn | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | Bottle | 0.766 | 0.985 | 0.974 | 0.7 | 0.761 | nan | - | - | - | - |
| Can | 0.493 | 0.821 | 0.509 | 0.514 | 0.489 | nan | - | - | - | - | |
| Chain | 0.676 | 0.963 | 0.813 | 0.701 | 0.524 | 0.704 | - | - | - | - | |
| Drink-carton | 0.65 | 0.995 | 0.827 | 0.643 | 0.664 | nan | - | - | - | - | |
| Hook | 0.655 | 0.927 | 0.84 | nan | 0.655 | nan | - | - | - | - | |
| Propeller | 0.652 | 0.945 | 0.745 | 0.5 | 0.655 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.724 | 0.964 | 0.894 | nan | 0.729 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.78 | 1 | 1 | nan | 0.78 | nan | - | - | - | - | |
| Tire | 0.812 | 0.989 | 0.945 | 0.1 | 0.824 | 0 | - | - | - | - | |
| Valve | 0.692 | 0.917 | 0.86 | 0 | 0.722 | 0 | - | - | - | - | |
| Wall | 0.716 | 0.935 | 0.85 | 0.413 | 0.723 | 0.778 | - | - | - | - | |
| all | 0.692 | 0.949 | 0.842 | 0.446 | 0.684 | 0.497 | 0.763 | 0.502 | 0.754 | 0.525 | |
| dspnet | Bottle | 0.746 | 0.985 | 0.933 | 0.7 | 0.742 | nan | - | - | - | - |
| Can | 0.428 | 0.808 | 0.447 | 0.417 | 0.444 | nan | - | - | - | - | |
| Chain | 0.667 | 0.964 | 0.833 | 0.635 | 0.462 | 0.709 | - | - | - | - | |
| Drink-carton | 0.632 | 0.993 | 0.782 | 0.626 | 0.658 | nan | - | - | - | - | |
| Hook | 0.675 | 0.934 | 0.818 | nan | 0.675 | nan | - | - | - | - | |
| Propeller | 0.595 | 0.966 | 0.678 | 0.3 | 0.603 | 0.5 | - | - | - | - | |
| Shampoo-bottle | 0.715 | 0.969 | 0.909 | nan | 0.718 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.73 | 1 | 0.901 | nan | 0.73 | nan | - | - | - | - | |
| Tire | 0.792 | 0.999 | 0.94 | 0.1 | 0.803 | 0.3 | - | - | - | - | |
| Valve | 0.676 | 0.917 | 0.806 | 0 | 0.705 | 0 | - | - | - | - | |
| Wall | 0.705 | 0.923 | 0.827 | 0.371 | 0.711 | 0.772 | - | - | - | - | |
| all | 0.669 | 0.951 | 0.807 | 0.394 | 0.659 | 0.513 | 0.74 | 0.487 | 0.731 | 0.536 | |
| b2ub- g25 | Bottle | 0.759 | 0.986 | 0.963 | 0.3 | 0.754 | nan | - | - | - | - |
| Can | 0.499 | 0.827 | 0.493 | 0.564 | 0.478 | nan | - | - | - | - | |
| Chain | 0.678 | 0.969 | 0.786 | 0.75 | 0.499 | 0.715 | - | - | - | - | |
| Drink-carton | 0.66 | 0.989 | 0.754 | 0.651 | 0.688 | nan | - | - | - | - | |
| Hook | 0.715 | 0.939 | 0.914 | nan | 0.715 | nan | - | - | - | - | |
| Propeller | 0.659 | 0.99 | 0.822 | 0.6 | 0.66 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.721 | 0.98 | 0.919 | nan | 0.73 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.712 | 1 | 0.811 | nan | 0.712 | nan | - | - | - | - | |
| Tire | 0.813 | 0.96 | 0.942 | 0 | 0.829 | 0 | - | - | - | - | |
| Valve | 0.684 | 0.918 | 0.879 | 0 | 0.713 | 0 | - | - | - | - | |
| Wall | 0.721 | 0.94 | 0.829 | 0.361 | 0.735 | 0.788 | - | - | - | - | |
| all | 0.693 | 0.954 | 0.828 | 0.403 | 0.683 | 0.467 | 0.762 | 0.48 | 0.754 | 0.492 | |
| b2ub- g5-50 | Bottle | 0.765 | 0.986 | 0.975 | 0.233 | 0.76 | nan | - | - | - | - |
| Can | 0.512 | 0.844 | 0.508 | 0.552 | 0.503 | nan | - | - | - | - | |
| Chain | 0.689 | 0.965 | 0.836 | 0.75 | 0.562 | 0.716 | - | - | - | - | |
| Drink-carton | 0.643 | 0.976 | 0.808 | 0.662 | 0.623 | nan | - | - | - | - | |
| Hook | 0.709 | 0.948 | 0.897 | nan | 0.709 | nan | - | - | - | - | |
| Propeller | 0.676 | 0.975 | 0.874 | 0.6 | 0.681 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.704 | 0.978 | 0.882 | nan | 0.72 | 0.4 | - | - | - | - | |
| Standing-bottle | 0.818 | 1 | 1 | nan | 0.818 | nan | - | - | - | - | |
| Tire | 0.824 | 0.989 | 0.938 | 0.3 | 0.836 | 0 | - | - | - | - | |
| Valve | 0.696 | 0.92 | 0.865 | 0 | 0.726 | 0 | - | - | - | - | |
| Wall | 0.716 | 0.914 | 0.847 | 0.324 | 0.73 | 0.784 | - | - | - | - | |
| all | 0.705 | 0.954 | 0.857 | 0.428 | 0.697 | 0.433 | 0.768 | 0.531 | 0.758 | 0.494 | |
| b2ub- p30 | Bottle | 0.766 | 0.986 | 0.965 | 0.5 | 0.762 | nan | - | - | - | - |
| Can | 0.489 | 0.829 | 0.517 | 0.534 | 0.481 | nan | - | - | - | - | |
| Chain | 0.686 | 0.97 | 0.846 | 0.75 | 0.526 | 0.718 | - | - | - | - | |
| Drink-carton | 0.639 | 0.977 | 0.808 | 0.657 | 0.636 | nan | - | - | - | - | |
| Hook | 0.706 | 0.938 | 0.88 | nan | 0.706 | nan | - | - | - | - | |
| Propeller | 0.669 | 0.977 | 0.78 | 0.6 | 0.671 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.711 | 0.957 | 0.896 | nan | 0.713 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.711 | 1 | 0.871 | nan | 0.771 | nan | - | - | - | - | |
| Tire | 0.807 | 0.979 | 0.939 | 0.2 | 0.818 | 0 | - | - | - | - | |
| Valve | 0.691 | 0.92 | 0.863 | 0 | 0.721 | 0 | - | - | - | - | |
| Wall | 0.717 | 0.934 | 0.821 | 0.36 | 0.725 | 0.78 | - | - | - | - | |
| all | 0.696 | 0.952 | 0.835 | 0.45 | 0.685 | 0.5 | 0.764 | 0.52 | 0.752 | 0.526 | |
| b2ub- p5-50 | Bottle | 0.749 | 0.986 | 0.965 | 0.25 | 0.745 | nan | - | - | - | - |
| Can | 0.492 | 0.838 | 0.501 | 0.528 | 0.486 | nan | - | - | - | - | |
| Chain | 0.686 | 0.963 | 0.84 | 0.734 | 0.554 | 0.718 | - | - | - | - | |
| Drink-carton | 0.641 | 0.978 | 0.832 | 0.663 | 0.619 | nan | - | - | - | - | |
| Hook | 0.695 | 0.954 | 0.935 | nan | 0.695 | nan | - | - | - | - | |
| Propeller | 0.663 | 0.974 | 0.773 | 0.6 | 0.667 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.702 | 0.957 | 0.843 | nan | 0.724 | 0.4 | - | - | - | - | |
| Standing-bottle | 0.725 | 1 | 0.871 | nan | 0.725 | nan | - | - | - | - | |
| Tire | 0.809 | 0.97 | 0.937 | 0.3 | 0.822 | 0 | - | - | - | - | |
| Valve | 0.689 | 0.918 | 0.821 | 0 | 0.72 | 0 | - | - | - | - | |
| Wall | 0.717 | 0.936 | 0.843 | 0.371 | 0.719 | 0.794 | - | - | - | - | |
| all | 0.688 | 0.952 | 0.833 | 0.431 | 0.679 | 0.435 | 0.761 | 0.515 | 0.753 | 0.461 | |
| nei- g25 | Bottle | 0.764 | 0.988 | 0.978 | 0.7 | 0.757 | nan | - | - | - | - |
| Can | 0.492 | 0.847 | 0.499 | 0.524 | 0.498 | nan | - | - | - | - | |
| Chain | 0.672 | 0.965 | 0.812 | 0.617 | 0.522 | 0.707 | - | - | - | - | |
| Drink-carton | 0.637 | 0.996 | 0.84 | 0.653 | 0.623 | nan | - | - | - | - | |
| Hook | 0.676 | 0.931 | 0.905 | nan | 0.676 | nan | - | - | - | - | |
| Propeller | 0.654 | 0.979 | 0.822 | 0.25 | 0.662 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.697 | 0.967 | 0.859 | nan | 0.721 | 0.267 | - | - | - | - | |
| Standing-bottle | 0.777 | 1 | 1 | nan | 0.777 | nan | - | - | - | - | |
| Tire | 0.802 | 0.979 | 0.949 | 0.1 | 0.816 | 0 | - | - | - | - | |
| Valve | 0.674 | 0.92 | 0.824 | 0 | 0.703 | 0 | - | - | - | - | |
| Wall | 0.711 | 0.926 | 0.834 | 0.399 | 0.719 | 0.773 | - | - | - | - | |
| all | 0.687 | 0.954 | 0.847 | 0.405 | 0.68 | 0.408 | 0.757 | 0.506 | 0.746 | 0.458 | |
| nei- g5-50 | Bottle | 0.762 | 0.984 | 0.964 | 0.25 | 0.758 | nan | - | - | - | - |
| Can | 0.471 | 0.83 | 0.471 | 0.491 | 0.469 | nan | - | - | - | - | |
| Chain | 0.685 | 0.961 | 0.882 | 0.8 | 0.529 | 0.714 | - | - | - | - | |
| Drink-carton | 0.628 | 0.975 | 0.821 | 0.639 | 0.626 | nan | - | - | - | - | |
| Hook | 0.718 | 0.954 | 0.906 | nan | 0.718 | nan | - | - | - | - | |
| Propeller | 0.692 | 0.979 | 0.835 | 0.6 | 0.695 | 0.8 | - | - | - | - | |
| Shampoo-bottle | 0.714 | 0.961 | 0.892 | nan | 0.716 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.794 | 1 | 1 | nan | 0.794 | nan | - | - | - | - | |
| Tire | 0.806 | 0.969 | 0.958 | 0 | 0.821 | 0 | - | - | - | - | |
| Valve | 0.687 | 0.935 | 0.838 | 0 | 0.717 | 0 | - | - | - | - | |
| Wall | 0.728 | 0.914 | 0.824 | 0.342 | 0.743 | 0.795 | - | - | - | - | |
| all | 0.699 | 0.951 | 0.854 | 0.39 | 0.689 | 0.501 | 0.77 | 0.472 | 0.758 | 0.525 | |
| nei- p30 | Bottle | 0.758 | 0.985 | 0.974 | 0.3 | 0.754 | nan | - | - | - | - |
| Can | 0.489 | 0.843 | 0.473 | 0.541 | 0.481 | nan | - | - | - | - | |
| Chain | 0.693 | 0.967 | 0.855 | 0.75 | 0.522 | 0.723 | - | - | - | - | |
| Drink-carton | 0.638 | 0.995 | 0.813 | 0.645 | 0.639 | nan | - | - | - | - | |
| Hook | 0.693 | 0.934 | 0.934 | nan | 0.693 | nan | - | - | - | - | |
| Propeller | 0.654 | 0.973 | 0.799 | 0.5 | 0.66 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.674 | 0.983 | 0.87 | nan | 0.686 | 0.5 | - | - | - | - | |
| Standing-bottle | 0.772 | 1 | 0.822 | nan | 0.772 | nan | - | - | - | - | |
| Tire | 0.804 | 0.97 | 0.948 | 0.2 | 0.817 | 0 | - | - | - | - | |
| Valve | 0.677 | 0.92 | 0.881 | 0 | 0.706 | 0 | - | - | - | - | |
| Wall | 0.705 | 0.915 | 0.813 | 0.332 | 0.717 | 0.773 | - | - | - | - | |
| all | 0.687 | 0.953 | 0.835 | 0.409 | 0.677 | 0.433 | 0.758 | 0.495 | 0.751 | 0.46 | |
| nei- p5-50 | Bottle | 0.76 | 0.987 | 0.965 | 0.4 | 0.757 | nan | - | - | - | - |
| Can | 0.474 | 0.831 | 0.481 | 0.514 | 0.462 | nan | - | - | - | - | |
| Chain | 0.689 | 0.964 | 0.878 | 0.7 | 0.533 | 0.719 | - | - | - | - | |
| Drink-carton | 0.63 | 0.977 | 0.789 | 0.662 | 0.59 | nan | - | - | - | - | |
| Hook | 0.682 | 0.945 | 0.897 | nan | 0.682 | nan | - | - | - | - | |
| Propeller | 0.658 | 0.972 | 0.845 | 0.6 | 0.658 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.712 | 0.964 | 0.964 | nan | 0.726 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.767 | 1 | 1 | nan | 0.767 | nan | - | - | - | - | |
| Tire | 0.811 | 0.977 | 0.925 | 0 | 0.826 | 0 | - | - | - | - | |
| Valve | 0.682 | 0.915 | 0.836 | 0 | 0.712 | 0 | - | - | - | - | |
| Wall | 0.725 | 0.921 | 0.844 | 0.37 | 0.738 | 0.795 | - | - | - | - | |
| all | 0.69 | 0.95 | 0.857 | 0.406 | 0.678 | 0.469 | 0.759 | 0.446 | 0.748 | 0.492 |
| Denoising | ssd300 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | Bottle | 0.753 | 0.986 | 0.975 | 0.7 | 0.748 | nan | - | - | - | - |
| Can | 0.452 | 0.779 | 0.452 | 0.538 | 0.436 | nan | - | - | - | - | |
| Chain | 0.683 | 0.968 | 0.844 | 0.825 | 0.506 | 0.708 | - | - | - | - | |
| Drink-carton | 0.594 | 0.972 | 0.753 | 0.59 | 0.624 | nan | - | - | - | - | |
| Hook | 0.719 | 0.952 | 0.952 | nan | 0.719 | nan | - | - | - | - | |
| Propeller | 0.677 | 0.993 | 0.835 | 0.4 | 0.688 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.709 | 0.95 | 0.902 | nan | 0.738 | 0.3 | - | - | - | - | |
| Standing-bottle | 0.726 | 1 | 0.822 | nan | 0.726 | nan | - | - | - | - | |
| Tire | 0.8 | 0.998 | 0.935 | 0.6 | 0.805 | 0.5 | - | - | - | - | |
| Valve | 0.715 | 0.949 | 0.848 | 0.3 | 0.741 | 0 | - | - | - | - | |
| Wall | 0.71 | 0.926 | 0.831 | 0.279 | 0.715 | 0.818 | - | - | - | - | |
| all | 0.685 | 0.952 | 0.832 | 0.529 | 0.677 | 0.488 | 0.749 | 0.573 | 0.744 | 0.505 | |
| dspnet | Bottle | 0.737 | 0.982 | 0.926 | 0.233 | 0.734 | nan | - | - | - | - |
| Can | 0.436 | 0.845 | 0.491 | 0.487 | 0.433 | nan | - | - | - | - | |
| Chain | 0.677 | 0.972 | 0.792 | 0.701 | 0.483 | 0.714 | - | - | - | - | |
| Drink-carton | 0.562 | 0.942 | 0.634 | 0.55 | 0.585 | nan | - | - | - | - | |
| Hook | 0.709 | 0.959 | 0.931 | nan | 0.709 | nan | - | - | - | - | |
| Propeller | 0.635 | 0.974 | 0.732 | 0.5 | 0.641 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.705 | 0.951 | 0.951 | nan | 0.706 | 0.9 | - | - | - | - | |
| Standing-bottle | 0.702 | 1 | 0.861 | nan | 0.702 | nan | - | - | - | - | |
| Tire | 0.784 | 0.998 | 0.919 | 0.5 | 0.791 | 0.3 | - | - | - | - | |
| Valve | 0.706 | 0.932 | 0.862 | 0 | 0.736 | 0 | - | - | - | - | |
| Wall | 0.709 | 0.925 | 0.831 | 0.283 | 0.724 | 0.796 | - | - | - | - | |
| all | 0.669 | 0.953 | 0.812 | 0.407 | 0.659 | 0.552 | 0.736 | 0.515 | 0.721 | 0.57 | |
| b2ub- g25 | Bottle | 0.762 | 0.986 | 0.975 | 0.9 | 0.755 | nan | - | - | - | - |
| Can | 0.444 | 0.751 | 0.426 | 0.551 | 0.429 | nan | - | - | - | - | |
| Chain | 0.683 | 0.972 | 0.822 | 0.735 | 0.527 | 0.71 | - | - | - | - | |
| Drink-carton | 0.604 | 0.992 | 0.736 | 0.606 | 0.614 | nan | - | - | - | - | |
| Hook | 0.724 | 0.953 | 0.953 | nan | 0.724 | nan | - | - | - | - | |
| Propeller | 0.683 | 0.983 | 0.845 | 0.35 | 0.685 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.729 | 0.959 | 0.903 | nan | 0.757 | 0.3 | - | - | - | - | |
| Standing-bottle | 0.769 | 1 | 1 | nan | 0.769 | nan | - | - | - | - | |
| Tire | 0.811 | 0.999 | 0.955 | 0.6 | 0.817 | 0.5 | - | - | - | - | |
| Valve | 0.709 | 0.918 | 0.872 | 0 | 0.74 | 0 | - | - | - | - | |
| Wall | 0.714 | 0.92 | 0.83 | 0.347 | 0.716 | 0.823 | - | - | - | - | |
| all | 0.694 | 0.948 | 0.847 | 0.511 | 0.685 | 0.506 | 0.76 | 0.605 | 0.753 | 0.521 | |
| b2ub- g5-50 | Bottle | 0.755 | 0.985 | 0.962 | 0.267 | 0.749 | nan | - | - | - | - |
| Can | 0.45 | 0.778 | 0.496 | 0.538 | 0.435 | nan | - | - | - | - | |
| Chain | 0.711 | 0.974 | 0.848 | 0.9 | 0.546 | 0.732 | - | - | - | - | |
| Drink-carton | 0.62 | 0.991 | 0.77 | 0.621 | 0.644 | nan | - | - | - | - | |
| Hook | 0.726 | 0.953 | 0.903 | nan | 0.726 | nan | - | - | - | - | |
| Propeller | 0.684 | 0.978 | 0.784 | 0.6 | 0.69 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.766 | 0.971 | 0.971 | nan | 0.771 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.745 | 1 | 1 | nan | 0.745 | nan | - | - | - | - | |
| Tire | 0.804 | 0.989 | 0.954 | 0.4 | 0.809 | 0.6 | - | - | - | - | |
| Valve | 0.698 | 0.93 | 0.837 | 0 | 0.729 | 0 | - | - | - | - | |
| Wall | 0.725 | 0.931 | 0.832 | 0.356 | 0.73 | 0.813 | - | - | - | - | |
| all | 0.699 | 0.953 | 0.851 | 0.46 | 0.689 | 0.574 | 0.767 | 0.568 | 0.762 | 0.59 | |
| b2ub- p30 | Bottle | 0.745 | 0.984 | 0.973 | 0.233 | 0.74 | nan | - | - | - | - |
| Can | 0.455 | 0.751 | 0.498 | 0.585 | 0.429 | nan | - | - | - | - | |
| Chain | 0.679 | 0.956 | 0.867 | 0.65 | 0.519 | 0.713 | - | - | - | - | |
| Drink-carton | 0.624 | 0.995 | 0.782 | 0.634 | 0.623 | nan | - | - | - | - | |
| Hook | 0.727 | 0.956 | 0.956 | nan | 0.727 | nan | - | - | - | - | |
| Propeller | 0.679 | 0.988 | 0.851 | 0.5 | 0.688 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.74 | 0.969 | 0.969 | nan | 0.74 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.774 | 1 | 0.822 | nan | 0.774 | nan | - | - | - | - | |
| Tire | 0.811 | 0.999 | 0.954 | 0.2 | 0.819 | 0.5 | - | - | - | - | |
| Valve | 0.697 | 0.919 | 0.869 | 0 | 0.726 | 0 | - | - | - | - | |
| Wall | 0.718 | 0.922 | 0.821 | 0.351 | 0.724 | 0.81 | - | - | - | - | |
| all | 0.695 | 0.949 | 0.851 | 0.394 | 0.683 | 0.57 | 0.759 | 0.493 | 0.746 | 0.588 | |
| b2ub- p5-50 | Bottle | 0.754 | 0.985 | 0.962 | 0.8 | 0.749 | nan | - | - | - | - |
| Can | 0.453 | 0.761 | 0.454 | 0.527 | 0.452 | nan | - | - | - | - | |
| Chain | 0.684 | 0.95 | 0.815 | 0.9 | 0.502 | 0.714 | - | - | - | - | |
| Drink-carton | 0.629 | 0.991 | 0.754 | 0.631 | 0.638 | nan | - | - | - | - | |
| Hook | 0.722 | 0.953 | 0.934 | nan | 0.722 | nan | - | - | - | - | |
| Propeller | 0.697 | 0.983 | 0.837 | 0.3 | 0.705 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.762 | 0.947 | 0.947 | nan | 0.767 | 0.9 | - | - | - | - | |
| Standing-bottle | 0.767 | 1 | 0.822 | nan | 0.767 | nan | - | - | - | - | |
| Tire | 0.816 | 0.989 | 0.956 | 0.5 | 0.823 | 0.5 | - | - | - | - | |
| Valve | 0.705 | 0.93 | 0.858 | 0 | 0.735 | 0 | - | - | - | - | |
| Wall | 0.721 | 0.921 | 0.832 | 0.314 | 0.724 | 0.83 | - | - | - | - | |
| all | 0.701 | 0.946 | 0.834 | 0.497 | 0.689 | 0.591 | 0.768 | 0.577 | 0.759 | 0.607 | |
| nei- g25 | Bottle | 0.744 | 0.986 | 0.953 | 0.7 | 0.738 | nan | - | - | - | - |
| Can | 0.459 | 0.794 | 0.498 | 0.589 | 0.435 | nan | - | - | - | - | |
| Chain | 0.702 | 0.955 | 0.822 | 0.75 | 0.514 | 0.733 | - | - | - | - | |
| Drink-carton | 0.614 | 0.997 | 0.734 | 0.613 | 0.624 | nan | - | - | - | - | |
| Hook | 0.735 | 0.954 | 0.903 | nan | 0.735 | nan | - | - | - | - | |
| Propeller | 0.708 | 0.981 | 0.824 | 0.4 | 0.717 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.745 | 0.954 | 0.954 | nan | 0.756 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.73 | 1 | 1 | nan | 0.73 | nan | - | - | - | - | |
| Tire | 0.809 | 0.998 | 0.955 | 0.6 | 0.814 | 0.4 | - | - | - | - | |
| Valve | 0.729 | 0.939 | 0.88 | 0.5 | 0.754 | 0 | - | - | - | - | |
| Wall | 0.715 | 0.93 | 0.826 | 0.344 | 0.709 | 0.836 | - | - | - | - | |
| all | 0.699 | 0.954 | 0.85 | 0.562 | 0.684 | 0.561 | 0.762 | 0.597 | 0.749 | 0.577 | |
| nei- g5-50 | Bottle | 0.757 | 0.986 | 0.974 | 0.35 | 0.752 | nan | - | - | - | - |
| Can | 0.439 | 0.752 | 0.446 | 0.503 | 0.431 | nan | - | - | - | - | |
| Chain | 0.698 | 0.97 | 0.839 | 0.8 | 0.501 | 0.727 | - | - | - | - | |
| Drink-carton | 0.611 | 0.993 | 0.739 | 0.612 | 0.625 | nan | - | - | - | - | |
| Hook | 0.727 | 0.952 | 0.952 | nan | 0.727 | nan | - | - | - | - | |
| Propeller | 0.693 | 0.977 | 0.861 | 0.6 | 0.696 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.759 | 0.974 | 0.912 | nan | 0.778 | 0.5 | - | - | - | - | |
| Standing-bottle | 0.749 | 1 | 1 | nan | 0.749 | nan | - | - | - | - | |
| Tire | 0.815 | 0.99 | 0.967 | 0.6 | 0.817 | 0.6 | - | - | - | - | |
| Valve | 0.711 | 0.93 | 0.868 | 0.2 | 0.741 | 0 | - | - | - | - | |
| Wall | 0.722 | 0.931 | 0.844 | 0.335 | 0.72 | 0.837 | - | - | - | - | |
| all | 0.698 | 0.95 | 0.855 | 0.5 | 0.685 | 0.561 | 0.763 | 0.592 | 0.754 | 0.576 | |
| nei- p30 | Bottle | 0.749 | 0.984 | 0.962 | 0.267 | 0.743 | nan | - | - | - | - |
| Can | 0.444 | 0.734 | 0.43 | 0.539 | 0.431 | nan | - | - | - | - | |
| Chain | 0.671 | 0.956 | 0.843 | 0.783 | 0.534 | 0.694 | - | - | - | - | |
| Drink-carton | 0.645 | 0.996 | 0.836 | 0.63 | 0.682 | nan | - | - | - | - | |
| Hook | 0.708 | 0.959 | 0.959 | nan | 0.708 | nan | - | - | - | - | |
| Propeller | 0.665 | 0.988 | 0.772 | 0.5 | 0.674 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.773 | 0.962 | 0.962 | nan | 0.776 | 0.8 | - | - | - | - | |
| Standing-bottle | 0.729 | 1 | 1 | nan | 0.729 | nan | - | - | - | - | |
| Tire | 0.823 | 0.999 | 0.962 | 0.6 | 0.829 | 0.5 | - | - | - | - | |
| Valve | 0.71 | 0.918 | 0.871 | 0 | 0.741 | 0 | - | - | - | - | |
| Wall | 0.723 | 0.931 | 0.821 | 0.359 | 0.72 | 0.835 | - | - | - | - | |
| all | 0.694 | 0.948 | 0.856 | 0.46 | 0.688 | 0.571 | 0.763 | 0.576 | 0.756 | 0.589 | |
| nei- p5-50 | Bottle | 0.754 | 0.984 | 0.951 | 0.4 | 0.748 | nan | - | - | - | - |
| Can | 0.461 | 0.766 | 0.436 | 0.53 | 0.448 | nan | - | - | - | - | |
| Chain | 0.698 | 0.955 | 0.855 | 0.834 | 0.544 | 0.723 | - | - | - | - | |
| Drink-carton | 0.624 | 0.993 | 0.811 | 0.615 | 0.64 | nan | - | - | - | - | |
| Hook | 0.717 | 0.954 | 0.906 | nan | 0.717 | nan | - | - | - | - | |
| Propeller | 0.698 | 0.986 | 0.804 | 0.4 | 0.71 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.746 | 0.955 | 0.9 | nan | 0.765 | 0.5 | - | - | - | - | |
| Standing-bottle | 0.76 | 1 | 1 | nan | 0.76 | nan | - | - | - | - | |
| Tire | 0.81 | 0.989 | 0.937 | 0.4 | 0.815 | 0.6 | - | - | - | - | |
| Valve | 0.706 | 0.931 | 0.8 | 0 | 0.736 | 0 | - | - | - | - | |
| Wall | 0.714 | 0.923 | 0.816 | 0.337 | 0.709 | 0.825 | - | - | - | - | |
| all | 0.699 | 0.949 | 0.838 | 0.44 | 0.69 | 0.541 | 0.76 | 0.537 | 0.754 | 0.558 |
| Denoising | YOLOX | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | Bottle | 0.74 | 0.986 | 0.964 | 0.3 | 0.735 | nan | - | - | - | - |
| Can | 0.503 | 0.857 | 0.594 | 0.541 | 0.492 | nan | - | - | - | - | |
| Chain | 0.678 | 0.959 | 0.785 | 0.75 | 0.51 | 0.708 | - | - | - | - | |
| Drink-carton | 0.621 | 0.999 | 0.729 | 0.619 | 0.632 | nan | - | - | - | - | |
| Hook | 0.725 | 0.967 | 0.933 | nan | 0.725 | nan | - | - | - | - | |
| Propeller | 0.681 | 0.968 | 0.774 | 0.125 | 0.688 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.616 | 0.971 | 0.701 | nan | 0.62 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.712 | 1 | 0.851 | nan | 0.712 | nan | - | - | - | - | |
| Tire | 0.799 | 0.99 | 0.937 | 0.8 | 0.804 | 0.4 | - | - | - | - | |
| Valve | 0.654 | 0.927 | 0.8 | 0 | 0.683 | 0 | - | - | - | - | |
| Wall | 0.69 | 0.939 | 0.845 | 0.384 | 0.685 | 0.756 | - | - | - | - | |
| all | 0.675 | 0.96 | 0.81 | 0.44 | 0.662 | 0.527 | 0.736 | 0.564 | 0.728 | 0.544 | |
| dspnet | Bottle | 0.735 | 0.985 | 0.974 | 0.233 | 0.73 | nan | - | - | - | - |
| Can | 0.432 | 0.748 | 0.489 | 0.489 | 0.424 | nan | - | - | - | - | |
| Chain | 0.664 | 0.944 | 0.819 | 0.651 | 0.527 | 0.699 | - | - | - | - | |
| Drink-carton | 0.569 | 0.956 | 0.608 | 0.566 | 0.603 | nan | - | - | - | - | |
| Hook | 0.7 | 0.955 | 0.855 | nan | 0.7 | nan | - | - | - | - | |
| Propeller | 0.658 | 0.945 | 0.838 | 0.25 | 0.663 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.632 | 0.951 | 0.776 | nan | 0.636 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.657 | 1 | 0.832 | nan | 0.657 | nan | - | - | - | - | |
| Tire | 0.798 | 0.979 | 0.936 | 0.7 | 0.805 | 0 | - | - | - | - | |
| Valve | 0.668 | 0.927 | 0.822 | 0 | 0.699 | 0 | - | - | - | - | |
| Wall | 0.666 | 0.929 | 0.83 | 0.332 | 0.666 | 0.734 | - | - | - | - | |
| all | 0.653 | 0.938 | 0.798 | 0.403 | 0.646 | 0.456 | 0.712 | 0.526 | 0.705 | 0.475 | |
| b2ub- g25 | Bottle | 0.746 | 0.984 | 0.964 | 0.6 | 0.742 | nan | - | - | - | - |
| Can | 0.475 | 0.838 | 0.485 | 0.552 | 0.461 | nan | - | - | - | - | |
| Chain | 0.687 | 0.947 | 0.814 | 0.8 | 0.543 | 0.71 | - | - | - | - | |
| Drink-carton | 0.633 | 0.998 | 0.772 | 0.622 | 0.662 | nan | - | - | - | - | |
| Hook | 0.726 | 0.947 | 0.913 | nan | 0.726 | nan | - | - | - | - | |
| Propeller | 0.686 | 0.964 | 0.833 | 0.5 | 0.695 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.657 | 0.971 | 0.971 | nan | 0.657 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.749 | 1 | 0.851 | nan | 0.749 | nan | - | - | - | - | |
| Tire | 0.802 | 0.99 | 0.938 | 0.7 | 0.809 | 0.1 | - | - | - | - | |
| Valve | 0.684 | 0.931 | 0.861 | 0 | 0.714 | 0 | - | - | - | - | |
| Wall | 0.686 | 0.931 | 0.836 | 0.38 | 0.672 | 0.765 | - | - | - | - | |
| all | 0.685 | 0.955 | 0.84 | 0.519 | 0.675 | 0.496 | 0.744 | 0.562 | 0.738 | 0.513 | |
| b2ub- g5-50 | Bottle | 0.724 | 0.985 | 0.955 | 0.3 | 0.72 | nan | - | - | - | - |
| Can | 0.487 | 0.825 | 0.562 | 0.539 | 0.481 | nan | - | - | - | - | |
| Chain | 0.676 | 0.946 | 0.775 | 0.725 | 0.538 | 0.705 | - | - | - | - | |
| Drink-carton | 0.611 | 0.995 | 0.692 | 0.61 | 0.616 | nan | - | - | - | - | |
| Hook | 0.717 | 0.944 | 0.9 | nan | 0.717 | nan | - | - | - | - | |
| Propeller | 0.713 | 0.985 | 0.841 | 0.15 | 0.721 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.715 | 0.954 | 0.954 | nan | 0.726 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.709 | 1 | 0.861 | nan | 0.709 | nan | - | - | - | - | |
| Tire | 0.82 | 0.99 | 0.949 | 0.9 | 0.826 | 0.2 | - | - | - | - | |
| Valve | 0.71 | 0.928 | 0.839 | 0 | 0.741 | 0 | - | - | - | - | |
| Wall | 0.701 | 0.932 | 0.836 | 0.338 | 0.69 | 0.779 | - | - | - | - | |
| all | 0.689 | 0.953 | 0.833 | 0.445 | 0.68 | 0.497 | 0.748 | 0.589 | 0.739 | 0.513 | |
| b2ub- p30 | Bottle | 0.754 | 0.986 | 0.964 | 0.7 | 0.749 | nan | - | - | - | - |
| Can | 0.485 | 0.836 | 0.544 | 0.571 | 0.472 | nan | - | - | - | - | |
| Chain | 0.688 | 0.947 | 0.809 | 0.735 | 0.524 | 0.72 | - | - | - | - | |
| Drink-carton | 0.607 | 0.979 | 0.706 | 0.608 | 0.612 | nan | - | - | - | - | |
| Hook | 0.732 | 0.958 | 0.936 | nan | 0.732 | nan | - | - | - | - | |
| Propeller | 0.725 | 0.984 | 0.888 | 0.2 | 0.733 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.691 | 0.964 | 0.884 | nan | 0.707 | 0.5 | - | - | - | - | |
| Standing-bottle | 0.748 | 1 | 1 | nan | 0.748 | nan | - | - | - | - | |
| Tire | 0.821 | 0.99 | 0.959 | 0.8 | 0.827 | 0.3 | - | - | - | - | |
| Valve | 0.683 | 0.923 | 0.835 | 0 | 0.713 | 0 | - | - | - | - | |
| Wall | 0.707 | 0.929 | 0.867 | 0.397 | 0.693 | 0.777 | - | - | - | - | |
| all | 0.695 | 0.954 | 0.854 | 0.501 | 0.683 | 0.5 | 0.752 | 0.6 | 0.74 | 0.517 | |
| b2ub- p5-50 | Bottle | 0.752 | 0.983 | 0.961 | 0.3 | 0.747 | nan | - | - | - | - |
| Can | 0.478 | 0.824 | 0.522 | 0.586 | 0.446 | nan | - | - | - | - | |
| Chain | 0.691 | 0.949 | 0.857 | 0.8 | 0.548 | 0.721 | - | - | - | - | |
| Drink-carton | 0.637 | 0.997 | 0.792 | 0.626 | 0.671 | nan | - | - | - | - | |
| Hook | 0.709 | 0.963 | 0.91 | nan | 0.709 | nan | - | - | - | - | |
| Propeller | 0.702 | 0.972 | 0.869 | 0.06 | 0.717 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.687 | 0.951 | 0.875 | nan | 0.694 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.72 | 1 | 0.891 | nan | 0.72 | nan | - | - | - | - | |
| Tire | 0.812 | 0.999 | 0.938 | 0.8 | 0.816 | 0.3 | - | - | - | - | |
| Valve | 0.686 | 0.946 | 0.836 | 0.65 | 0.701 | 0 | - | - | - | - | |
| Wall | 0.701 | 0.93 | 0.835 | 0.369 | 0.702 | 0.769 | - | - | - | - | |
| all | 0.689 | 0.956 | 0.844 | 0.524 | 0.679 | 0.515 | 0.747 | 0.66 | 0.739 | 0.533 | |
| nei- g25 | Bottle | 0.759 | 0.985 | 0.964 | 0.25 | 0.755 | nan | - | - | - | - |
| Can | 0.483 | 0.793 | 0.534 | 0.58 | 0.461 | nan | - | - | - | - | |
| Chain | 0.701 | 0.947 | 0.83 | 0.651 | 0.627 | 0.72 | - | - | - | - | |
| Drink-carton | 0.629 | 0.997 | 0.715 | 0.622 | 0.65 | nan | - | - | - | - | |
| Hook | 0.747 | 0.963 | 0.94 | nan | 0.747 | nan | - | - | - | - | |
| Propeller | 0.688 | 0.971 | 0.834 | 0.1 | 0.7 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.693 | 0.954 | 0.954 | nan | 0.699 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.7 | 1 | 1 | nan | 0.7 | nan | - | - | - | - | |
| Tire | 0.829 | 0.999 | 0.959 | 0.6 | 0.835 | 0.4 | - | - | - | - | |
| Valve | 0.676 | 0.925 | 0.818 | 0 | 0.706 | 0 | - | - | - | - | |
| Wall | 0.717 | 0.936 | 0.863 | 0.384 | 0.707 | 0.791 | - | - | - | - | |
| all | 0.693 | 0.952 | 0.856 | 0.398 | 0.69 | 0.535 | 0.754 | 0.509 | 0.752 | 0.552 | |
| nei- g5-50 | Bottle | 0.755 | 0.985 | 0.962 | 0.6 | 0.75 | nan | - | - | - | - |
| Can | 0.49 | 0.847 | 0.58 | 0.572 | 0.466 | nan | - | - | - | - | |
| Chain | 0.686 | 0.963 | 0.824 | 0.801 | 0.526 | 0.717 | - | - | - | - | |
| Drink-carton | 0.639 | 0.998 | 0.81 | 0.63 | 0.662 | nan | - | - | - | - | |
| Hook | 0.73 | 0.959 | 0.924 | nan | 0.73 | nan | - | - | - | - | |
| Propeller | 0.69 | 0.975 | 0.808 | 0.12 | 0.7 | 0.6 | - | - | - | - | |
| Shampoo-bottle | 0.672 | 0.944 | 0.793 | nan | 0.682 | 0.5 | - | - | - | - | |
| Standing-bottle | 0.743 | 1 | 0.88 | nan | 0.743 | nan | - | - | - | - | |
| Tire | 0.832 | 0.999 | 0.966 | 0.6 | 0.84 | 0.2 | - | - | - | - | |
| Valve | 0.683 | 0.926 | 0.879 | 0 | 0.713 | 0 | - | - | - | - | |
| Wall | 0.669 | 0.928 | 0.807 | 0.403 | 0.669 | 0.729 | - | - | - | - | |
| all | 0.69 | 0.957 | 0.839 | 0.466 | 0.68 | 0.458 | 0.75 | 0.561 | 0.739 | 0.479 | |
| nei- p30 | Bottle | 0.737 | 0.987 | 0.954 | 0.6 | 0.731 | nan | - | - | - | - |
| Can | 0.481 | 0.811 | 0.555 | 0.562 | 0.46 | nan | - | - | - | - | |
| Chain | 0.697 | 0.946 | 0.82 | 0.825 | 0.55 | 0.723 | - | - | - | - | |
| Drink-carton | 0.634 | 0.999 | 0.78 | 0.635 | 0.643 | nan | - | - | - | - | |
| Hook | 0.724 | 0.958 | 0.958 | nan | 0.724 | nan | - | - | - | - | |
| Propeller | 0.692 | 0.976 | 0.848 | 0.25 | 0.698 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.668 | 0.961 | 0.961 | nan | 0.676 | 0.6 | - | - | - | - | |
| Standing-bottle | 0.728 | 1 | 0.79 | nan | 0.728 | nan | - | - | - | - | |
| Tire | 0.818 | 0.999 | 0.957 | 0.7 | 0.824 | 0.2 | - | - | - | - | |
| Valve | 0.655 | 0.92 | 0.813 | 0 | 0.683 | 0 | - | - | - | - | |
| Wall | 0.701 | 0.936 | 0.853 | 0.383 | 0.693 | 0.769 | - | - | - | - | |
| all | 0.685 | 0.954 | 0.844 | 0.494 | 0.674 | 0.499 | 0.738 | 0.569 | 0.728 | 0.515 | |
| nei- p5-50 | Bottle | 0.739 | 0.986 | 0.952 | 0.35 | 0.733 | nan | - | - | - | - |
| Can | 0.488 | 0.801 | 0.594 | 0.587 | 0.465 | nan | - | - | - | - | |
| Chain | 0.681 | 0.944 | 0.827 | 0.768 | 0.523 | 0.704 | - | - | - | - | |
| Drink-carton | 0.615 | 0.999 | 0.691 | 0.596 | 0.655 | nan | - | - | - | - | |
| Hook | 0.696 | 0.965 | 0.89 | nan | 0.696 | nan | - | - | - | - | |
| Propeller | 0.694 | 0.984 | 0.814 | 0.2 | 0.7 | 0.7 | - | - | - | - | |
| Shampoo-bottle | 0.635 | 0.951 | 0.849 | nan | 0.632 | 0.7 | - | - | - | - | |
| Standing-bottle | 0.701 | 1 | 0.871 | nan | 0.701 | nan | - | - | - | - | |
| Tire | 0.792 | 0.99 | 0.938 | 0.8 | 0.797 | 0.2 | - | - | - | - | |
| Valve | 0.675 | 0.926 | 0.835 | 0 | 0.704 | 0 | - | - | - | - | |
| Wall | 0.672 | 0.937 | 0.823 | 0.344 | 0.678 | 0.723 | - | - | - | - | |
| all | 0.672 | 0.953 | 0.826 | 0.456 | 0.662 | 0.505 | 0.729 | 0.605 | 0.724 | 0.522 |
| Denoising | ssd_mobilenetv2 | ||||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |||
| noisy | victim | 0.825 | 1 | 0.96 | nan | 0.713 | 0.828 | - | - | - | - | ||
| plane | 0.779 | 1 | 0.953 | nan | 0.781 | 0.725 | - | - | - | - | |||
| boat | 0.733 | 0.99 | 0.938 | 0.9 | 0.732 | nan | - | - | - | - | |||
| all | 0.779 | 0.996 | 0.95 | 0.9 | 0.742 | 0.776 | 0.82 | 0.9 | 0.774 | 0.804 | |||
| dspnet | victim | 0.824 | 1 | 0.962 | nan | 0.707 | 0.827 | - | - | - | - | ||
| plane | 0.772 | 1 | 0.95 | nan | 0.773 | 0.801 | - | - | - | - | |||
| boat | 0.717 | 0.989 | 0.91 | 0.8 | 0.718 | nan | - | - | - | - | |||
| all | 0.771 | 0.996 | 0.941 | 0.8 | 0.733 | 0.814 | 0.819 | 0.8 | 0.77 | 0.832 | |||
| b2ub- g25 | victim | 0.812 | 1 | 0.951 | nan | 0.647 | 0.817 | - | - | - | - | ||
| plane | 0.738 | 1 | 0.929 | nan | 0.74 | 0.7 | - | - | - | - | |||
| boat | 0.592 | 0.98 | 0.653 | 0.7 | 0.592 | nan | - | - | - | - | |||
| all | 0.714 | 0.993 | 0.844 | 0.7 | 0.66 | 0.758 | 0.758 | 0.7 | 0.7 | 0.825 | |||
| b2ub- g5-50 | victim | 0.831 | 1 | 0.963 | nan | 0.769 | 0.833 | - | - | - | - | ||
| plane | 0.777 | 1 | 0.965 | nan | 0.778 | 0.75 | - | - | - | - | |||
| boat | 0.716 | 0.99 | 0.913 | 0.8 | 0.716 | nan | - | - | - | - | |||
| all | 0.775 | 0.997 | 0.947 | 0.8 | 0.754 | 0.792 | 0.818 | 0.8 | 0.788 | 0.807 | |||
| b2ub- p30 | victim | 0.778 | 0.999 | 0.921 | nan | 0.565 | 0.784 | - | - | - | - | ||
| plane | 0.7 | 0.989 | 0.944 | nan | 0.7 | 0.767 | - | - | - | - | |||
| boat | 0.59 | 0.989 | 0.655 | 0.8 | 0.589 | nan | - | - | - | - | |||
| all | 0.689 | 0.993 | 0.84 | 0.8 | 0.618 | 0.775 | 0.737 | 0.8 | 0.657 | 0.809 | |||
| b2ub- p5-50 | victim | 0.807 | 1 | 0.936 | nan | 0.653 | 0.809 | - | - | - | - | ||
| plane | 0.742 | 1 | 0.942 | nan | 0.744 | 0.759 | - | - | - | - | |||
| boat | 0.669 | 0.99 | 0.823 | 0.7 | 0.668 | nan | - | - | - | - | |||
| all | 0.739 | 0.997 | 0.901 | 0.7 | 0.688 | 0.784 | 0.787 | 0.7 | 0.74 | 0.846 | |||
| nei- g25 | victim | 0.809 | 1 | 0.967 | nan | 0.671 | 0.813 | - | - | - | - | ||
| plane | 0.725 | 1 | 0.918 | nan | 0.731 | 0.683 | - | - | - | - | |||
| boat | 0.58 | 0.989 | 0.688 | 0.5 | 0.58 | nan | - | - | - | - | |||
| all | 0.705 | 0.996 | 0.858 | 0.5 | 0.661 | 0.748 | 0.754 | 0.5 | 0.705 | 0.775 | |||
| nei- g5-50 | ship | 0.516 | 0.783 | 0.603 | 0 | 0.466 | 0.55 | - | - | - | - | ||
| aircraft | 0.368 | 0.739 | 0.347 | nan | 0.552 | 0.326 | - | - | - | - | |||
| human | 0.35 | 0.487 | 0.446 | 0.8 | 0.314 | 0.303 | - | - | - | - | |||
| all | 0.411 | 0.67 | 0.465 | 0.4 | 0.444 | 0.393 | 0.477 | 0.4 | 0.452 | 0.46 | |||
| nei- p30 | victim | 0.751 | 0.989 | 0.899 | nan | 0.533 | 0.756 | - | - | - | - | ||
| plane | 0.696 | 0.999 | 0.866 | nan | 0.695 | 0.701 | - | - | - | - | |||
| boat | 0.539 | 0.988 | 0.526 | 0.2 | 0.54 | nan | - | - | - | - | |||
| all | 0.662 | 0.992 | 0.764 | 0.2 | 0.59 | 0.728 | 0.712 | 0.2 | 0.665 | 0.745 | |||
| nei- p5-50 | victim | 0.827 | 1 | 0.971 | nan | 0.681 | 0.832 | - | - | - | - | ||
| plane | 0.796 | 1 | 0.968 | nan | 0.798 | 0.801 | - | - | - | - | |||
| boat | 0.74 | 0.99 | 0.94 | 0.7 | 0.741 | nan | - | - | - | - | |||
| all | 0.788 | 0.997 | 0.96 | 0.7 | 0.74 | 0.816 | 0.826 | 0.7 | 0.778 | 0.831 | |||
| Denoising | Faster r-cnn | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | victim | 0.807 | 0.99 | 0.935 | nan | 0.666 | 0.809 | - | - | - | - |
| plane | 0.879 | 1 | 0.987 | nan | 0.88 | 0.725 | - | - | - | - | |
| boat | 0.849 | 0.987 | 0.987 | 0.9 | 0.849 | nan | - | - | - | - | |
| all | 0.845 | 0.992 | 0.969 | 0.9 | 0.798 | 0.767 | 0.887 | 0.9 | 0.827 | 0.803 | |
| dspnet | victim | 0.802 | 1 | 0.928 | nan | 0.532 | 0.807 | - | - | - | - |
| plane | 0.87 | 1 | 1 | nan | 0.873 | 0.801 | - | - | - | - | |
| boat | 0.833 | 0.987 | 0.987 | 0.9 | 0.833 | nan | - | - | - | - | |
| all | 0.835 | 0.996 | 0.972 | 0.9 | 0.746 | 0.804 | 0.876 | 0.9 | 0.795 | 0.825 | |
| b2ub- g25 | victim | 0.81 | 1 | 0.947 | nan | 0.641 | 0.815 | - | - | - | - |
| plane | 0.88 | 1 | 0.986 | nan | 0.883 | 0.75 | - | - | - | - | |
| boat | 0.84 | 0.988 | 0.988 | 0.8 | 0.84 | nan | - | - | - | - | |
| all | 0.843 | 0.996 | 0.974 | 0.8 | 0.788 | 0.783 | 0.878 | 0.8 | 0.816 | 0.825 | |
| b2ub- g5-50 | victim | 0.815 | 1 | 0.942 | nan | 0.714 | 0.817 | - | - | - | - |
| plane | 0.88 | 1 | 1 | nan | 0.881 | 0.75 | - | - | - | - | |
| boat | 0.855 | 0.987 | 0.987 | 0.9 | 0.855 | nan | - | - | - | - | |
| all | 0.85 | 0.996 | 0.977 | 0.9 | 0.817 | 0.784 | 0.89 | 0.9 | 0.843 | 0.83 | |
| b2ub- p30 | victim | 0.822 | 1 | 0.954 | nan | 0.661 | 0.825 | - | - | - | - |
| plane | 0.885 | 1 | 0.987 | nan | 0.886 | 0.85 | - | - | - | - | |
| boat | 0.855 | 0.987 | 0.987 | 0.9 | 0.855 | nan | - | - | - | - | |
| all | 0.854 | 0.996 | 0.976 | 0.9 | 0.801 | 0.838 | 0.89 | 0.9 | 0.829 | 0.854 | |
| b2ub- p5-50 | victim | 0.804 | 0.99 | 0.947 | nan | 0.516 | 0.811 | - | - | - | - |
| plane | 0.86 | 1 | 1 | nan | 0.861 | 0.8 | - | - | - | - | |
| boat | 0.836 | 0.988 | 0.978 | 0.9 | 0.836 | nan | - | - | - | - | |
| all | 0.833 | 0.993 | 0.975 | 0.9 | 0.737 | 0.806 | 0.87 | 0.9 | 0.768 | 0.822 | |
| nei- g25 | victim | 0.804 | 1 | 0.939 | nan | 0.638 | 0.809 | - | - | - | - |
| plane | 0.875 | 1 | 0.986 | nan | 0.877 | 0.75 | - | - | - | - | |
| boat | 0.844 | 0.988 | 0.988 | 0.9 | 0.844 | nan | - | - | - | - | |
| all | 0.841 | 0.996 | 0.971 | 0.9 | 0.786 | 0.78 | 0.881 | 0.9 | 0.82 | 0.799 | |
| nei- g5-50 | ship | 0.822 | 1 | 0.954 | nan | 0.653 | 0.826 | - | - | - | - |
| aircraft | 0.881 | 1 | 0.99 | nan | 0.882 | 0.75 | - | - | - | - | |
| human | 0.844 | 0.986 | 0.986 | 0.9 | 0.844 | nan | - | - | - | - | |
| all | 0.849 | 0.995 | 0.977 | 0.9 | 0.793 | 0.788 | 0.889 | 0.9 | 0.822 | 0.833 | |
| nei- p30 | victim | 0.797 | 0.99 | 0.927 | nan | 0.612 | 0.803 | - | - | - | - |
| plane | 0.866 | 1 | 0.988 | nan | 0.867 | 0.75 | - | - | - | - | |
| boat | 0.839 | 0.986 | 0.986 | 0.9 | 0.839 | nan | - | - | - | - | |
| all | 0.834 | 0.992 | 0.967 | 0.9 | 0.773 | 0.777 | 0.873 | 0.9 | 0.8 | 0.795 | |
| nei- p5-50 | victim | 0.827 | 1 | 0.965 | nan | 0.599 | 0.834 | - | - | - | - |
| plane | 0.891 | 1 | 0.987 | nan | 0.893 | 0.75 | - | - | - | - | |
| boat | 0.851 | 0.986 | 0.986 | 0.9 | 0.851 | nan | - | - | - | - | |
| all | 0.856 | 0.995 | 0.979 | 0.9 | 0.781 | 0.792 | 0.89 | 0.9 | 0.804 | 0.832 | |
| Denoising | ssd300 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | victim | 0.805 | 1 | 0.956 | nan | 0.634 | 0.81 | - | - | - | - |
| plane | 0.842 | 1 | 0.975 | nan | 0.843 | 0.8 | - | - | - | - | |
| boat | 0.813 | 0.989 | 0.974 | 0.85 | 0.814 | nan | - | - | - | - | |
| all | 0.82 | 0.996 | 0.968 | 0.85 | 0.764 | 0.805 | 0.861 | 0.9 | 0.8 | 0.848 | |
| dspnet | victim | 0.802 | 1 | 0.938 | nan | 0.66 | 0.806 | - | - | - | - |
| plane | 0.833 | 1 | 0.973 | nan | 0.835 | 0.775 | - | - | - | - | |
| boat | 0.797 | 0.989 | 0.977 | 0.8 | 0.797 | nan | - | - | - | - | |
| all | 0.811 | 0.996 | 0.963 | 0.8 | 0.764 | 0.79 | 0.851 | 0.8 | 0.799 | 0.872 | |
| b2ub- g25 | victim | 0.805 | 1 | 0.956 | nan | 0.613 | 0.811 | - | - | - | - |
| plane | 0.847 | 1 | 0.977 | nan | 0.85 | 0.767 | - | - | - | - | |
| boat | 0.811 | 0.99 | 0.978 | 0.9 | 0.811 | nan | - | - | - | - | |
| all | 0.821 | 0.997 | 0.97 | 0.9 | 0.758 | 0.789 | 0.863 | 0.9 | 0.795 | 0.851 | |
| b2ub- g5-50 | victim | 0.814 | 1 | 0.963 | nan | 0.678 | 0.817 | - | - | - | - |
| plane | 0.85 | 1 | 0.987 | nan | 0.851 | 0.8 | - | - | - | - | |
| boat | 0.814 | 0.99 | 0.975 | 0.9 | 0.814 | nan | - | - | - | - | |
| all | 0.826 | 0.997 | 0.975 | 0.9 | 0.781 | 0.809 | 0.864 | 0.9 | 0.815 | 0.851 | |
| b2ub- p30 | victim | 0.807 | 1 | 0.947 | nan | 0.678 | 0.811 | - | - | - | - |
| plane | 0.857 | 1 | 0.989 | nan | 0.859 | 0.725 | - | - | - | - | |
| boat | 0.817 | 0.989 | 0.976 | 0.9 | 0.817 | nan | - | - | - | - | |
| all | 0.827 | 0.996 | 0.971 | 0.9 | 0.785 | 0.768 | 0.866 | 0.9 | 0.813 | 0.801 | |
| b2ub- p5-50 | victim | 0.791 | 1 | 0.952 | nan | 0.626 | 0.796 | - | - | - | - |
| plane | 0.805 | 1 | 0.974 | nan | 0.805 | 0.725 | - | - | - | - | |
| boat | 0.78 | 0.989 | 0.975 | 0.8 | 0.781 | nan | - | - | - | - | |
| all | 0.792 | 0.996 | 0.967 | 0.8 | 0.737 | 0.761 | 0.83 | 0.8 | 0.784 | 0.789 | |
| nei- g25 | victim | 0.811 | 1 | 0.966 | nan | 0.699 | 0.813 | - | - | - | - |
| plane | 0.85 | 1 | 0.988 | nan | 0.853 | 0.719 | - | - | - | - | |
| boat | 0.814 | 0.989 | 0.974 | 0.7 | 0.815 | nan | - | - | - | - | |
| all | 0.825 | 0.996 | 0.976 | 0.7 | 0.789 | 0.766 | 0.864 | 0.7 | 0.823 | 0.827 | |
| nei- g5-50 | ship | 0.805 | 1 | 0.975 | nan | 0.616 | 0.81 | - | - | - | - |
| aircraft | 0.845 | 1 | 0.989 | nan | 0.848 | 0.717 | - | - | - | - | |
| human | 0.809 | 0.99 | 0.975 | 0.9 | 0.809 | nan | - | - | - | - | |
| all | 0.82 | 0.997 | 0.979 | 0.9 | 0.758 | 0.764 | 0.86 | 0.9 | 0.794 | 0.8 | |
| nei- p30 | victim | 0.788 | 1 | 0.942 | nan | 0.63 | 0.792 | - | - | - | - |
| plane | 0.838 | 1 | 0.988 | nan | 0.839 | 0.846 | - | - | - | - | |
| boat | 0.804 | 0.99 | 0.976 | 0.9 | 0.804 | nan | - | - | - | - | |
| all | 0.81 | 0.997 | 0.969 | 0.9 | 0.757 | 0.819 | 0.85 | 0.9 | 0.785 | 0.865 | |
| nei- p5-50 | victim | 0.812 | 1 | 0.961 | nan | 0.667 | 0.815 | - | - | - | - |
| plane | 0.854 | 1 | 0.987 | nan | 0.856 | 0.742 | - | - | - | - | |
| boat | 0.815 | 0.989 | 0.963 | 0.8 | 0.815 | nan | - | - | - | - | |
| all | 0.827 | 0.996 | 0.97 | 0.8 | 0.779 | 0.864 | 0.8 | 0.811 | 0.825 | 0.825 | |
| Denoising | YOLOX | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | victim | 0.83 | 1 | 0.964 | nan | 0.627 | 0.838 | - | - | - | - |
| plane | 0.853 | 1 | 0.98 | nan | 0.853 | 0.817 | - | - | - | - | |
| boat | 0.825 | 0.99 | 0.979 | 0.8 | 0.825 | nan | - | - | - | - | |
| all | 0.836 | 0.997 | 0.974 | 0.8 | 0.768 | 0.827 | 0.863 | 0.8 | 0.788 | 0.855 | |
| dspnet | victim | 0.811 | 1 | 0.967 | nan | 0.648 | 0.814 | - | - | - | - |
| plane | 0.852 | 1 | 0.979 | nan | 0.852 | 0.783 | - | - | - | - | |
| boat | 0.816 | 0.989 | 0.989 | 0.9 | 0.817 | nan | - | - | - | - | |
| all | 0.826 | 0.996 | 0.978 | 0.9 | 0.772 | 0.799 | 0.855 | 0.9 | 0.796 | 0.82 | |
| b2ub- g25 | victim | 0.793 | 1 | 0.956 | nan | 0.585 | 0.798 | - | - | - | - |
| plane | 0.832 | 1 | 0.978 | nan | 0.832 | 0.725 | - | - | - | - | |
| boat | 0.774 | 0.99 | 0.979 | 0.7 | 0.775 | nan | - | - | - | - | |
| all | 0.799 | 0.997 | 0.971 | 0.7 | 0.731 | 0.762 | 0.829 | 0.7 | 0.752 | 0.787 | |
| b2ub- g5-50 | victim | 0.818 | 1 | 0.968 | nan | 0.715 | 0.82 | - | - | - | - |
| plane | 0.865 | 1 | 0.99 | nan | 0.866 | 0.723 | - | - | - | - | |
| boat | 0.816 | 0.99 | 0.99 | 0.7 | 0.816 | nan | - | - | - | - | |
| all | 0.833 | 0.997 | 0.983 | 0.7 | 0.799 | 0.771 | 0.861 | 0.7 | 0.823 | 0.824 | |
| b2ub- p30 | victim | 0.816 | 1 | 0.954 | nan | 0.628 | 0.821 | - | - | - | - |
| plane | 0.866 | 1 | 0.98 | nan | 0.866 | 0.717 | - | - | - | - | |
| boat | 0.837 | 0.99 | 0.99 | 0.9 | 0.837 | nan | - | - | - | - | |
| all | 0.839 | 0.997 | 0.975 | 0.9 | 0.777 | 0.769 | 0.865 | 0.9 | 0.799 | 0.799 | |
| b2ub- p5-50 | victim | 0.843 | 1 | 0.975 | nan | 0.699 | 0.846 | - | - | - | - |
| plane | 0.859 | 1 | 0.98 | nan | 0.86 | 0.8 | - | - | - | - | |
| boat | 0.847 | 0.99 | 0.99 | 0.9 | 0.847 | nan | - | - | - | - | |
| all | 0.85 | 0.997 | 0.982 | 0.9 | 0.802 | 0.823 | 0.875 | 0.9 | 0.82 | 0.834 | |
| nei- g25 | victim | 0.812 | 1 | 0.953 | nan | 0.687 | 0.815 | - | - | - | - |
| plane | 0.833 | 1 | 0.98 | nan | 0.834 | 0.717 | - | - | - | - | |
| boat | 0.793 | 0.99 | 0.99 | 0.8 | 0.793 | nan | - | - | - | - | |
| all | 0.812 | 0.997 | 0.974 | 0.8 | 0.771 | 0.766 | 0.843 | 0.8 | 0.797 | 0.795 | |
| nei- g5-50 | ship | 0.819 | 1 | 0.974 | nan | 0.713 | 0.822 | - | - | - | - |
| aircraft | 0.85 | 1 | 0.979 | nan | 0.852 | 0.723 | - | - | - | - | |
| human | 0.824 | 0.99 | 0.98 | 0.8 | 0.824 | nan | - | - | - | - | |
| all | 0.831 | 0.997 | 0.978 | 0.8 | 0.796 | 0.773 | 0.857 | 0.8 | 0.819 | 0.823 | |
| nei- p30 | victim | 0.821 | 1 | 0.977 | nan | 0.666 | 0.826 | - | - | - | - |
| plane | 0.863 | 1 | 0.98 | nan | 0.864 | 0.7 | - | - | - | - | |
| boat | 0.83 | 0.989 | 0.989 | 0.8 | 0.831 | nan | - | - | - | - | |
| all | 0.838 | 0.996 | 0.982 | 0.8 | 0.787 | 0.763 | 0.866 | 0.8 | 0.808 | 0.825 | |
| nei- p5-50 | victim | 0.823 | 1 | 0.965 | nan | 0.718 | 0.825 | - | - | - | - |
| plane | 0.856 | 1 | 0.98 | nan | 0.857 | 0.75 | - | - | - | - | |
| boat | 0.821 | 0.989 | 0.989 | 0.8 | 0.822 | nan | - | - | - | - | |
| all | 0.833 | 0.996 | 0.978 | 0.8 | 0.799 | 0.788 | 0.859 | 0.8 | 0.825 | 0.824 | |
| Denoising | ssd_mobilenetv2 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | human body | 0.475 | 0.972 | 0.345 | 0 | 0.471 | 0.552 | - | - | - | - |
| ball | 0.436 | 0.934 | 0.313 | nan | 0.437 | 0.214 | - | - | - | - | |
| circle cage | 0.415 | 0.898 | 0.354 | nan | 0.412 | 0.474 | - | - | - | - | |
| square cage | 0.428 | 0.922 | 0.317 | 0 | 0.431 | nan | - | - | - | - | |
| tyre | 0.373 | 0.863 | 0.257 | 0 | 0.348 | 0.564 | - | - | - | - | |
| cube | 0.474 | 0.934 | 0.404 | 0.169 | 0.478 | 0.496 | - | - | - | - | |
| cylinder | 0.313 | 0.816 | 0.12 | 0.583 | 0.309 | nan | - | - | - | - | |
| metal bucket | 0.44 | 0.89 | 0.275 | nan | 0.443 | 0.363 | - | - | - | - | |
| plane | 0.562 | 0.974 | 0.586 | nan | 0.535 | 0.605 | - | - | - | - | |
| rov | 0.303 | 0.855 | 0.119 | 0 | 0.305 | nan | - | - | - | - | |
| all | 0.422 | 0.906 | 0.309 | 0.125 | 0.417 | 0.467 | 0.516 | 0.13 | 0.512 | 0.517 | |
| dspnet | human body | 0.484 | 0.964 | 0.411 | 0 | 0.481 | 0.564 | - | - | - | - |
| ball | 0.452 | 0.95 | 0.321 | nan | 0.451 | 0.327 | - | - | - | - | |
| circle cage | 0.433 | 0.858 | 0.341 | nan | 0.429 | 0.528 | - | - | - | - | |
| square cage | 0.447 | 0.933 | 0.356 | 0 | 0.45 | nan | - | - | - | - | |
| tyre | 0.372 | 0.896 | 0.234 | 0 | 0.355 | 0.507 | - | - | - | - | |
| cube | 0.462 | 0.917 | 0.38 | 0.348 | 0.462 | 0.524 | - | - | - | - | |
| cylinder | 0.329 | 0.795 | 0.219 | 0.288 | 0.331 | nan | - | - | - | - | |
| metal bucket | 0.456 | 0.906 | 0.37 | nan | 0.455 | 0.451 | - | - | - | - | |
| plane | 0.547 | 0.975 | 0.488 | nan | 0.502 | 0.622 | - | - | - | - | |
| rov | 0.32 | 0.841 | 0.108 | 0.1 | 0.323 | nan | - | - | - | - | |
| all | 0.43 | 0.904 | 0.323 | 0.123 | 0.424 | 0.503 | 0.523 | 0.136 | 0.518 | 0.552 | |
| b2ub- g25 | human body | 0.453 | 0.957 | 0.355 | 0 | 0.449 | 0.549 | - | - | - | - |
| ball | 0.425 | 0.931 | 0.243 | nan | 0.425 | 0.277 | - | - | - | - | |
| circle cage | 0.404 | 0.876 | 0.243 | nan | 0.399 | 0.525 | - | - | - | - | |
| square cage | 0.404 | 0.906 | 0.282 | 0 | 0.408 | nan | - | - | - | - | |
| tyre | 0.333 | 0.833 | 0.192 | 0 | 0.314 | 0.499 | - | - | - | - | |
| cube | 0.439 | 0.931 | 0.281 | 0.283 | 0.439 | 0.543 | - | - | - | - | |
| cylinder | 0.284 | 0.767 | 0.161 | 0.222 | 0.288 | nan | - | - | - | - | |
| metal bucket | 0.391 | 0.832 | 0.248 | nan | 0.393 | 0.376 | - | - | - | - | |
| plane | 0.551 | 0.987 | 0.545 | nan | 0.514 | 0.615 | - | - | - | - | |
| rov | 0.296 | 0.832 | 0.098 | 0 | 0.297 | nan | - | - | - | - | |
| all | 0.398 | 0.885 | 0.265 | 0.084 | 0.393 | 0.483 | 0.494 | 0.091 | 0.49 | 0.514 | |
| b2ub- g5-50 | human body | 0.474 | 0.962 | 0.362 | 0 | 0.471 | 0.55 | - | - | - | - |
| ball | 0.446 | 0.938 | 0.325 | nan | 0.446 | 0.252 | - | - | - | - | |
| circle cage | 0.406 | 0.868 | 0.283 | nan | 0.399 | 0.546 | - | - | - | - | |
| square cage | 0.45 | 0.959 | 0.276 | 0 | 0.453 | nan | - | - | - | - | |
| tyre | 0.363 | 0.893 | 0.217 | 0 | 0.352 | 0.488 | - | - | - | - | |
| cube | 0.471 | 0.938 | 0.374 | 0.342 | 0.472 | 0.478 | - | - | - | - | |
| cylinder | 0.317 | 0.819 | 0.182 | 0.533 | 0.315 | nan | - | - | - | - | |
| metal bucket | 0.469 | 0.903 | 0.425 | nan | 0.467 | 0.564 | - | - | - | - | |
| plane | 0.57 | 0.999 | 0.524 | nan | 0.539 | 0.622 | - | - | - | - | |
| rov | 0.323 | 0.918 | 0.121 | 0.2 | 0.323 | nan | - | - | - | - | |
| all | 0.429 | 0.92 | 0.309 | 0.179 | 0.424 | 0.5 | 0.523 | 0.179 | 0.516 | 0.554 | |
| b2ub- p30 | human body | 0.469 | 0.974 | 0.312 | 0 | 0.464 | 0.579 | - | - | - | - |
| ball | 0.436 | 0.938 | 0.293 | nan | 0.437 | 0.277 | - | - | - | - | |
| circle cage | 0.419 | 0.888 | 0.317 | nan | 0.414 | 0.524 | - | - | - | - | |
| square cage | 0.426 | 0.951 | 0.248 | 0 | 0.428 | nan | - | - | - | - | |
| tyre | 0.373 | 0.857 | 0.254 | 0 | 0.357 | 0.506 | - | - | - | - | |
| cube | 0.462 | 0.923 | 0.402 | 0.223 | 0.465 | 0.498 | - | - | - | - | |
| cylinder | 0.301 | 0.801 | 0.153 | 0.417 | 0.3 | nan | - | - | - | - | |
| metal bucket | 0.41 | 0.894 | 0.284 | nan | 0.411 | 0.403 | - | - | - | - | |
| plane | 0.549 | 0.984 | 0.528 | nan | 0.51 | 0.61 | - | - | - | - | |
| rov | 0.304 | 0.832 | 0.135 | 0 | 0.306 | nan | - | - | - | - | |
| all | 0.415 | 0.904 | 0.292 | 0.107 | 0.409 | 0.485 | 0.512 | 0.121 | 0.506 | 0.534 | |
| b2ub- p5-50 | human body | 0.456 | 0.955 | 0.308 | 0 | 0.451 | 0.553 | - | - | - | - |
| ball | 0.43 | 0.935 | 0.301 | nan | 0.43 | 0.218 | - | - | - | - | |
| circle cage | 0.43 | 0.882 | 0.363 | nan | 0.429 | 0.475 | - | - | - | - | |
| square cage | 0.422 | 0.936 | 0.27 | 0 | 0.426 | nan | - | - | - | - | |
| tyre | 0.372 | 0.889 | 0.252 | 0 | 0.35 | 0.542 | - | - | - | - | |
| cube | 0.468 | 0.941 | 0.379 | 0.362 | 0.468 | 0.493 | - | - | - | - | |
| cylinder | 0.312 | 0.791 | 0.143 | 0.32 | 0.314 | nan | - | - | - | - | |
| metal bucket | 0.44 | 0.917 | 0.271 | nan | 0.447 | 0.307 | - | - | - | - | |
| plane | 0.568 | 0.979 | 0.645 | nan | 0.547 | 0.597 | - | - | - | - | |
| rov | 0.294 | 0.84 | 0.115 | 0 | 0.296 | nan | - | - | - | - | |
| all | 0.419 | 0.906 | 0.305 | 0.114 | 0.416 | 0.455 | 0.513 | 0.121 | 0.512 | 0.5 | |
| nei- g25 | human body | 0.418 | 0.944 | 0.275 | 0 | 0.408 | 0.592 | - | - | - | - |
| ball | 0.414 | 0.916 | 0.243 | nan | 0.414 | 0.202 | - | - | - | - | |
| circle cage | 0.381 | 0.827 | 0.315 | nan | 0.38 | 0.407 | - | - | - | - | |
| square cage | 0.413 | 0.932 | 0.247 | 0 | 0.416 | nan | - | - | - | - | |
| tyre | 0.328 | 0.814 | 0.197 | 0 | 0.307 | 0.474 | - | - | - | - | |
| cube | 0.439 | 0.919 | 0.31 | 0.23 | 0.44 | 0.509 | - | - | - | - | |
| cylinder | 0.315 | 0.807 | 0.13 | 0.5 | 0.312 | nan | - | - | - | - | |
| metal bucket | 0.444 | 0.906 | 0.32 | nan | 0.449 | 0.378 | - | - | - | - | |
| plane | 0.572 | 0.979 | 0.581 | nan | 0.534 | 0.633 | - | - | - | - | |
| rov | 0.29 | 0.853 | 0.099 | 0 | 0.292 | nan | - | - | - | - | |
| all | 0.401 | 0.89 | 0.272 | 0.122 | 0.395 | 0.456 | 0.496 | 0.129 | 0.491 | 0.491 | |
| nei- g5-50 | human body | 0.457 | 0.955 | 0.318 | 0 | 0.454 | 0.521 | - | - | - | - |
| ball | 0.417 | 0.926 | 0.253 | nan | 0.417 | 0.235 | - | - | - | - | |
| circle cage | 0.419 | 0.895 | 0.295 | nan | 0.419 | 0.48 | - | - | - | - | |
| square cage | 0.434 | 0.935 | 0.267 | 0 | 0.435 | nan | - | - | - | - | |
| tyre | 0.318 | 0.844 | 0.122 | 0 | 0.313 | 0.372 | - | - | - | - | |
| cube | 0.439 | 0.915 | 0.339 | 0.244 | 0.441 | 0.508 | - | - | - | - | |
| cylinder | 0.306 | 0.776 | 0.174 | 0.466 | 0.304 | nan | - | - | - | - | |
| metal bucket | 0.418 | 0.855 | 0.265 | nan | 0.425 | 0.326 | - | - | - | - | |
| plane | 0.537 | 0.997 | 0.5 | nan | 0.515 | 0.579 | - | - | - | - | |
| rov | 0.27 | 0.8 | 0.075 | 0.2 | 0.271 | nan | - | - | - | - | |
| all | 0.401 | 0.89 | 0.261 | 0.152 | 0.399 | 0.432 | 0.495 | 0.16 | 0.494 | 0.471 | |
| nei- p30 | human body | 0.506 | 0.976 | 0.42 | 0 | 0.507 | 0.544 | - | - | - | - |
| ball | 0.535 | 0.968 | 0.521 | nan | 0.535 | 0.303 | - | - | - | - | |
| circle cage | 0.514 | 0.935 | 0.498 | nan | 0.518 | 0.524 | - | - | - | - | |
| square cage | 0.488 | 0.946 | 0.385 | 0 | 0.491 | nan | - | - | - | - | |
| tyre | 0.5 | 0.92 | 0.479 | 0 | 0.492 | 0.586 | - | - | - | - | |
| cube | 0.501 | 0.942 | 0.461 | 0.307 | 0.502 | 0.549 | - | - | - | - | |
| cylinder | 0.352 | 0.785 | 0.207 | 0.655 | 0.348 | nan | - | - | - | - | |
| metal bucket | 0.491 | 0.915 | 0.47 | nan | 0.5 | 0.295 | - | - | - | - | |
| plane | 0.573 | 0.997 | 0.604 | nan | 0.549 | 0.614 | - | - | - | - | |
| rov | 0.429 | 0.977 | 0.243 | 0 | 0.433 | nan | - | - | - | - | |
| all | 0.489 | 0.936 | 0.429 | 0.16 | 0.488 | 0.488 | 0.572 | 0.171 | 0.571 | 0.527 | |
| nei- p5-50 | human body | 0.491 | 0.961 | 0.413 | 0 | 0.487 | 0.578 | - | - | - | - |
| ball | 0.461 | 0.948 | 0.35 | nan | 0.46 | 0.269 | - | - | - | - | |
| circle cage | 0.438 | 0.904 | 0.366 | nan | 0.437 | 0.479 | - | - | - | - | |
| square cage | 0.439 | 0.96 | 0.267 | 0 | 0.441 | nan | - | - | - | - | |
| tyre | 0.391 | 0.906 | 0.232 | 0 | 0.378 | 0.518 | - | - | - | - | |
| cube | 0.471 | 0.945 | 0.407 | 0.343 | 0.47 | 0.559 | - | - | - | - | |
| cylinder | 0.295 | 0.808 | 0.108 | 0.294 | 0.296 | nan | - | - | - | - | |
| metal bucket | 0.438 | 0.905 | 0.358 | nan | 0.436 | 0.467 | - | - | - | - | |
| plane | 0.56 | 0.974 | 0.576 | nan | 0.52 | 0.62 | - | - | - | - | |
| rov | 0.336 | 0.911 | 0.144 | 0 | 0.338 | nan | - | - | - | - | |
| all | 0.432 | 0.922 | 0.322 | 0.106 | 0.426 | 0.499 | 0.524 | 0.123 | 0.519 | 0.556 |
| Denoising | Faster r-cnn | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | human body | 0.527 | 0.977 | 0.519 | 0 | 0.531 | 0.547 | - | - | - | - |
| ball | 0.55 | 0.959 | 0.558 | nan | 0.551 | 0.252 | - | - | - | - | |
| circle cage | 0.527 | 0.904 | 0.552 | nan | 0.52 | 0.613 | - | - | - | - | |
| square cage | 0.528 | 0.962 | 0.502 | 0 | 0.53 | nan | - | - | - | - | |
| tyre | 0.54 | 0.957 | 0.532 | 0 | 0.53 | 0.638 | - | - | - | - | |
| cube | 0.537 | 0.955 | 0.574 | 0.343 | 0.538 | 0.578 | - | - | - | - | |
| cylinder | 0.446 | 0.895 | 0.411 | 0.734 | 0.437 | nan | - | - | - | - | |
| metal bucket | 0.523 | 0.932 | 0.509 | nan | 0.533 | 0.37 | - | - | - | - | |
| plane | 0.651 | 0.99 | 0.773 | nan | 0.606 | 0.702 | - | - | - | - | |
| rov | 0.504 | 0.968 | 0.398 | 0 | 0.507 | nan | - | - | - | - | |
| all | 0.533 | 0.95 | 0.533 | 0.179 | 0.529 | 0.528 | 0.605 | 0.182 | 0.603 | 0.563 | |
| dspnet | human body | 0.531 | 0.978 | 0.501 | 0 | 0.536 | 0.549 | - | - | - | - |
| ball | 0.549 | 0.955 | 0.555 | nan | 0.549 | 0.303 | - | - | - | - | |
| circle cage | 0.544 | 0.923 | 0.595 | nan | 0.544 | 0.574 | - | - | - | - | |
| square cage | 0.51 | 0.959 | 0.446 | 0 | 0.512 | nan | - | - | - | - | |
| tyre | 0.551 | 0.96 | 0.607 | 0 | 0.544 | 0.632 | - | - | - | - | |
| cube | 0.532 | 0.954 | 0.542 | 0.343 | 0.533 | 0.561 | - | - | - | - | |
| cylinder | 0.453 | 0.884 | 0.387 | 0.717 | 0.446 | nan | - | - | - | - | |
| metal bucket | 0.515 | 0.927 | 0.5 | nan | 0.528 | 0.334 | - | - | - | - | |
| plane | 0.638 | 0.999 | 0.68 | nan | 0.612 | 0.674 | - | - | - | - | |
| rov | 0.53 | 0.986 | 0.443 | 0.1 | 0.533 | nan | - | - | - | - | |
| all | 0.535 | 0.952 | 0.526 | 0.193 | 0.534 | 0.518 | 0.606 | 0.196 | 0.604 | 0.554 | |
| b2ub- g25 | human body | 0.484 | 0.946 | 0.421 | 0 | 0.486 | 0.499 | - | - | - | - |
| ball | 0.523 | 0.953 | 0.503 | nan | 0.523 | 0.252 | - | - | - | - | |
| circle cage | 0.505 | 0.891 | 0.466 | nan | 0.497 | 0.59 | - | - | - | - | |
| square cage | 0.474 | 0.929 | 0.377 | 0 | 0.477 | nan | - | - | - | - | |
| tyre | 0.466 | 0.88 | 0.429 | 0 | 0.455 | 0.566 | - | - | - | - | |
| cube | 0.512 | 0.943 | 0.482 | 0.284 | 0.514 | 0.598 | - | - | - | - | |
| cylinder | 0.42 | 0.841 | 0.338 | 0.634 | 0.413 | nan | - | - | - | - | |
| metal bucket | 0.471 | 0.917 | 0.432 | nan | 0.488 | 0.127 | - | - | - | - | |
| plane | 0.643 | 0.998 | 0.745 | nan | 0.606 | 0.691 | - | - | - | - | |
| rov | 0.508 | 0.963 | 0.438 | 0.3 | 0.507 | nan | - | - | - | - | |
| all | 0.5 | 0.926 | 0.463 | 0.203 | 0.497 | 0.475 | 0.577 | 0.236 | 0.575 | 0.506 | |
| b2ub- g5-50 | human body | 0.537 | 0.968 | 0.525 | 0 | 0.535 | 0.591 | - | - | - | - |
| ball | 0.555 | 0.96 | 0.571 | nan | 0.555 | 0.303 | - | - | - | - | |
| circle cage | 0.535 | 0.923 | 0.583 | nan | 0.537 | 0.53 | - | - | - | - | |
| square cage | 0.517 | 0.953 | 0.473 | 0 | 0.519 | nan | - | - | - | - | |
| tyre | 0.542 | 0.945 | 0.57 | 0 | 0.535 | 0.628 | - | - | - | - | |
| cube | 0.538 | 0.958 | 0.54 | 0.322 | 0.542 | 0.547 | - | - | - | - | |
| cylinder | 0.469 | 0.892 | 0.426 | 0.734 | 0.463 | nan | - | - | - | - | |
| metal bucket | 0.532 | 0.938 | 0.553 | nan | 0.537 | 0.452 | - | - | - | - | |
| plane | 0.635 | 0.99 | 0.741 | nan | 0.603 | 0.678 | - | - | - | - | |
| rov | 0.52 | 0.987 | 0.429 | 0 | 0.525 | nan | - | - | - | - | |
| all | 0.538 | 0.951 | 0.541 | 0.176 | 0.535 | 0.533 | 0.607 | 0.176 | 0.605 | 0.569 | |
| b2ub- p30 | human body | 0.528 | 0.975 | 0.512 | 0 | 0.528 | 0.574 | - | - | - | - |
| ball | 0.555 | 0.955 | 0.593 | nan | 0.555 | 0.303 | - | - | - | - | |
| circle cage | 0.53 | 0.914 | 0.537 | nan | 0.532 | 0.538 | - | - | - | - | |
| square cage | 0.519 | 0.962 | 0.501 | 0 | 0.521 | nan | - | - | - | - | |
| tyre | 0.541 | 0.928 | 0.579 | 0 | 0.531 | 0.635 | - | - | - | - | |
| cube | 0.539 | 0.942 | 0.566 | 0.302 | 0.541 | 0.624 | - | - | - | - | |
| cylinder | 0.443 | 0.885 | 0.349 | 0.7 | 0.44 | nan | - | - | - | - | |
| metal bucket | 0.494 | 0.929 | 0.478 | nan | 0.498 | 0.427 | - | - | - | - | |
| plane | 0.629 | 0.999 | 0.739 | nan | 0.605 | 0.662 | - | - | - | - | |
| rov | 0.535 | 0.967 | 0.522 | 0.2 | 0.538 | nan | - | - | - | - | |
| all | 0.531 | 0.946 | 0.538 | 0.2 | 0.529 | 0.538 | 0.605 | 0.206 | 0.603 | 0.574 | |
| b2ub- p5-50 | human body | 0.528 | 0.977 | 0.479 | 0 | 0.528 | 0.576 | - | - | - | - |
| ball | 0.554 | 0.967 | 0.593 | nan | 0.555 | 0.252 | - | - | - | - | |
| circle cage | 0.548 | 0.932 | 0.598 | nan | 0.548 | 0.55 | - | - | - | - | |
| square cage | 0.512 | 0.952 | 0.49 | 0 | 0.514 | nan | - | - | - | - | |
| tyre | 0.543 | 0.949 | 0.589 | 0 | 0.54 | 0.593 | - | - | - | - | |
| cube | 0.542 | 0.957 | 0.551 | 0.362 | 0.542 | 0.598 | - | - | - | - | |
| cylinder | 0.457 | 0.897 | 0.438 | 0.701 | 0.451 | nan | - | - | - | - | |
| metal bucket | 0.492 | 0.905 | 0.454 | nan | 0.5 | 0.33 | - | - | - | - | |
| plane | 0.634 | 0.989 | 0.741 | nan | 0.605 | 0.678 | - | - | - | - | |
| rov | 0.514 | 0.966 | 0.408 | 0 | 0.516 | nan | - | - | - | - | |
| all | 0.532 | 0.949 | 0.534 | 0.177 | 0.53 | 0.511 | 0.606 | 0.177 | 0.605 | 0.553 | |
| nei- g25 | human body | 0.53 | 0.982 | 0.52 | 0 | 0.526 | 0.583 | - | - | - | - |
| ball | 0.541 | 0.957 | 0.558 | nan | 0.541 | 0.252 | - | - | - | - | |
| circle cage | 0.517 | 0.913 | 0.499 | nan | 0.519 | 0.525 | - | - | - | - | |
| square cage | 0.505 | 0.955 | 0.468 | 0 | 0.508 | nan | - | - | - | - | |
| tyre | 0.507 | 0.926 | 0.508 | 0 | 0.495 | 0.61 | - | - | - | - | |
| cube | 0.527 | 0.945 | 0.536 | 0.263 | 0.532 | 0.537 | - | - | - | - | |
| cylinder | 0.45 | 0.865 | 0.407 | 0.672 | 0.445 | nan | - | - | - | - | |
| metal bucket | 0.513 | 0.915 | 0.538 | nan | 0.522 | 0.326 | - | - | - | - | |
| plane | 0.654 | 0.99 | 0.777 | nan | 0.617 | 0.7 | - | - | - | - | |
| rov | 0.512 | 0.976 | 0.442 | 0.3 | 0.514 | nan | - | - | - | - | |
| all | 0.526 | 0.943 | 0.525 | 0.206 | 0.522 | 0.505 | 0.595 | 0.21 | 0.593 | 0.535 | |
| nei- g5-50 | human body | 0.524 | 0.979 | 0.482 | 0 | 0.525 | 0.56 | - | - | - | - |
| ball | 0.556 | 0.966 | 0.588 | nan | 0.556 | 0.303 | - | - | - | - | |
| circle cage | 0.526 | 0.931 | 0.535 | nan | 0.526 | 0.551 | - | - | - | - | |
| square cage | 0.506 | 0.954 | 0.437 | 0 | 0.51 | nan | - | - | - | - | |
| tyre | 0.548 | 0.948 | 0.585 | 0 | 0.548 | 0.59 | - | - | - | - | |
| cube | 0.534 | 0.95 | 0.537 | 0.343 | 0.535 | 0.581 | - | - | - | - | |
| cylinder | 0.454 | 0.895 | 0.423 | 0.723 | 0.449 | nan | - | - | - | - | |
| metal bucket | 0.52 | 0.928 | 0.481 | nan | 0.524 | 0.478 | - | - | - | - | |
| plane | 0.649 | 0.98 | 0.709 | nan | 0.608 | 0.699 | - | - | - | - | |
| rov | 0.521 | 0.983 | 0.446 | 0.3 | 0.521 | nan | - | - | - | - | |
| all | 0.534 | 0.951 | 0.522 | 0.228 | 0.53 | 0.537 | 0.606 | 0.246 | 0.603 | 0.574 | |
| nei- p30 | human body | 0.569 | 0.977 | 0.623 | 0 | 0.568 | 0.618 | - | - | - | - |
| ball | 0.587 | 0.973 | 0.663 | nan | 0.587 | 0.353 | - | - | - | - | |
| circle cage | 0.532 | 0.912 | 0.574 | nan | 0.531 | 0.57 | - | - | - | - | |
| square cage | 0.513 | 0.953 | 0.496 | 0 | 0.515 | nan | - | - | - | - | |
| tyre | 0.553 | 0.946 | 0.601 | 0 | 0.548 | 0.632 | - | - | - | - | |
| cube | 0.563 | 0.96 | 0.607 | 0.423 | 0.563 | 0.642 | - | - | - | - | |
| cylinder | 0.476 | 0.941 | 0.366 | 0.634 | 0.472 | nan | - | - | - | - | |
| metal bucket | 0.509 | 0.931 | 0.516 | nan | 0.522 | 0.251 | - | - | - | - | |
| plane | 0.658 | 1 | 0.742 | nan | 0.631 | 0.69 | - | - | - | - | |
| rov | 0.525 | 0.999 | 0.456 | 0.2 | 0.527 | nan | - | - | - | - | |
| all | 0.549 | 0.959 | 0.959 | 0.209 | 0.546 | 0.537 | 0.61 | 0.209 | 0.609 | 0.565 | |
| nei- p5-50 | human body | 0.537 | 0.978 | 0.538 | 0 | 0.536 | 0.581 | - | - | - | - |
| ball | 0.554 | 0.969 | 0.577 | nan | 0.553 | 0.303 | - | - | - | - | |
| circle cage | 0.536 | 0.93 | 0.548 | nan | 0.538 | 0.535 | - | - | - | - | |
| square cage | 0.517 | 0.95 | 0.467 | 0 | 0.519 | nan | - | - | - | - | |
| tyre | 0.551 | 0.928 | 0.615 | 0 | 0.543 | 0.631 | - | - | - | - | |
| cube | 0.541 | 0.957 | 0.544 | 0.329 | 0.542 | 0.545 | - | - | - | - | |
| cylinder | 0.457 | 0.873 | 0.418 | 0.634 | 0.453 | nan | - | - | - | - | |
| metal bucket | 0.511 | 0.933 | 0.451 | nan | 0.523 | 0.326 | - | - | - | - | |
| plane | 0.653 | 0.999 | 0.708 | nan | 0.616 | 0.701 | - | - | - | - | |
| rov | 0.512 | 0.976 | 0.44 | 0.1 | 0.513 | nan | - | - | - | - | |
| all | 0.537 | 0.949 | 0.531 | 0.177 | 0.534 | 0.517 | 0.608 | 0.184 | 0.607 | 0.549 |
| Denoising | ssd300 | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | human body | 0.481 | 0.966 | 0.402 | 0 | 0.482 | 0.485 | - | - | - | - |
| ball | 0.517 | 0.963 | 0.497 | nan | 0.516 | 0.349 | - | - | - | - | |
| circle cage | 0.496 | 0.92 | 0.504 | nan | 0.486 | 0.582 | - | - | - | - | |
| square cage | 0.473 | 0.942 | 0.361 | 0 | 0.477 | nan | - | - | - | - | |
| tyre | 0.463 | 0.908 | 0.425 | 0 | 0.451 | 0.574 | - | - | - | - | |
| cube | 0.499 | 0.939 | 0.448 | 0.363 | 0.499 | 0.587 | - | - | - | - | |
| cylinder | 0.355 | 0.823 | 0.19 | 0.666 | 0.352 | nan | - | - | - | - | |
| metal bucket | 0.484 | 0.926 | 0.43 | nan | 0.492 | 0.337 | - | - | - | - | |
| plane | 0.582 | 0.996 | 0.643 | nan | 0.542 | 0.644 | - | - | - | - | |
| rov | 0.442 | 0.976 | 0.321 | 0 | 0.447 | nan | - | - | - | - | |
| all | 0.479 | 0.936 | 0.422 | 0.172 | 0.474 | 0.508 | 0.561 | 0.171 | 0.558 | 0.586 | |
| dspnet | human body | 0.476 | 0.957 | 0.389 | 0 | 0.478 | 0.503 | - | - | - | - |
| ball | 0.512 | 0.962 | 0.443 | nan | 0.511 | 0.435 | - | - | - | - | |
| circle cage | 0.494 | 0.931 | 0.434 | nan | 0.486 | 0.578 | - | - | - | - | |
| square cage | 0.446 | 0.912 | 0.273 | 0 | 0.45 | nan | - | - | - | - | |
| tyre | 0.457 | 0.864 | 0.424 | 0 | 0.443 | 0.582 | - | - | - | - | |
| cube | 0.487 | 0.928 | 0.458 | 0.343 | 0.486 | 0.586 | - | - | - | - | |
| cylinder | 0.379 | 0.838 | 0.24 | 0.623 | 0.375 | nan | - | - | - | - | |
| metal bucket | 0.5 | 0.924 | 0.494 | nan | 0.51 | 0.327 | - | - | - | - | |
| plane | 0.578 | 0.996 | 0.557 | nan | 0.55 | 0.617 | - | - | - | - | |
| rov | 0.426 | 0.946 | 0.257 | 0 | 0.437 | nan | - | - | - | - | |
| all | 0.476 | 0.926 | 0.397 | 0.161 | 0.473 | 0.518 | 0.558 | 0.182 | 0.555 | 0.569 | |
| b2ub- g25 | human body | 0.454 | 0.974 | 0.312 | 0 | 0.455 | 0.498 | - | - | - | - |
| ball | 0.506 | 0.956 | 0.425 | nan | 0.507 | 0.384 | - | - | - | - | |
| circle cage | 0.477 | 0.92 | 0.468 | nan | 0.469 | 0.546 | - | - | - | - | |
| square cage | 0.458 | 0.929 | 0.351 | 0 | 0.461 | nan | - | - | - | - | |
| tyre | 0.445 | 0.89 | 0.412 | 0 | 0.43 | 0.58 | - | - | - | - | |
| cube | 0.486 | 0.939 | 0.412 | 0.261 | 0.484 | 0.596 | - | - | - | - | |
| cylinder | 0.357 | 0.847 | 0.194 | 0.434 | 0.356 | nan | - | - | - | - | |
| metal bucket | 0.458 | 0.922 | 0.316 | nan | 0.47 | 0.201 | - | - | - | - | |
| plane | 0.577 | 0.977 | 0.638 | nan | 0.559 | 0.606 | - | - | - | - | |
| rov | 0.422 | 0.954 | 0.219 | 0 | 0.426 | nan | - | - | - | - | |
| all | 0.464 | 0.931 | 0.375 | 0.116 | 0.462 | 0.487 | 0.546 | 0.116 | 0.545 | 0.538 | |
| b2ub- g5-50 | human body | 0.508 | 0.971 | 0.438 | 0 | 0.51 | 0.524 | - | - | - | - |
| ball | 0.534 | 0.967 | 0.519 | nan | 0.534 | 0.372 | - | - | - | - | |
| circle cage | 0.495 | 0.934 | 0.408 | nan | 0.491 | 0.543 | - | - | - | - | |
| square cage | 0.49 | 0.942 | 0.411 | 0 | 0.494 | nan | - | - | - | - | |
| tyre | 0.48 | 0.913 | 0.464 | 0 | 0.469 | 0.59 | - | - | - | - | |
| cube | 0.504 | 0.942 | 0.481 | 0.322 | 0.505 | 0.588 | - | - | - | - | |
| cylinder | 0.364 | 0.839 | 0.201 | 0.655 | 0.359 | nan | - | - | - | - | |
| metal bucket | 0.492 | 0.928 | 0.487 | nan | 0.503 | 0.271 | - | - | - | - | |
| plane | 0.58 | 0.979 | 0.612 | nan | 0.549 | 0.626 | - | - | - | - | |
| rov | 0.428 | 0.958 | 0.262 | 0 | 0.435 | nan | - | - | - | - | |
| all | 0.488 | 0.937 | 0.428 | 0.163 | 0.485 | 0.502 | 0.57 | 0.168 | 0.568 | 0.589 | |
| b2ub- p30 | human body | 0.501 | 0.97 | 0.456 | 0 | 0.504 | 0.522 | - | - | - | - |
| ball | 0.531 | 0.966 | 0.505 | nan | 0.531 | 0.382 | - | - | - | - | |
| circle cage | 0.519 | 0.927 | 0.564 | nan | 0.514 | 0.57 | - | - | - | - | |
| square cage | 0.495 | 0.948 | 0.452 | 0 | 0.497 | nan | - | - | - | - | |
| tyre | 0.476 | 0.9 | 0.46 | 0 | 0.466 | 0.572 | - | - | - | - | |
| cube | 0.505 | 0.936 | 0.484 | 0.236 | 0.506 | 0.599 | - | - | - | - | |
| cylinder | 0.375 | 0.803 | 0.253 | 0.578 | 0.371 | nan | - | - | - | - | |
| metal bucket | 0.501 | 0.916 | 0.492 | nan | 0.508 | 0.35 | - | - | - | - | |
| plane | 0.585 | 0.976 | 0.62 | nan | 0.547 | 0.647 | - | - | - | - | |
| rov | 0.43 | 0.96 | 0.251 | 0 | 0.434 | nan | - | - | - | - | |
| all | 0.492 | 0.93 | 0.454 | 0.136 | 0.488 | 0.521 | 0.575 | 0.156 | 0.573 | 0.596 | |
| b2ub- p5-50 | human body | 0.509 | 0.974 | 0.431 | 0 | 0.512 | 0.522 | - | - | - | - |
| ball | 0.536 | 0.973 | 0.509 | nan | 0.535 | 0.468 | - | - | - | - | |
| circle cage | 0.512 | 0.924 | 0.515 | nan | 0.508 | 0.555 | - | - | - | - | |
| square cage | 0.487 | 0.931 | 0.4 | 0 | 0.49 | nan | - | - | - | - | |
| tyre | 0.483 | 0.917 | 0.404 | 0 | 0.472 | 0.593 | - | - | - | - | |
| cube | 0.497 | 0.944 | 0.442 | 0.323 | 0.497 | 0.57 | - | - | - | - | |
| cylinder | 0.383 | 0.837 | 0.225 | 0.7 | 0.378 | nan | - | - | - | - | |
| metal bucket | 0.501 | 0.929 | 0.518 | nan | 0.51 | 0.276 | - | - | - | - | |
| plane | 0.573 | 0.996 | 0.59 | nan | 0.559 | 0.598 | - | - | - | - | |
| rov | 0.44 | 0.97 | 0.289 | 0 | 0.444 | nan | - | - | - | - | |
| all | 0.492 | 0.94 | 0.432 | 0.17 | 0.49 | 0.512 | 0.572 | 0.17 | 0.571 | 0.568 | |
| nei- g25 | human body | 0.47 | 0.971 | 0.333 | 0 | 0.471 | 0.513 | - | - | - | - |
| ball | 0.507 | 0.961 | 0.449 | nan | 0.507 | 0.252 | - | - | - | - | |
| circle cage | 0.495 | 0.921 | 0.494 | nan | 0.487 | 0.58 | - | - | - | - | |
| square cage | 0.465 | 0.939 | 0.344 | 0 | 0.468 | nan | - | - | - | - | |
| tyre | 0.446 | 0.879 | 0.387 | 0 | 0.433 | 0.564 | - | - | - | - | |
| cube | 0.489 | 0.935 | 0.449 | 0.323 | 0.487 | 0.666 | - | - | - | - | |
| cylinder | 0.392 | 0.863 | 0.322 | 0.634 | 0.388 | nan | - | - | - | - | |
| metal bucket | 0.475 | 0.937 | 0.391 | nan | 0.487 | 0.2 | - | - | - | - | |
| plane | 0.591 | 0.995 | 0.621 | nan | 0.568 | 0.63 | - | - | - | - | |
| rov | 0.46 | 0.965 | 0.35 | 0 | 0.465 | nan | - | - | - | - | |
| all | 0.479 | 0.937 | 0.414 | 0.159 | 0.476 | 0.486 | 0.558 | 0.159 | 0.556 | 0.532 | |
| nei- g5-50 | human body | 0.505 | 0.96 | 0.474 | 0 | 0.508 | 0.508 | - | - | - | - |
| ball | 0.543 | 0.969 | 0.558 | nan | 0.543 | 0.329 | - | - | - | - | |
| circle cage | 0.511 | 0.934 | 0.526 | nan | 0.503 | 0.576 | - | - | - | - | |
| square cage | 0.486 | 0.946 | 0.394 | 0 | 0.489 | nan | - | - | - | - | |
| tyre | 0.505 | 0.926 | 0.446 | 0 | 0.498 | 0.594 | - | - | - | - | |
| cube | 0.508 | 0.946 | 0.505 | 0.343 | 0.507 | 0.603 | - | - | - | - | |
| cylinder | 0.389 | 0.845 | 0.287 | 0.606 | 0.386 | nan | - | - | - | - | |
| metal bucket | 0.496 | 0.925 | 0.508 | nan | 0.506 | 0.276 | - | - | - | - | |
| plane | 0.593 | 0.998 | 0.648 | nan | 0.564 | 0.634 | - | - | - | - | |
| rov | 0.438 | 0.975 | 0.289 | 0 | 0.444 | nan | - | - | - | - | |
| all | 0.498 | 0.942 | 0.464 | 0.158 | 0.495 | 0.503 | 0.575 | 0.173 | 0.573 | 0.573 | |
| nei- p30 | human body | 0.506 | 0.976 | 0.42 | 0 | 0.507 | 0.544 | - | - | - | - |
| ball | 0.535 | 0.968 | 0.521 | nan | 0.535 | 0.303 | - | - | - | - | |
| circle cage | 0.514 | 0.935 | 0.498 | nan | 0.518 | 0.524 | - | - | - | - | |
| square cage | 0.488 | 0.946 | 0.385 | 0 | 0.491 | nan | - | - | - | - | |
| tyre | 0.5 | 0.92 | 0.479 | 0 | 0.492 | 0.586 | - | - | - | - | |
| cube | 0.501 | 0.942 | 0.461 | 0.307 | 0.502 | 0.549 | - | - | - | - | |
| cylinder | 0.352 | 0.785 | 0.207 | 0.655 | 0.348 | nan | - | - | - | - | |
| metal bucket | 0.491 | 0.915 | 0.47 | nan | 0.5 | 0.295 | - | - | - | - | |
| plane | 0.573 | 0.997 | 0.604 | nan | 0.549 | 0.614 | - | - | - | - | |
| rov | 0.429 | 0.977 | 0.243 | 0 | 0.433 | nan | - | - | - | - | |
| all | 0.489 | 0.936 | 0.429 | 0.16 | 0.488 | 0.488 | 0.572 | 0.171 | 0.571 | 0.527 | |
| nei- p5-50 | human body | 0.52 | 0.97 | 0.499 | 0 | 0.519 | 0.559 | - | - | - | - |
| ball | 0.534 | 0.965 | 0.512 | nan | 0.534 | 0.374 | - | - | - | - | |
| circle cage | 0.53 | 0.935 | 0.613 | nan | 0.524 | 0.593 | - | - | - | - | |
| square cage | 0.501 | 0.946 | 0.473 | 0 | 0.503 | nan | - | - | - | - | |
| tyre | 0.505 | 0.904 | 0.477 | 0 | 0.501 | 0.572 | - | - | - | - | |
| cube | 0.511 | 0.935 | 0.504 | 0.362 | 0.51 | 0.624 | - | - | - | - | |
| cylinder | 0.379 | 0.821 | 0.214 | 0.667 | 0.375 | nan | - | - | - | - | |
| metal bucket | 0.49 | 0.912 | 0.468 | nan | 0.497 | 0.435 | - | - | - | - | |
| plane | 0.581 | 0.997 | 0.654 | nan | 0.549 | 0.629 | - | - | - | - | |
| rov | 0.44 | 0.967 | 0.291 | 0 | 0.444 | nan | - | - | - | - | |
| all | 0.499 | 0.935 | 0.47 | 0.172 | 0.496 | 0.541 | 0.578 | 0.171 | 0.575 | 0.607 |
| Denoising | YOLOX | ||||||||||
| class | mAP | mAP_50 | mAP_75 | mAP_s | mAP_m | mAP_l | AR | AR_s | AR_m | AR_l | |
| noisy | human body | 0.563 | 0.977 | 0.562 | 0 | 0.56 | 0.639 | - | - | - | - |
| ball | 0.571 | 0.968 | 0.639 | nan | 0.571 | 0.328 | - | - | - | - | |
| circle cage | 0.537 | 0.917 | 0.598 | nan | 0.529 | 0.63 | - | - | - | - | |
| square cage | 0.519 | 0.956 | 0.502 | 0 | 0.522 | nan | - | - | - | - | |
| tyre | 0.537 | 0.935 | 0.549 | 0 | 0.526 | 0.65 | - | - | - | - | |
| cube | 0.56 | 0.959 | 0.591 | 0.424 | 0.559 | 0.647 | - | - | - | - | |
| cylinder | 0.475 | 0.925 | 0.355 | 0.666 | 0.471 | nan | - | - | - | - | |
| metal bucket | 0.508 | 0.927 | 0.497 | nan | 0.522 | 0.201 | - | - | - | - | |
| plane | 0.663 | 1 | 0.783 | nan | 0.637 | 0.695 | - | - | - | - | |
| rov | 0.516 | 0.977 | 0.472 | 0.1 | 0.519 | nan | - | - | - | - | |
| all | 0.545 | 0.954 | 0.555 | 0.198 | 0.542 | 0.542 | 0.604 | 0.203 | 0.602 | 0.567 | |
| dspnet | human body | 0.559 | 0.978 | 0.591 | 0 | 0.556 | 0.622 | - | - | - | - |
| ball | 0.577 | 0.971 | 0.649 | nan | 0.577 | 0.353 | - | - | - | - | |
| circle cage | 0.528 | 0.923 | 0.567 | nan | 0.528 | 0.555 | - | - | - | - | |
| square cage | 0.514 | 0.953 | 0.504 | 0 | 0.515 | nan | - | - | - | - | |
| tyre | 0.539 | 0.942 | 0.589 | 0 | 0.531 | 0.63 | - | - | - | - | |
| cube | 0.553 | 0.961 | 0.584 | 0.362 | 0.554 | 0.634 | - | - | - | - | |
| cylinder | 0.473 | 0.927 | 0.415 | 0.55 | 0.471 | nan | - | - | - | - | |
| metal bucket | 0.508 | 0.918 | 0.479 | nan | 0.52 | 0.27 | - | - | - | - | |
| plane | 0.657 | 1 | 0.768 | nan | 0.637 | 0.684 | - | - | - | - | |
| rov | 0.551 | 0.988 | 0.541 | 0 | 0.557 | nan | - | - | - | - | |
| all | 0.546 | 0.956 | 0.569 | 0.152 | 0.545 | 0.536 | 0.608 | 0.154 | 0.608 | 0.566 | |
| b2ub- g25 | human body | 0.538 | 0.978 | 0.548 | 0 | 0.538 | 0.584 | - | - | - | - |
| ball | 0.573 | 0.971 | 0.633 | nan | 0.573 | 0.303 | - | - | - | - | |
| circle cage | 0.531 | 0.905 | 0.565 | nan | 0.523 | 0.616 | - | - | - | - | |
| square cage | 0.502 | 0.96 | 0.419 | 0 | 0.505 | nan | - | - | - | - | |
| tyre | 0.498 | 0.928 | 0.509 | 0 | 0.493 | 0.588 | - | - | - | - | |
| cube | 0.543 | 0.957 | 0.566 | 0.363 | 0.539 | 0.695 | - | - | - | - | |
| cylinder | 0.481 | 0.926 | 0.401 | 0.699 | 0.476 | nan | - | - | - | - | |
| metal bucket | 0.481 | 0.917 | 0.401 | nan | 0.493 | 0.203 | - | - | - | - | |
| plane | 0.65 | 0.99 | 0.738 | nan | 0.619 | 0.689 | - | - | - | - | |
| rov | 0.542 | 0.989 | 0.539 | 0 | 0.545 | nan | - | - | - | - | |
| all | 0.534 | 0.952 | 0.532 | 0.177 | 0.53 | 0.525 | 0.595 | 0.177 | 0.593 | 0.554 | |
| b2ub- g5-50 | human body | 0.523 | 0.976 | 0.511 | 0 | 0.52 | 0.595 | - | - | - | - |
| ball | 0.55 | 0.962 | 0.556 | nan | 0.55 | 0.303 | - | - | - | - | |
| circle cage | 0.507 | 0.892 | 0.492 | nan | 0.506 | 0.567 | - | - | - | - | |
| square cage | 0.479 | 0.947 | 0.396 | 0 | 0.482 | nan | - | - | - | - | |
| tyre | 0.484 | 0.899 | 0.459 | 0 | 0.47 | 0.596 | - | - | - | - | |
| cube | 0.533 | 0.963 | 0.477 | 0.363 | 0.533 | 0.625 | - | - | - | - | |
| cylinder | 0.442 | 0.91 | 0.332 | 0.6 | 0.439 | nan | - | - | - | - | |
| metal bucket | 0.484 | 0.924 | 0.357 | nan | 0.497 | 0.227 | - | - | - | - | |
| plane | 0.602 | 1 | 0.714 | nan | 0.583 | 0.627 | - | - | - | - | |
| rov | 0.488 | 0.969 | 0.449 | 0 | 0.491 | nan | - | - | - | - | |
| all | 0.509 | 0.944 | 0.474 | 0.161 | 0.507 | 0.506 | 0.575 | 0.16 | 0.574 | 0.544 | |
| b2ub- p30 | human body | 0.562 | 0.973 | 0.62 | 0 | 0.559 | 0.627 | - | - | - | - |
| ball | 0.579 | 0.97 | 0.679 | nan | 0.58 | 0.328 | - | - | - | - | |
| circle cage | 0.557 | 0.924 | 0.652 | nan | 0.556 | 0.599 | - | - | - | - | |
| square cage | 0.529 | 0.946 | 0.534 | 0 | 0.532 | nan | - | - | - | - | |
| tyre | 0.555 | 0.961 | 0.588 | 0 | 0.546 | 0.65 | - | - | - | - | |
| cube | 0.563 | 0.967 | 0.604 | 0.424 | 0.563 | 0.687 | - | - | - | - | |
| cylinder | 0.481 | 0.938 | 0.392 | 0.665 | 0.478 | nan | - | - | - | - | |
| metal bucket | 0.51 | 0.923 | 0.488 | nan | 0.523 | 0.251 | - | - | - | - | |
| plane | 0.645 | 0.99 | 0.777 | nan | 0.611 | 0.683 | - | - | - | - | |
| rov | 0.534 | 0.979 | 0.532 | 0 | 0.537 | nan | - | - | - | - | |
| all | 0.552 | 0.957 | 0.587 | 0.182 | 0.548 | 0.546 | 0.614 | 0.181 | 0.612 | 0.581 | |
| b2ub- p5-50 | human body | 0.569 | 0.977 | 0.623 | 0 | 0.568 | 0.618 | - | - | - | - |
| ball | 0.587 | 0.973 | 0.663 | nan | 0.587 | 0.353 | - | - | - | - | |
| circle cage | 0.532 | 0.912 | 0.574 | nan | 0.531 | 0.57 | - | - | - | - | |
| square cage | 0.513 | 0.953 | 0.496 | 0 | 0.515 | nan | - | - | - | - | |
| tyre | 0.553 | 0.946 | 0.601 | 0 | 0.548 | 0.632 | - | - | - | - | |
| cube | 0.563 | 0.96 | 0.607 | 0.423 | 0.563 | 0.642 | - | - | - | - | |
| cylinder | 0.476 | 0.941 | 0.366 | 0.634 | 0.472 | nan | - | - | - | - | |
| metal bucket | 0.509 | 0.931 | 0.516 | nan | 0.522 | 0.251 | - | - | - | - | |
| plane | 0.658 | 1 | 0.742 | nan | 0.631 | 0.69 | - | - | - | - | |
| rov | 0.525 | 0.999 | 0.456 | 0.2 | 0.527 | nan | - | - | - | - | |
| all | 0.549 | 0.959 | 0.564 | 0.209 | 0.546 | 0.537 | 0.61 | 0.209 | 0.609 | 0.565 | |
| nei- g25 | human body | 0.43 | 0.917 | 0.353 | 0 | 0.421 | 0.565 | - | - | - | - |
| ball | 0.505 | 0.936 | 0.455 | nan | 0.505 | 0.303 | - | - | - | - | |
| circle cage | 0.451 | 0.877 | 0.393 | nan | 0.442 | 0.565 | - | - | - | - | |
| square cage | 0.438 | 0.93 | 0.337 | 0 | 0.44 | nan | - | - | - | - | |
| tyre | 0.422 | 0.816 | 0.392 | 0 | 0.412 | 0.528 | - | - | - | - | |
| cube | 0.491 | 0.944 | 0.422 | 0.283 | 0.493 | 0.579 | - | - | - | - | |
| cylinder | 0.422 | 0.896 | 0.317 | 0.6 | 0.417 | nan | - | - | - | - | |
| metal bucket | 0.424 | 0.857 | 0.376 | nan | 0.43 | 0.277 | - | - | - | - | |
| plane | 0.649 | 1 | 0.736 | nan | 0.606 | 0.697 | - | - | - | - | |
| rov | 0.511 | 0.978 | 0.387 | 0 | 0.515 | nan | - | - | - | - | |
| all | 0.474 | 0.915 | 0.417 | 0.147 | 0.468 | 0.502 | 0.541 | 0.152 | 0.537 | 0.539 | |
| nei- g5-50 | human body | 0.554 | 0.986 | 0.577 | 0 | 0.552 | 0.601 | - | - | - | - |
| ball | 0.581 | 0.972 | 0.687 | nan | 0.582 | 0.353 | - | - | - | - | |
| circle cage | 0.537 | 0.928 | 0.594 | nan | 0.536 | 0.578 | - | - | - | - | |
| square cage | 0.522 | 0.947 | 0.48 | 0 | 0.524 | nan | - | - | - | - | |
| tyre | 0.541 | 0.952 | 0.546 | 0 | 0.534 | 0.615 | - | - | - | - | |
| cube | 0.553 | 0.967 | 0.578 | 0.379 | 0.553 | 0.657 | - | - | - | - | |
| cylinder | 0.468 | 0.939 | 0.358 | 0.601 | 0.465 | nan | - | - | - | - | |
| metal bucket | 0.512 | 0.935 | 0.45 | nan | 0.524 | 0.276 | - | - | - | - | |
| plane | 0.655 | 1 | 0.756 | nan | 0.623 | 0.695 | - | - | - | - | |
| rov | 0.521 | 0.968 | 0.448 | 0 | 0.524 | nan | - | - | - | - | |
| all | 0.544 | 0.959 | 0.547 | 0.163 | 0.542 | 0.539 | 0.608 | 0.167 | 0.606 | 0.572 | |
| nei- p30 | human body | 0.557 | 0.975 | 0.603 | 0 | 0.556 | 0.615 | - | - | - | - |
| ball | 0.593 | 0.971 | 0.683 | nan | 0.593 | 0.353 | - | - | - | - | |
| circle cage | 0.556 | 0.935 | 0.642 | nan | 0.555 | 0.582 | - | - | - | - | |
| square cage | 0.525 | 0.956 | 0.497 | 0 | 0.527 | nan | - | - | - | - | |
| tyre | 0.547 | 0.946 | 0.587 | 0 | 0.545 | 0.624 | - | - | - | - | |
| cube | 0.565 | 0.972 | 0.568 | 0.423 | 0.563 | 0.662 | - | - | - | - | |
| cylinder | 0.471 | 0.933 | 0.36 | 0.566 | 0.469 | nan | - | - | - | - | |
| metal bucket | 0.51 | 0.926 | 0.48 | nan | 0.524 | 0.227 | - | - | - | - | |
| plane | 0.651 | 1 | 0.746 | nan | 0.635 | 0.675 | - | - | - | - | |
| rov | 0.538 | 0.99 | 0.537 | 0.1 | 0.542 | nan | - | - | - | - | |
| all | 0.551 | 0.96 | 0.57 | 0.182 | 0.551 | 0.534 | 0.612 | 0.187 | 0.612 | 0.567 | |
| nei- p5-50 | human body | 0.559 | 0.989 | 0.582 | 0 | 0.564 | 0.569 | - | - | - | - |
| ball | 0.585 | 0.966 | 0.687 | nan | 0.586 | 0.353 | - | - | - | - | |
| circle cage | 0.53 | 0.911 | 0.578 | nan | 0.529 | 0.561 | - | - | - | - | |
| square cage | 0.517 | 0.945 | 0.474 | 0 | 0.52 | nan | - | - | - | - | |
| tyre | 0.556 | 0.965 | 0.61 | 0 | 0.551 | 0.636 | - | - | - | - | |
| cube | 0.56 | 0.968 | 0.607 | 0.412 | 0.559 | 0.645 | - | - | - | - | |
| cylinder | 0.487 | 0.933 | 0.432 | 0.633 | 0.485 | nan | - | - | - | - | |
| metal bucket | 0.51 | 0.923 | 0.43 | nan | 0.517 | 0.351 | - | - | - | - | |
| plane | 0.658 | 0.989 | 0.785 | nan | 0.627 | 0.693 | - | - | - | - | |
| rov | 0.546 | 0.978 | 0.571 | 0 | 0.55 | nan | - | - | - | - | |
| all | 0.551 | 0.957 | 0.576 | 0.174 | 0.549 | 0.544 | 0.614 | 0.181 | 0.613 | 0.575 |